### v1

# Extremely Amenable Automorphism Groups of Countable Structures

Abstract.
In this paper we address the question: How many pairwise non-isomorphic extremely amenable groups are there which are seprable metrizable or even Polish? We show that there are continuum many such groups. In fact we construct continuum many pairwise non-isomorphic extremely amenable groups as automorphism groups of countable structures. We also consider this classicfication problem from the point of view of descriptive set theory by showing that the class of all extremely amenable closed subgroups of $S_{\infty}$ is Borel and by obtaining some lower bounds for their isomorphism relation in the Borel reducibility hierarchy.
The second and the third authors acknowledge the partial support of their research by the National Natural Science Foundation of China (NSFC) grants 12250710128 and 12271263.
1. Introduction
Ever since Mitchell [17] asked in 1970 about the existence of non-trivial extremely amenable topological groups, many examples of such groups have been discovered.
The first examples of extremely amenable groups were given by Herer–Christensen [11]. They constructed such groups which are abelian Polish groups of the form $L_{0}(\mu,\mathbb{R})$, which is the additive group of all $\mathbb{R}$-valued $\mu$-measurable functions, where $\mu$ is a pathological submeasure. Later, other examples of the more general form $L_{0}(\mu,G)$, where $G$ is a topological group, were shown to be extremely amenable. Among them, Furstenberg–Weiss (in unpublished work) and independently Glasner [9] showed that the group of measurable maps from the unit interval $\mathbb{I}$ with the Lebesgue measure $\lambda$ to the unit circle $\mathbb{T}$ is extremely amenable. Farah–Solecki [5] generalized it by showing that $L_{0}(\mu,G)$ is extremely amenable when $G$ is a compact solvable topological group and $\mu$ is a so-called diffused submeasure. Sabok [28] generalized further to arbitrary solvable topological groups $G$ and diffused submeaures $\mu$.
In a different thread of research, many operator groups were also shown to be extremely amenable. Famously, Gromov–Milman [10] showed that the unitary group $U(H)$ of the infinite dimensional separable Hilbert space $H$ is extremely amenable. Pestov proved the extreme amenability for the group $\mbox{Homeo}_{+}(\mathbb{I})$ of all order-preserving homeomorphisms of the unit interval ([24]) and for the group $\mbox{\rm Iso}(\mathbb{U})$ of all isometries of the universal Urysohn metric space $\mathbb{U}$ ([25]). Giordano–Pestov [8] showed the extreme amenability for $\mbox{\rm Aut}(\mathbb{I},\lambda)$, the group of all measure-preserving automorphisms of the Lebesgue space.
A third class of topological groups known to be extremely amenable in the literature comes from automorphism groups of countable structures. Pestov [24] gave the first example of such a group, which is ${\rm Aut}(\mathbb{Q},<)$, the automorphism group of a countable dense linear order without endpoints. Motivated by the idea in [24], Kechris–Pestov–Todorcevic [15] gave a complete characterization of when the automorphism group of a countable structure is extemely amenable, which is now called the KPT correspondence. The KPT correspondence connects the extreme amenability to the so-called structural Ramsey property of Fraïssé classes and gives a number of examples of extremely amenable groups. Through the work of Nešetřril–Rödl [19] [20] and Nešetřil [18] on Ramsey classes, we now have extremely amenable groups of the form $\mbox{\rm Aut}(M)$, where $M$ is the Fraïssé limit of a Fraïssé class of countable structures. Such structures include ordered graphs, ordered $K_{n}$-free graphs, ordered $\mathbb{Q}$-metric spaces, and ordered $\mathbb{N}$-metric spaces, etc.
In this paper we are concerned with the question: How many pairwise non-isomorphic extremely amenable groups are there? The only result that explicitly addresses this question is a theorem of Pestov in [24], which states that for each infinite cardinal $\kappa$, there is an extremely amenable topological group of weight $\kappa$. Since the weight is an isomorphism invariant, we obtain many pairwise non-isomorphic extremely amenable groups. However, we would like to study this question in the context of more familiar groups, in particular separable metrizable groups and Polish groups. The theorem of Pestov gives only one such example. Despite the diversity of examples mentioned above, many are seprable metrizable and even Polish, we found no results in the literature that differentiate the isomorphism types of the examples, and therefore some work needs to be done before one can claim that there are infinitely many pairwise non-isomorphic extremely amenable groups that are separable metrizable or Polish.
The first main result of this paper confirms that there are continuum many pairwise non-isomorphic extremely amenable groups that are separable and metrizable. For each countable distance value set $\Delta$ (to be defined in Section 2), we consider the $\Delta$-metric Urysohn space $\mathbb{U}_{\Delta}$. The group $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ of all isometries of $\mathbb{U}_{\Delta}$, equipped with the topology of pointwise convergence in metric, is a separable metrizable group. We show the following theorems.
Theorem 1.1.
Let $\Delta$ be a countable distance value set. Then the following hold:
  1. (1)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable iff $\inf\Delta=0$.
  2. (2)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is Polish iff $\inf\Delta>0$.
  3. (3)
    If $\inf\Delta=0$, then $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of either $\mbox{\rm Iso}(\mathbb{U})$ or $\mbox{\rm Iso}(\mathbb{U}_{1})$, where $\mathbb{U}_{1}$ is the Urysohn sphere, i.e., the unit sphere of $\mathbb{U}$.
Theorem 1.2.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as topogoical groups.
  2. (ii)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as abstract groups.
  3. (iii)
    $\Delta$ and $\Lambda$ are equivalent.
Theorem 1.3.
There are continnum many pairwise inequivalent countable distance value sets $\Delta$ that are dense in $(0,+\infty)$. Consequently, there are continuum many pairwise non-isomorphic extremely amenable groups $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ that are separable metrizable.
In fact, all groups in this class are dense in $\mbox{\rm Iso}(\mathbb{U})$; in particular, none of them are Polish.
We then study the case of Polish groups. Here we consider the Fraïssé class of all finite ordered $\Delta$-metric spaces and its Fraïssé limit $\mathbb{U}^{<}_{\Delta}$. Regarding $\mathbb{U}^{<}_{\Delta}$ as a countable structure, the automorphism group $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$, equipped with the topology of pointwise convergence where the underlying space has the discrete topology, is a Polish group. In fact, it is isomorphic to a closed subgroup of $S_{\infty}$, the infinite permutation group. We use the following consequence of a theorem of Nešetřil [18] (for a more explicit restatement of Nešetřil’s theorem, see Theorem 13 of Nguyen Van Thé [22]).
Theorem 1.4.
Let $\Delta$ be a countable distance value set. Then $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ is extremely amenable.
We show the following theorems.
Theorem 1.5.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Lambda})$ are isomorphic as topological groups.
  2. (ii)
    $\Delta$ and $\Lambda$ are equivalent.
Corollary 1.6.
There are continnum many pairwise non-isomorphic extremely amenable Polish groups $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$.
The proofs of Theorems 1.2 and 1.5 follow the outline of the proof of a similar theorem in [4].
In addition, we consider some related questions from the point of view of descriptive set theory. Following a program started by Kechris–Nies–Tent [14], we show that the class of extremely amenable closed subgroups of $S_{\infty}$ is a Borel class, and that the isomorphism relation on this class is an equivalence relation that is not smooth. Using some recent work of Calderoni–Marker–Motto Ros–Shani [2], we prove that a non-optimal lower bound for this equivalence relation in the Borel reducibility hierarchy is $=^{+}$, the equality relation of countable sets of real numbers. From the point of view of descriptive set theory, this is a strengthening of Corollary 1.6.
Theorem 1.7.
The equality relation of countable sets of real numbers, $=^{+}$, is strictly Borel reducible to the isomorphism relation among extremely amenable Polish groups that are isomorphic to closed subgroups of $S_{\infty}$.
The rest of the paper is organized as follows. In Section 2 we define the basic concepts and notions and give a review of the results we need to cite in this paper. In Section 3 we prove Theorems 1.1 and 1.4. In Section 4 we study the notion of equivalence between distance value sets, and prove the first part of Theorem 1.3. In Section 5 we prove Theorems 1.2 and 1.5, which will allow us to derive the second part of Theorem 1.3 and Corollary 1.6. In Section 6 we prove that the class of extremely amenable closed subgroups of $S_{\infty}$ is Borel and state some consequences of our results about the isomorphism relation of extremely amenable groups. Finally, in Section 7 we make some concluding remarks and state some open questions.
Acknowledgment. We thank Filippo Calderoni, Ruiwen Li, Marcin Sabok, Assaf Shani, and Simon Thomas for their inspiring comments on an earlier version of the paper.
2. Preliminaries
In this section we review the basic concepts and notation used in the rest of the paper.
2.1. Topological dynamics
Our standard reference for topological dynamics is [1].
Our topological spaces and topological groups are all Hausdorff. Let $G$ be a topological group and let $X$ be a compact Hausdorff space. Let $a\colon G\times X\to X$ be a continuous action of $G$ on $X$. For $g\in G$ and $x\in X$, we also write $g\cdot x$ for $a(g,x)$. We say that $X$ is a $G$-flow. For $x\in X$, the orbit of $x$, denoted $[x]_{G}$, is the set $\{g\cdot x\colon g\in G\}$. We say that the $G$-flow $X$ is minimal if for every $x\in X$, $[x]_{G}$ is dense in $X$.
Suppose $G$ is a topological group and $X$, $Y$ are two $G$-flows. A continuous map $\varphi\colon X\to Y$ is equivariant or a homomorphism if for all $g\in G$ and $x\in X$, $\varphi(g\cdot x)=g\cdot\varphi(x)$. An isomorphism or a topological conjugacy from $X$ to $Y$ is an equivariant homeomorphism from $X$ onto $Y$. As usual, $X$ and $Y$ are isomorphic or topologically conjugate if there is an isomorphism from $X$ to $Y$.
A general result of topological dynamics states that for every topological group $G$, there is a universal minimal $G$-flow, i.e., a minimal $G$-flow which admits continuous equivariant maps onto all other minimal $G$-flows. Moreover, the universal minimal $G$-flow for a topological group $G$ is uniquely determined up to topological conjugacy.
A topological group $G$ is extremely amenable if every $G$-flow $X$ has a fixed point, i.e., there is $x\in X$ such that $g\cdot x=x$ for all $g\in G$. Equivalently, a topological group $G$ is extremely amenable iff the universal minimal $G$-flow is a singleton. By a theorem of Veech [31], no non-trivial locally compact topological groups can be extremely amenable. Thus extreme amenability is a largeness property for topological groups.
2.2. Fraïssé Theory
Our standard reference for model theory and Fraïssé theory is [13].
A language $\mathcal{L}$ is a set consisting of the following:
  1. (a)
    a collection $(R_{i})_{i\in I}$ of relation symbols; for each $i\in I$, there is an associated arity $n(i)$, which is a positive integer, such that $R_{i}$ is an $n(i)$-ary relation symbol;
  2. (b)
    a collection $(F_{j})_{j\in J}$ of function symbols; for each $j\in J$, there is an associate arity $m(j)$, which is a natural number, such that $F_{j}$ is an $m(j)$-ary function symbol.
The 00-ary function symbols, if any, are also called constant symbols. We say $\mathcal{L}$ is countable if $I$ and $J$ are countable.
If $\mathcal{L}=((R_{i})_{i\in I},(F_{j})_{j\in J})$ is a language, an $\mathcal{L}$-structure is a tuple
$\mathcal{A}=(A;(R_{i}^{\mathcal{A}})_{i\in I},(F_{j}^{\mathcal{A}})_{j\in J})$
where
  1. (i)
    $A$ is a set known as the universe of $\mathcal{A}$ and denoted as $|\mathcal{A}|$;
  2. (ii)
    for each $i\in I$, $R_{i}^{\mathcal{A}}$ is the interpretation of $R_{i}$ in $\mathcal{A}$, which is an $n(i)$-ary relation on $A$, i.e., $R_{i}^{\mathcal{A}}\subseteq A^{n(i)}$;
  3. (iii)
    for each $j\in J$, $F_{j}^{\mathcal{A}}$ is the interpretation of $F_{j}$ in $\mathcal{A}$, which is an $m(j)$-ary function from $A^{m(j)}$ to $A$, i.e., $F_{j}^{\mathcal{A}}\colon A^{m(j)}\to A$.
If $m(j)=0$, $F_{j}^{\mathcal{A}}\in A$ is a constant. $\mathcal{A}$ is finite or countable if $|\mathcal{A}|$ is finite or countable, respectively.
Let $\mathcal{L}$ be a language and let $\mathcal{A}$ and $\mathcal{B}$ be $\mathcal{L}$-structures. By an embedding from $\mathcal{A}$ to $\mathcal{B}$, we mean an injection $f\colon|\mathcal{A}|\rightarrow|\mathcal{B}|$ such that for all $i\in I$ and $a_{1},\dots,a_{n(i)}\in A$,
$R_{i}^{\mathcal{A}}(a_{1},\dots,a_{n(i)})\iff R_{i}^{\mathcal{B}}(f(a_{1}),\dots,f(a_{n(i)})),$
and for all $j\in J$ and $a_{1},\dots,a_{m(j)}\in A$,
$f(F_{j}^{\mathcal{A}}(a_{1},\dots,a_{m(j)}))=F_{j}^{\mathcal{B}}(f(a_{1}),\dots,f(a_{m(j)})).$
If $f$ is the inclusion map, then we say $\mathcal{A}$ is a substructure of $\mathcal{B}$, denoted by $\mathcal{A}\subseteq\mathcal{B}$. If $S\subseteq|\mathcal{B}|$, then the substructure of $\mathcal{B}$ generated by $S$ is the smallest substructure of $\mathcal{B}$ whose universe contains $S$ as a subset. In this case $S$ is called a generating set for the substructure generated by $S$. A substructure is finitely generated if it has a finite generating set.
If there is an embedding from $\mathcal{A}$ to $\mathcal{B}$, we say $\mathcal{A}$ is embeddable to $\mathcal{B}$, denoted by $\mathcal{A}\leq\mathcal{B}$. Finally, if $f$ is bijective, we say it is an isomorphism and if there is an isomorphism between $\mathcal{A}$ and $\mathcal{B}$, we say they are isomorphic, denoted by $\mathcal{A}\cong\mathcal{B}$.
For the rest of the discussion, fix a countable language $\mathcal{L}$. A class $\mathcal{K}$ of finite $\mathcal{L}$-structures is said to be a Fraïssé class if it satisfies the following conditions:
  1. (1)
    Up to isomorphism, there are only countably many different structures in $\mathcal{K}$.
  2. (2)
    There are structures of arbitarily large finite cardinality in $\mathcal{K}$.
  3. (3)
    Hereditary property (HP): If $\mathcal{A}\in\mathcal{K}$, then every finite substructure of $\mathcal{A}$ is in $\mathcal{K}$.
  4. (4)
    Joint embedding property (JEP): If $\mathcal{A},\mathcal{B}\in\mathcal{K}$, then there is a $\mathcal{C}\in\mathcal{K}$ such that $\mathcal{A}\leq\mathcal{C}$ and $\mathcal{B}\leq\mathcal{C}$.
  5. (5)
    Amalgamation property (AP): For any embeddings $f:\mathcal{A}\rightarrow\mathcal{B}$ and $g:\mathcal{A}\rightarrow\mathcal{C}$, there is a $\mathcal{D}\in\mathcal{K}$ and embeddings $r:\mathcal{B}\rightarrow\mathcal{D}$ and $s:\mathcal{C}\rightarrow\mathcal{D}$ such that $r\circ f=s\circ g$.
A countable $\mathcal{L}$-structure $\mathcal{A}$ is locally finite if every finitely generated substructure of $\mathcal{A}$ is finite. The age of $\mathcal{A}$, denoted by ${\rm Age}(\mathcal{A})$, is the class of finite substructures of $\mathcal{A}$. We say $\mathcal{A}$ is ultrahomogeneous if every isomorphism between two finite substructures can be extended to an automorphism of $\mathcal{A}$. We call $\mathcal{A}$ a Fraïssé structure if it is locally finite, countably infinite and ultrahomogeneous.
Fraïssé [6] (c.f. [13]) gave a correspondence between Fraïssé classes and Fraïssé structures, as follows. If $\mathcal{K}$ is a Fraïssé class, then there is a unique Fraïssé structure $\mathcal{A}$, up to isomorphism, such that ${\rm Age}(\mathcal{A})=\mathcal{K}$. This unique structure is called the Fraïssé limit of $\mathcal{K}$, denoted by $\mbox{Flim}\,{\mathcal{K}}$. Conversely, if $\mathcal{A}$ is a Fraïssé structure, then ${\rm Age}(\mathcal{A})$ is a Fraïssé class.
We remark here that the original setting of Fraïssé theory allows a more general definition of Fraïssé classes than we presented above. For example, in the above definition of Fraïssé classes, every mention of finite substructures can be replaced by finitely generated substructures and condition (2) can be omitted. Correspondingly, in the definition of Fraïssé structures, the structure does not need to be locally finite or infinite. We define these notions as above just because the KPT correspondence requires them.
The countable language $\mathcal{L}$ we will consider in this paper are all relational, i.e., they consists only of relation symbols and have no function symbols. In this case, every structure is locally finite, and the age of a structure consists of exactly all finitely generated substructures.
We note the following equivalent condition for a countable structure to be ultrahomogeneous.
Definition 2.1 ([13]).
Let $\mathcal{L}$ be a countable language. A countable $\mathcal{L}$-structure $\mathcal{A}$ is weakly homogeneous if for all $\mathcal{X}$ and $\mathcal{Y}$ in ${\rm Age}(\mathcal{A})$ with $\mathcal{X}\subseteq\mathcal{Y}$ and embedding $f$ from $\mathcal{X}$ to $\mathcal{A}$, $f$ can be extended to an embedding $g$ from $\mathcal{Y}$ to $\mathcal{A}$.
It turns out that for countable structures, ultrahomogeneity and weak homogeneity are equivalent. This equivalent condition will be used later.
2.3. The KPT correspondence
What has become known as the KPT correspondence was developed in Kechris–Pestov–Todorcevic [15]. We review the basic definitions and results here.
A Polish space is a separable, completely metrizable topological space. A Polish group is a topological group whose topology is Polish. An example of a Polish group is the infinite permutation group $S_{\infty}$. $S_{\infty}$ consists of all permutations of $\mathbb{N}$, where the group operation is composition. The topology on $S_{\infty}$ is the pointwise convergence topology, where the topology on $\mathbb{N}$ is discrete. Equivalently, $S_{\infty}$ has the subspace topology inherited from $\mathbb{N}^{\mathbb{N}}$, where $\mathbb{N}^{\mathbb{N}}$ has the product topology. Both the topologies on $\mathbb{N}^{\mathbb{N}}$ and on $S_{\infty}$ are Polish.
Throughout the rest of this subsection we fix a countable language $\mathcal{L}$. For a countable $\mathcal{L}$-structure $\mathcal{A}$, we use ${\rm Aut}(\mathcal{A})$ to denote the group of all automorphisms of $\mathcal{A}$, where the group operation is composition. Considering the discrete topology on $|\mathcal{A}|$, and equipping $\mbox{\rm Aut}(\mathcal{A})$ with the topology of pointwise convergence, $\mbox{\rm Aut}(\mathcal{A})$ becomes a Polish group. In fact, it can be regarded as a closed subgroup of the permutation group $S_{|\mathcal{A}|}$. When $|\mathcal{A}|$ is infinite, $S_{|\mathcal{A}|}$ is isomorphic to $S_{\infty}$. Conversely, if $G$ is a closed subgroup of $S_{\infty}$, then there is a countable structure $\mathcal{A}$ such that $G$ and $\mbox{\rm Aut}(\mathcal{A})$ are isomorphic as topological groups.
A Fraïssé class $\mathcal{K}$ is said to be ordered if there is a binary relation symbol $<$ in $\mathcal{L}$ and, in every structure in $\mathcal{K}$, the interpretaton of $<$ is a linear order. The KPT correspondence unravels a deep connection between the property of a Fraïssé ordered class $\mathcal{K}$ and the extreme amenability of $\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$. This property is called the (structural) Ramsey property, which we define below.
Let $\mathcal{K}$ be a class of finite structures and let $\mathcal{A},\mathcal{B}\in\mathcal{K}$. We use the following notation to denote the set of all substructures of $\mathcal{B}$ which are isomorphic to $\mathcal{A}$:
$\begin{pmatrix}\mathcal{B}\\ \mathcal{A}\end{pmatrix}=\{\mathcal{A^{\prime}}\colon\mathcal{A^{\prime}}\subseteq\mathcal{B},\mathcal{A^{\prime}}\cong\mathcal{A}\}.$
Then for $\mathcal{A}\leq\mathcal{B}\leq\mathcal{C}$ and $k\geq 2$, we use the Erdős-Rado arrow,
$\mathcal{C}\rightarrow(\mathcal{B})^{\mathcal{A}}_{k}$
to mean that for any function $c:\begin{pmatrix}\mathcal{C}\\ \mathcal{A}\end{pmatrix}\rightarrow k$, there is a $\mathcal{B^{\prime}}\in\begin{pmatrix}\mathcal{C}\\ \mathcal{B}\end{pmatrix}$ such that $c$ is constant on $\begin{pmatrix}\mathcal{B^{\prime}}\\ \mathcal{A}\end{pmatrix}.$ Finally, we say $\mathcal{K}$ has the Ramsey property if for any $\mathcal{A}\leq\mathcal{B}\in\mathcal{K}$ and $k\geq 2$, there is a $\mathcal{C}\in\mathcal{K}$ such that $\mathcal{C}\rightarrow(\mathcal{B})^{\mathcal{A}}_{k}$.
Theorem 2.1 (Kechris-Pestov-Todorcevic [15]).
Let $\mathcal{K}$ be a Fraïssé ordered class and let $\mathcal{A}$ be the Fraïssé limit of $\mathcal{K}$. Then the following are equivalent:
  1. (1)
    ${\rm Aut}(\mathcal{A})$ is extremely amenable.
  2. (2)
    $\mathcal{K}$ has the Ramsey property.
The simplest example of a Fraïssé class with the Ramsey property is the class of all finite sets. In fact, the Ramsey property for this class is equivalent to the classical Ramsey theorem. Also, the class of all finite linear orders has the Ramsey property, which is again equivalent to the classical Ramsey theorem; since the Fraïssé limit of this class is the countable dense linear order without endpoints, it follows from the above theorem that ${\rm Aut}(\mathbb{Q},<)$ is extremely amenable. (Historically, Pestov [24] proved the extreme amenability of $\mbox{\rm Aut}(\mathbb{Q},<)$ first, which motivated the above theorem.) Many more examples of Fraïssé ordered classes with the Ramsey property are now known and the automorphism groups of their Fraïssé limits are all extremely amenable by the KPT correspondence. For an excellent survey of these results, see Nguyen Van Thé [23].
We will also use the following strengthening of Theorem 2.1, which is Theorem 3 of Nguyen Van Thé [23].
Theorem 2.2 ([23]).
Let $\mathcal{L}$ be a countable language, let $\mathcal{K}$ be a Fraïssé class of finite $\mathcal{L}$-structures, and let $\mathcal{A}$ be the Fraïssé limit of $\mathcal{K}$. Then the following are equivalent:
  1. (1)
    $\mbox{\rm Aut}(\mathcal{A})$ is extremely amenable.
  2. (2)
    $\mathcal{K}$ has the Ramsey property and every element of $\mathcal{K}$ is rigid, i.e., has trivial automorphism group.
3. The Extreme Amenability of Automorphism Groups
3.1. $\Delta$-metric spaces
We work with the following notion.
Definition 3.1 ([4]).
Let $\Delta$ be a set of positive real numbers. We call $\Delta$ a distance value set if for all $x,y\in\Delta$,
$\min\{x+y,\sup\Delta\}\in\Delta.$
Briefly speaking, a distance value set is exactly an initial segment of a subsemigroup of positive reals. Given a distance value set $\Delta$, a metric space $(X,d)$ is a $\Delta$-metric space if $d(x,y)\in\Delta$ for all $x\neq y\in X$.
Fix a countable distance value set $\Delta$. We can view each countable $\Delta$-metric space as a countable structure, as follows. Let $\mathcal{L}_{\Delta}=\{R_{s}\colon s\in\Delta\}$, where each $R_{s}$ is a binary relation symbol. Given any countable $\Delta$-metric space $(X,d)$, set $R^{X}_{s}(x,y)$ iff $d(x,y)=s$. Thus we turn $(X,d)$ into a countable $\mathcal{L}_{\Delta}$-structure $(X;(R^{X}_{s})_{s\in\Delta})$. Let $\mbox{\rm Aut}(X,d)$, or simply $\mbox{\rm Aut}(X)$ if $d$ is understood, denote the automorphism group $\mbox{\rm Aut}(X;(R^{X}_{s})_{s\in\Delta})$. Then $\mbox{\rm Aut}(X)$ is a Polish group.
However, in this section we will consider another topology, and thus we will use a different notation. Let $\mbox{\rm Iso}(X,d)$ or simply $\mbox{\rm Iso}(X)$ if $d$ is understood, denote the group of all autoisometries of $X$, i.e., $f\in\mbox{\rm Iso}(X)$ if $f\colon X\to X$ satisfies $d(f(x),f(y))=d(x,y)$ for all $x,y\in X$. Equip $\mbox{\rm Iso}(X)$ with the topology of pointwise convergence in metric, i.e., the topology given by basic open sets of the form
$\{f\in\mbox{\rm Iso}(X)\colon d(f(x_{1}),y_{1})<\epsilon_{1},\dots,d(f(x_{n}),y_{n})<\epsilon_{n}\}$
where $x_{1},\dots,x_{n},y_{1},\dots,y_{n}\in X$ and $\epsilon_{1},\dots,\epsilon_{n}$ are positive real numbers. Then $\mbox{\rm Iso}(X)$ is a separable metrizable group, and it is Polish when $d$ is a complete metric on $X$. Let $(\bar{X},\bar{d})$ be the completion of $(X,d)$. Then $\mbox{\rm Iso}(X)$ is isomorphic to a subgroup of $\mbox{\rm Iso}(\bar{X})$, and $\mbox{\rm Iso}(\bar{X})$ is Polish.
Importantly, note that $\mbox{\rm Aut}(X)$ and $\mbox{\rm Iso}(X)$ are isomorphic as abstract groups.
Fix a countable distance value set $\Delta$. Let $\mathcal{K}_{\Delta}$ be the class of all finite $\Delta$-metric spaces. Then $\mathcal{K}_{\Delta}$ is a Fraïssé class. We denote the Fraïssé limit of $\mathcal{K}_{\Delta}$ by $\mathbb{U}_{\Delta}$ and call it the $\Delta$-Urysohn space. In the case that $\Delta=\mathbb{Q}$, the resulting Fraïssé limit $\mathbb{U}_{\mathbb{Q}}$ is the well-known rational Urysohn space and has been denoted by $\mathbb{QU}$ in the literature. The completion of $\mathbb{QU}$ is the universal Urysohn space and has been denoted by $\mathbb{U}$ in the literature. $\mathbb{U}$ is universal for all separable complete metric spaces, i.e., for any separable complete metric space $(X,d)$, there is an isometric embedding from $X$ into $\mathbb{U}$.
3.2. The extreme amenability of $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$
To apply the KPT correspondence, we need to consider a Fraïssé ordered class. Let $\Delta$ be a countable distance value set. Let $\mathcal{K}_{\Delta}^{<}$ be the class of all finite ordered $\Delta$-metric spaces, i.e., each element of $\mathcal{K}_{\Delta}^{<}$ is a structure $(X;(R_{s}^{X})_{s\in\Delta},<^{X})$, where $<^{X}$ is a linear order on $X$. It is easy to see that $\mathcal{K}_{\Delta}^{<}$ is a Fraïssé (ordered) class. We denote its Fraïssé limit by $\mathbb{U}_{\Delta}^{<}$. It can be shown that $\mathbb{U}_{\Delta}^{<}\cong(\mathbb{U}_{\Delta},<^{\mathbb{U}_{\Delta}})$ where $<^{\mathbb{U}_{\Delta}}$ is a linear order isomorphic to $(\mathbb{Q},<)$ (see, e.g., Proposition 5.2 of [15]).
The following is a theorem of Nešetřil [18]; Ngyen Van Thé [22] stated the result more explicitly (as Theorem 13 of [22]) and gave a full proof.
Theorem 3.1 (Nešetřil).
Let $\Delta$ be a countable distance value set. Then $\mathcal{K}_{\Delta}^{<}$ has the Ramsey property.
By the KPT correspondence, we get that $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$ is extremely amenable, which is Theorem 1.4.
3.3. The extreme amenability of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
In this subsection we prove Theorem 1.1. We first consider a countable distance value set $\Delta$ with $\inf\Delta=0$. Note that this condition is equivalent to that $\Delta$ is dense in $(0,\sup\Delta)$. In this case we show that $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable. Our proof follows the outline of the proof of Theorem 6.17 of [15], which argues the extreme amenability of ${\rm Iso}(\mathbb{U})$ from the extreme amenability of ${\rm Aut}(\mathbb{U}_{\mathbb{Q}}^{<})$. We use the following basic fact (Lemma 6.18 of [15]).
Lemma 3.2 ([15]).
Let $G,H$ be topological groups and let $f$ be a continuous group homomorphism from $G$ into $H$ with dense image. If $G$ is extremely amenable, so is $H$.
Note that every automorphism of $\mathbb{U}_{\Delta}^{<}$ is actually an isomertry of $\mathbb{U}_{\Delta}$. So we consider the identity map from ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ to ${\rm Iso}(\mathbb{U}_{\Delta})$. This map is obviously continuous. The following proposition implies that $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$.
Proposition 3.3.
Suppose $\Delta$ is a countable distance value set with $\inf\Delta=0$. Given $x_{1},\dots,x_{n},y_{1},\dots,y_{n}\in\mathbb{U}_{\Delta}$ such that $x_{i}\mapsto y_{i}$, $1\leq i\leq n$, is a partial isometry, and given $\epsilon>0$, there exist $y_{1}^{\prime},\dots,y_{n}^{\prime}\in\mathbb{U}_{\Delta}$ so that $x_{i}\mapsto y_{i}^{\prime}$, $1\leq i\leq n$, is an order-preserving (with respect to $<^{\mathbb{U}_{\Delta}}$) partial isometry, and for all $1\leq i\leq n$, $d(y_{i},y_{i}^{\prime})<\epsilon$.
Proof.
Let $Z=\{y_{1},\dots,y_{n},z_{1},\dots,z_{n}\}$ be a finite ordered $\Delta$-metric space with the following properties:
  1. (i)
    $d_{Z}(y_{i},y_{j})=d_{Z}(z_{i},z_{j})=d_{\mathbb{U}_{\Delta}}(y_{i},y_{j})$ for $1\leq i<j\leq n$,
  2. (ii)
    $d_{Z}(y_{i},z_{j})=\delta+d_{\mathbb{U}_{\Delta}}(y_{i},y_{j})$ for $1\leq i\leq j\leq n$,
  3. (iii)
    $y_{i}<^{Z}y_{j}$ iff $y_{i}<y_{j}$ for $1\leq i<j\leq n$,
  4. (iv)
    $y_{i}<^{Z}z_{j}$ for $1\leq i,j\leq n$,
  5. (v)
    $z_{i}<^{Z}z_{j}$ iff $x_{i}<x_{j}$ for $1\leq i<j\leq n$,
where $\delta\in\Delta$ and $\delta<\epsilon$. Note that $(Z,d_{Z})$ is a metric space. Thus, by the universality and the ultrahomogeneity of $\mathbb{U}_{\Delta}^{<}\cong(\mathbb{U}_{\Delta},<^{\mathbb{U}_{\Delta}})$, there are $\{y_{i}^{\prime}\}_{i=1}^{n}$ in $\mathbb{U}_{\Delta}$ such that $y_{i}\mapsto y_{i}$ and $z_{i}\mapsto y_{i}^{\prime}$ is an order-preserving isometry from $Z$ into $\mathbb{U}_{\Delta}^{<}$. These $y_{i}^{\prime}$ for $1\leq i\leq n$ are as desired. ∎
We are now ready to prove Theorem 1.1, which we restate below.
Theorem 3.4.
Let $\Delta$ be a countable distance value set. Then:
  1. (1)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable iff $\inf\Delta=0$.
  2. (2)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is Polish iff $\inf\Delta>0$.
  3. (3)
    If $\inf\Delta=0$, then ${\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of either ${\rm Iso}(\mathbb{U})$ or $\mbox{\rm Iso}(\mathbb{U}_{1})$.
Proof.
For (1) first assume that $\inf\Delta=0$. Then from Proposition 3.3, we conclude that $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$ is dense in $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$. Now by Theorems 3.1 and 2.1, $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ is extremely amenable. Thus $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable by Lemma 3.2.
On the other hand, assume $\inf\Delta>0$. Then $\mathbb{U}_{\Delta}$ has discrete topology, and it follows that $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ are isomorphic as topological groups. It is clear that not every finite $\Delta$-metric space is rigid. By Theorem 2.2, $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ is not extremely amenable. Hence $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is not extremely amenable.
Next we prove (3). It was proved in [3] that if $\Delta=\mathbb{Q}$, then $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U})$. The argument in [3] can be generalized to prove that if $\Delta$ is dense in $(0,+\infty)$ then ${\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of ${\rm Iso}(\mathbb{U})$. Then same argument also shows that when $\inf\Delta=0$ and $\Delta$ is bounded, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{a})$, where $a=\sup\Delta$ and $\mathbb{U}_{a}$ is a sphere of $\mathbb{U}$ with diameter $a$. However, for any $0<a<\infty$, $\mbox{\rm Iso}(\mathbb{U}_{a})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$ are isomorphic. Hence in this case $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{1})$.
(2) If $\inf\Delta>0$, then ${\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ are isomorphic as topological groups, while the latter is Polish. Hence $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is Polish. On the other hand, suppose $\inf\Delta=0$. First suppose $\Delta$ is unbounded, i.e., $\Delta$ is dense in $(0,+\infty)$. Then by the above argument ${\rm Iso}(\mathbb{U}_{\Delta})$ is dense in ${\rm Iso}(\mathbb{U})$. Note that $\mbox{\rm Iso}(\mathbb{U})$ is Polish. So if ${\rm Iso}(\mathbb{U}_{\Delta})$ was Polish, it would be a closed subgroup of $\mbox{\rm Iso}(\mathbb{U})$ and thus $\mbox{\rm Iso}(\mathbb{U}_{\Delta})=\mbox{\rm Iso}(\mathbb{U})$. This is impossible since we can easily find an isomertry of $\mathbb{U}$ not setwise fixing $\mathbb{U}_{\Delta}$. Next suppose $\Delta$ is bounded and let $a=\sup\Delta$. Then by the above argument $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U}_{a})$. If $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ was Polish, we again have $\mbox{\rm Iso}(\mathbb{U}_{\Delta})=\mbox{\rm Iso}(\mathbb{U}_{a})$, which yields a similar contradiction. ∎
By Theorem 3.4 (1) and (2), for any countable distance value set $\Delta$, the separable metrizable group ${\rm Iso}(\mathbb{U}_{\Delta})$ can never be both Polish and extremely amenable. When $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is not Polish, one typically takes the metric completion of $\mathbb{U}_{\Delta}$ and considers its isometry group, which is Polish. Vershik [32] (Theorem 3) has shown that if $\Delta$ is dense in $(0,+\infty)$, then $\mathbb{U}_{\Delta}$ is dense in $\mathbb{U}$, hence the metric completion of $\mathbb{U}_{\Delta}$ is $\mathbb{U}$. Similarly, if $\Delta$ is bounded and dense in $(0,\sup\Delta)$, then the metric completion of $\mathbb{U}_{\Delta}$ is just $\mathbb{U}_{a}$ where $a=\sup\Delta$. Thus, clause (3) of Theorem 3.4 says that the above-mentioned consideration obtains essentially only two Polish groups, $\mbox{\rm Iso}(\mathbb{U})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$, both known to be extremely amenable by Pestov [25].
4. Equivalence of Distance Value Sets
In this section, we deviate from the extremely amenable groups and study a notion of equivalence between distance value sets.
Definition 4.1 ([4]).
Let $\Delta$ and $\Lambda$ be distance value sets.
  1. (i)
    A triple $(x,y,z)$ of positive real numbers is a $\Delta$-triangle if $x,y,z\in\Delta$ and $|x-y|\leq z\leq x+y$.
  2. (ii)
    We say that $\Delta$ and $\Lambda$ are equivalent if there is a bijection $f\colon\Delta\to\Lambda$ such that for any positive real numbers $x,y,z$, $(x,y,z)$ is a $\Delta$-triangle iff $(f(x),f(y),f(z))$ is a $\Lambda$-triangle.
Let $\Delta$ and $\Lambda$ be equivalent distance value sets. Assume that $\inf\Delta=0$. Suppose $f:\Delta\rightarrow\Lambda$ witnesses that they are equivalent. Fix a sequence $(\delta_{i})_{i\in\mathbb{N}}$ in $\Delta$ such that $\delta_{i+1}<{\delta_{i}}/{2}$ for all $i$.
Lemma 4.1.
$f(\delta_{i+1})<{f(\delta_{i})}/{2}$ for all $i$. Consequently, $\inf\Lambda=0$.
Proof.
For each $i$, consider the triple $(\delta_{i+1},\delta_{i+1},\delta_{i})$. This triple is not a $\Delta$-triangle since $\delta_{i}>\delta_{i+1}+\delta_{i+1}$. By the property of $f$, the triple $(f(\delta_{i+1}),f(\delta_{i+1}),f(\delta_{i}))$ is not a $\Lambda$-triangle. But then the only failed triangle inequality must be $f(\delta_{i})>f(\delta_{i+1})+f(\delta_{i+1})$. ∎
Lemma 4.2.
$f$ is continuous.
Proof.
Fix $x\in\Delta$. For any $\epsilon>0$, find an $i$ such that $\delta_{i}<x$ and $f(\delta_{i})<\epsilon$. For any $y\in\Delta$ with $|x-y|<\delta_{i}$, $(x,y,\delta_{i})$ is a $\Delta$-triangle. Thus $(f(x),f(y),f(\delta_{i}))$ is a $\Lambda$-triangle. In particular, $|f(x)-f(y)|<f(\delta_{i})<\epsilon$. ∎
The above lemma actually shows that for all positive $\delta$, $f$ is uniformly continuous on $\Delta^{>\delta}\triangleq\Delta\cap(\delta,+\infty)$. The following lemma shows that $f$ is in fact linear. In the proof we use notation such as $\Delta^{>\delta}$, $\Delta^{<\delta}$, and $\Delta^{[a,b]}$ with obvious meanings.
Lemma 4.3.
If $x,y,x+y\in\Delta$, then $f(x+y)=f(x)+f(y)$.
Proof.
Fix such $x,y\in\Delta$. We claim that
  1. (a)
    $f(\Delta^{[|x-y|,x+y]})\subseteq\Lambda^{[|f(x)-f(y)|,f(x)+f(y)]}$,
  2. (b)
    $f(\Delta^{>x+y})\subseteq\Lambda^{>f(x)+f(y)}\mbox{ or }f(\Delta^{>x+y})\subseteq\Lambda^{<|f(x)-f(y)|}$,
  3. (c)
    $f(\Delta^{<|x-y|})\subseteq\Lambda^{>f(x)+f(y)}\mbox{ or }f(\Delta^{<|x-y|})\subseteq\Lambda^{<|f(x)-f(y)|}$.
Let $z\in\Delta$. (a) is equivalent to saying that if $(x,y,z)$ is a $\Delta$-triangle, then $(f(x),f(y),f(z))$ is a $\Lambda$-triangle. If $z>x+y$, then $(x,y,z)$ is not a $\Delta$-triangle. Thus $(f(x),f(y),f(z))$ is not a $\Lambda$-triangle, and so either $f(z)>f(x)+f(y)$ or $f(z)<|f(x)-f(y)|$. Similarly for $z<|x-y|$.
To prove (b), assume $z_{1},z_{2}>x+y$ such that $f(z_{1})<|f(x)-f(y)|$ and $f(z_{2})>f(x)+f(y)$. Without loss of generality assume $z_{1}<z_{2}$ (the argument for the case $z_{1}>z_{2}$ is similar). By the continuity of $f$ and the density of both $\Delta$ and $\Lambda$, we may find $z_{1}=a_{1},\dots,a_{n}=z_{2}$ in $\Delta$ such that $a_{k}>x+y$ and $|f(a_{k+1})-f(a_{k})|<f(x)+f(y)-|f(x)-f(y)|$ for all $1\leq k\leq n$. Then there must be an $a_{k}$ with $|f(x)-f(y)|\leq f(a_{k})\leq f(x)+f(y)$, contradicting $a_{k}>x+y$. This proves (b).
A similar argument for $z_{1},z_{2}<|x-y|$ proves (c).
Since $f$ is a bijection from $\Delta$ to $\Lambda$, the claimed inclusions are actually equalities. By considering sufficiently small $\delta_{i}$, we see that $f(\Delta^{<|x-y|})=\Lambda^{<|f(x)-f(y)|}$, and thus $f(\Delta^{>x+y})=\Lambda^{>f(x)+f(y)}$. By the continuty of $f$ again, we have $f(x+y)=f(x)+f(y)$. ∎
Lemma 4.4.
Let $x,t\in\Delta$. We have $f(x)/x=f(t)/t$.
Proof.
For each $k>0$, find $0<\delta_{k}<1/2^{k}$ by the continuity of $f$ such that for all $z\in\Delta$, $|t-z|<\delta_{k}$ implies $|f(t)-f(z)|<1/k$. Now for each $k$, find positive rational number $n_{k}/m_{k}$ such that $20<t-(n_{k}/m_{k})x<\delta_{k}/$ 2. Note that by the density of $\Delta$ and bijectivity of $f$, $\delta_{k}\to 0$ as $k\to\infty$. Thus $n_{k}/m_{k}\to t/x$ as $k\to\infty$.
Fix $k$. By the density of $\Delta$, find an increasing sequence $(t_{j})_{j\in\mathbb{N}}$ with $t_{j}\in\Delta$ and $t_{j}\to x/m_{k}$ as $j\to\infty$. Then $m_{k}t_{j}\to x$ as $j\to\infty$, and $m_{k}t_{j}<x\in\Delta$. Thus $m_{k}t_{j}\in\Delta$ and $m_{k}f(t_{j})=f(m_{k}t_{j})\to f(x)$, which gives $f(t_{j})\to f(x)/m_{k}$. On the other hand, $n_{k}t_{j}<(n_{k}/m_{k})x<t\in\Delta$ and $n_{k}t_{j}\to(n_{k}/m_{k})x$ as $j\to\infty$. Hence $n_{k}t_{j}\in\Delta$ and $f(n_{k}t_{j})=n_{k}f(t_{j})\to(n_{k}/m_{k})f(x)$ as $j\to\infty$. Thus we can find a $j$ such that $|n_{k}t_{j}-(n_{k}/m_{k})x|<\delta_{k}/2$ and $|f(n_{k}t_{j})-(n_{k}/m_{k})f(x)|<1/k$. For this fixed $j$, we have $|t-n_{k}t_{j}|\leq|t-(n_{k}/m_{k})x|+|n_{k}t_{j}-(n_{k}/m_{k})x|<\delta_{k}$. Thus $|f(t)-f(n_{k}t_{j})|<1/k$, and consequently $|f(t)-(n_{k}/m_{k})f(x)|<2/k$.
Letting $k\to\infty$, we get $f(t)=(t/x)f(x)$ and thus $f(x)/x=f(t)/t$. ∎
Thus we have shown the following theorem.
Theorem 4.5.
Let $\Delta$ and $\Lambda$ be distance value sets with $\inf\Delta=\inf\Lambda=0$. Then the following are equivalent:
  1. (1)
    $\Delta$ and $\Lambda$ are equivalent.
  2. (2)
    There is a positive real number $\alpha$ such that $\Lambda=\alpha\cdot\Delta$.
Moreover, if $\Delta$ and $\Lambda$ are equivalent, then every witness of their equivalence has the form of multiplication by a positive real number.
We consider a special kind of countable distance value sets for which the equivalence is more explicit.
Definition 4.2.
Let $\alpha$ be a positve irrational number. Let
$\Delta_{\alpha}=\{p\alpha+q:p,q\in\mathbb{Q}\}\cap(0,+\infty).$
Define an action of $GL_{2}(\mathbb{Q})$ on the space of irrational numbers by the linear fractional transfomation:
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot\alpha=\frac{a\alpha+b}{c\alpha+d}.$
We denote the equivalence relation induced by this action by $E_{GL_{2}(\mathbb{Q})}^{IR}$.
For any positive irrational number $\alpha$, $\Delta_{\alpha}$ is a countable distance value set with $\inf\Delta_{\alpha}=0$. The equivalence between distance value sets of the form $\Delta_{\alpha}$ can be characterized by $E_{GL_{2}(\mathbb{Q})}^{IR}$, as the following theorem shows.
Theorem 4.6.
Let $\alpha,\beta$ be positive irrational numbers. Then the following are equivalent:
  1. (1)
    $\Delta_{\alpha}$ and $\Delta_{\beta}$ are equivalent.
  2. (2)
    $\alpha$ and $\beta$ are $E_{GL_{2}(\mathbb{Q})}^{IR}$-equivalent.
Proof.
For the implication of (2) from (1), let $r\in\mathbb{R}^{+}$ be such that the function defined by $f(t)=r\cdot t$ witnesses the equivalence of $\Delta_{\alpha}$ and $\Delta_{\beta}$. Then
$\beta=\frac{\beta}{1}=\frac{f(\beta)}{f(1)}=\frac{a\alpha+b}{c\alpha+d},$
where $a,b,c,d\in\mathbb{Q}$. Note that the matrix formed by $a,b,c,d$ is invertible; otherwise there is an $s\in\mathbb{Q}$ such that $a=sc$ and $b=sd$, and then $\beta=s$, contradicting that $\beta$ is irrational.
For the other implication, suppose $\alpha=({a\beta+b})/({c\beta+d})$, where $a,b,c,d\in\mathbb{Q}$. We can assume $c\beta+d>0$. It is easy to verify that the function $f(t)=(c\beta+d)t$ is a triangle-preserving bijection from $\Delta_{\alpha}$ to $\Delta_{\beta}$. ∎
Corollary 4.7.
There are continuum many positive irrational numbers which are pairwise $E_{GL_{2}(\mathbb{Q})}^{IR}$-inequivalent. Consequently, there are continnum many pairwise inequivalent countable distance value sets each without a positive lower bound.
Proof.
Note that each $E_{GL_{2}(\mathbb{Q})}^{IR}$-equivalence class is countable. Thus there are continuum many pairwise $E_{GL_{2}(\mathbb{Q})}^{IR}$-inequivalent positive irrational numbers. ∎
5. Isomorphism Types of Extremely Amenable Groups
5.1. Isomorphism types of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
In this subsection we show that there are continuum many pairwise non-isomorphic extremely amenable groups that are separable metrizable. The class of such groups are of the form $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ for countable distance value sets $\Delta$ with $\inf\Delta=0$. We use the following theorem from [4].
Theorem 5.1 ([4]).
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    ${\rm Aut}(\mathbb{U}_{\Delta})$ and ${\rm Aut}(\mathbb{U}_{\Lambda})$ are isomorphic as abstract groups.
  2. (ii)
    ${\rm Aut}(\mathbb{U}_{\Delta})$ and ${\rm Aut}(\mathbb{U}_{\Lambda})$ are isomorphic as topological groups.
  3. (iii)
    $\Delta$ and $\Lambda$ are equivalent.
Our result is similar. The following is a restatement of Theorem 1.2.
Theorem 5.2.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ and ${\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as abstract groups.
  2. (ii)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ and ${\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as topological groups.
  3. (iii)
    $\Delta$ and $\Lambda$ are equivalent.
Proof.
To see that (i) implies (iii), the key oberservation is that $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ are identical abstract groups for any $\Delta$. Thus this implication follows from the implication of (iii) from (i) in Theorem 5.1.
The implication of (i) from (ii) is obvious.
To see that (iii) implies (ii), fix a bijection $f$ from $\Delta$ to $\Lambda$ which witnesses the equivalence of the two distance value sets. Let $d$ be the metric on $\mathbb{U}_{\Delta}$. Define a new metric $d_{f}$ on $\mathbb{U}_{\Delta}$ by setting $d_{f}(x,y)=f(d(x,y))$. Then $(\mathbb{U}_{\Delta},d_{f})$ is isometric to $\mathbb{U}_{\Lambda}$, since $(\mathbb{U}_{\Delta},d_{f})$ is a Fraïssé structure and its age is exactly the class of all finite $\Lambda$-metric spaces. Now the identity map from $(\mathbb{U}_{\Delta},d)$ to $(\mathbb{U}_{\Delta},d_{f})$ induces a topological isomorphism from $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ to $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$. ∎
Corollary 5.3.
There are continuum many pairwise non-isomorphic extremely amenable groups which are separable metrizable but not Polish.
Proof.
This follows immediately from Theorem 5.2, Corollary 4.7, and Theorem 3.4. ∎
5.2. Isomorphism types of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$
In this subsection we show that there are continuum many pairwise non-isomorphic extremely amenable Polish groups. This is done by considering $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ for countable distance value sets $\Delta$. Our main theorem of this subsection is the following.
Theorem 5.4.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ and ${\rm Aut}(\mathbb{U}_{\Lambda}^{<})$ are isomorphic as topological groups.
  2. (ii)
    $\Delta$ and $\Lambda$ are equivalent.
Granting the theorem, we immediately get the following corollary from Corollary 4.7.
Corollary 5.5.
There are continuum many pairwise non-isomorphic extremely amenable Polish groups which are automorphism groups of countable structures, or equivalently closed subgroups of $S_{\infty}$.
The rest of this subsection is devoted to a proof of Theorem 5.4. The outline of the proof follows the proof of Theorem 5.1 in [4].
The argument for the implication of (i) from (ii) is similar to the proof of Theorem 5.2.
To prove the other direction, we need some results due to Slutsky [29]. We work with the following concept from [29] in the Fraïssé theory of finite metric spaces.
Definition 5.1.
Let $\mathcal{A},\mathcal{B},\mathcal{C}$ be finite metric spaces. For simplicity assume $\mathcal{A}$ is a substructure of both $\mathcal{B}$ and $\mathcal{C}$. Also assume $|\mathcal{B}|\cap|\mathcal{C}|=|\mathcal{A}|$. Define a finite metric space $\mathcal{D}$ as follows. Let $|\mathcal{D}|$ = $|\mathcal{B}|\cup|\mathcal{C}|$. For $x,y\in|\mathcal{D}|$, define $d_{\mathcal{D}}(x,y)$ as follows. If $x,y\in|\mathcal{B}|$, then let $d_{\mathcal{D}}(x,y)=d_{\mathcal{B}}(x,y)$. Similarly, if $x,y\in\mathcal{C}$, then let $d_{\mathcal{D}}(x,y)=d_{\mathcal{C}}(x,y)$. If $x\in|\mathcal{B}|$ and $y\in|\mathcal{C}|$, then let
$d_{\mathcal{D}}(x,y)=\left\{\begin{array}[]{ll}\min\{d_{\mathcal{B}}(x,z)+d_{\mathcal{C}}(z,y)\colon z\in|\mathcal{A}|\},&\mbox{ if $|\mathcal{A}|\neq\varnothing$,}\\ \mbox{diam}(\mathcal{B})+\mbox{diam}(\mathcal{C}),&\mbox{ if $|\mathcal{A}|=\varnothing$.}\end{array}\right.$
Then $\mathcal{D}$ is called the free amalgam of $\mathcal{B}$ and $\mathcal{C}$ over $\mathcal{A}$.
Since (ordered) metric spaces can be regarded as relational structures, for an (ordered) metric space $\mathcal{B}$, a substructure $\mathcal{A}$ of $\mathcal{B}$ generated by a set $A\subseteq|\mathcal{B}|$ satisfies $|\mathcal{A}|=A$. Thus we will identify $\mathcal{A}$ with $A$, and speak of $A$ as a substructure. This does not cause confusion, and will simplify our notation.
Definition 5.2.
Let $A$ be a metric space and $p$ a partial isometry on $A$. A point $x\in{\rm dom}(p)$ is called periodic if there is a natural number $n>0$ such that
$x,p(x),\dots,p^{n}(x)\in{\rm dom}(p)\ {\rm and}\ p^{n}(x)=x.$
The set of all periodic points of $p$ is denoted by $Z(p)$. A point $x\in{\rm dom}(p)$ is called a fixed point if $p(x)=x$. The set of all fixed points of $p$ is denoted by $F(p)$.
Lemma 5.6 ([29]).
Let $D$ be a finite ordered metric space. Let $p_{1}$ and $p_{2}$ be order-preserving partial isometries of $D$ such that $Z(p_{1})=F(p_{1})$ and $Z(p_{2})=F(p_{2})$. Suppose $F(p_{1})\cap F(p_{2})\neq\emptyset$. Then there exist a finite ordered metric space $D^{\prime}$ extending $D$, order-preserving partial isometries $q_{1}$ and $q_{2}$ of $D^{\prime}$ extending $p_{1},p_{2}$ respectively, and an element $w\in\mathsf{F}(s,t)$, the free group generated by ${s,t}$, such that
  1. (1)
    $Z(q_{1})=Z(p_{1})$ and $Z(q_{2})=Z(p_{2})$;
  2. (2)
    $D\subseteq{\rm dom}(q_{1})$;
  3. (3)
    ${\rm dom}(q_{1})\cup w(q_{1},q_{2})(D)$ is the free amalgam of ${\rm dom}(q_{1})$ and $w(q_{1},q_{2})(D)$ over $F(p_{1})\cap F(p_{2})$.
Moreover, the distances in $D^{\prime}$ are taken from the additive semigroup generated by the distances in $D$.
This result was originally proved for metric spaces as Theorem 4.12 of [29]. Nevertheless, as Slutsky remarked (Remark 4.13 of [29]), the above ordered version also holds.
We use the following notation in our further discussions. For a countable structure $\mathcal{A}$ and subset $B\subseteq|\mathcal{A}|$, we use ${\rm Aut}_{B}(\mathcal{A})$ to denote the subgroup of $\mbox{\rm Aut}(\mathcal{A})$ consisting of the automorphisms of $\mathcal{A}$ which fix each element of $B$. Note that if $B$ is finite, then ${\rm Aut}_{B}(\mathcal{A})$ is an open subgroup of $\mbox{\rm Aut}(\mathcal{A})$.
The following two lemmas are key technical results in our proof. The combination of their unordered versions was a theorem of Slutsky [29].
Lemma 5.7.
Let $\Delta$ be a countable distance value set, and let $A,B\subseteq\mathbb{U}_{\Delta}^{<}$ be finite subsets. Suppose $A\cap B\neq\varnothing$. Then
$\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle\ is\ dense\ in\ {\rm Aut}_{A\cap B}(\mathbb{U}_{\Delta}^{<}).$
Proof.
For notational simplicity, we assume $A\cap B=\{x_{0}\}$. Suppose $A=\{a_{0},\dots,a_{n}\}$ and $B=\{b_{0},\dots,b_{m}\}$. Again, for definiteness and notational simplicity, assume $m=n$ and $a_{0}=b_{0}=x_{0}<a_{1}<b_{1}<\dots<a_{n}<b_{n}$. Note that in general, $A\cap B=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{0}<x_{1}<\cdots<x_{k}$ and the rest of the points in $A,B$ can be in the gaps between $x_{i}$s including below $x_{0}$ and above $x_{k}$. For the general case, the argument is similar.
Let $C$ be a finite subset of $\mathbb{U}_{\Delta}^{<}$ and let $p\in{\rm Aut}_{A\cap B}(\mathbb{U}_{\Delta}^{<})$. It suffices to show that there is a $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ such that $p\upharpoonright C=q\upharpoonright C$. Without loss of generality, we may assume $A\subseteq C$ and $B\subseteq C$. Let $D=C\cup p(C)$. We find (to be explicitly defined later)
$q=w(q_{1},q_{2})^{-1}g_{1}^{-1}\dots g_{2n}^{-1}rg_{2n}\dots g_{1}w(q_{1},q_{2}),$
where $g_{2i-1}\in{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})$ and $g_{2i}\in{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})$. Roughly speaking, we start from an isometric copy of $D$ in $\mathbb{U}_{\Delta}^{<}$ where all the new points (other than $A\cap B$) are appropriately distributed between $x_{i}$s. Using the combination of $g_{i}$s, we propagate the points in the initial copy of $D$ to create the correct ordering of points in the final copy of $D$.
As a substructure of $\mathbb{U}^{<}_{\Delta}$, $D$ is a finite ordered metric space. Define two order-preserving partial isometries $p_{1},p_{2}$ of $D$ by letting $p_{1}(x)=x$ for all $x\in A$ and $p_{2}(x)=x$ for all $x\in B$. Now applying Lemma 5.6 to $D$ and $p_{1},p_{2}$, we get a finite ordered metric space $D^{\prime}$ extending $D$, order-preserving partial isometries $q_{1},q_{2}$ of $D^{\prime}$ which are extensions of $p_{1},p_{2}$, respectively, and a word $w\in\mathsf{F}(s,t)$ such that
  1. (1)
    $Z(q_{1})=Z(p_{1})=F(p_{1})=A$ and $Z(q_{2})=Z(p_{2})=F(p_{2})=B$;
  2. (2)
    $D\subseteq{\rm dom}(q_{1})$;
  3. (3)
    ${\rm dom}(q_{1})\cup w(q_{1},q_{2})(D)$ is the free amalgam of $\mbox{dom}(q_{1})$ and $w(q_{1},q_{2})(D)$ over $\{x_{0}\}$.
For notational simplicity we write $w(q_{1},q_{2})$ as $w$.
We note that $(A\cup B)\cap w(D)=\{x_{0}\}$. This is because $x_{0}\in A\cup B\subseteq C\subseteq D\subseteq\mbox{dom}(q_{1})$ and by (3), $\mbox{dom}(q_{1})\cup w(D)$ is the free amalgam of $\mbox{dom}(q_{1})$ and $w(D)$ over $\{x_{0}\}$.
Let $w_{1}(D)$ be an isomorphic copy of $w(D)$ so that $w_{1}(D)$ and $w(D)$ have exactly one common point $x_{0}$. For each $y\in D$, the point in $w_{1}(D)$ corresponding to $w(y)\in w(D)$ will be denoted by $w_{1}(y)$. Let $D_{1}^{\prime}$ be the free amalgam of $D^{\prime}$ and $w_{1}(D)$ over $\{x_{0}\}$. Figure 1 illustrates the containment relationship among the sets considered so far.
$D^{\prime}$$w(D)$$w_{1}(D)$$D$$C$$A$$B$$\mbox{dom}(q_{1})$$x_{0}$
Figure 1. The composition of $D_{1}^{\prime}$.
Extend the ordering on $D^{\prime}$ to $D_{1}^{\prime}$ so that the following hold:
  1. (a)
    for any $y_{1},y_{2}\in D$, $w_{1}(y_{1})<w_{1}(y_{2})$ iff $w(y_{1})<w(y_{2})$;
  2. (b)
    for any $u\in A\cup B$ and $y\in D$, if $w(y)<b_{n}=\max(A\cup B)$, then $u<w_{1}(y)$ iff $u<w(y)$;
  3. (c)
    for any $y\in D$, if $w(y)>b_{n}$, then $a_{n}<w_{1}(y)<b_{n}$.
Such an extension clearly exists.
Now define $g_{1}\colon A\cup w(D)\to A\cup w_{1}(D)$ by letting $g_{1}(a)=a$ for all $a\in A$ and $g_{1}(w(y))=w_{1}(y)$ for all $y\in D$. Then $g_{1}$ is order-preserving by the above requirements (a) through (c). To see that $g_{1}$ is a partial isometry, let $a\in A$ and $y\in D$. Since $\mbox{dom}(q_{1})\cup w(D)$ is the free amalgam of $\mbox{dom}(q_{1})$ and $w(D)$ over $\{x_{0}\}$ and $D_{1}^{\prime}$ is the free amalgam of $D^{\prime}$ and $w_{1}(D)$ over $\{x_{0}\}$, we have
$d(a,w(y))=d(a,x_{0})+d(x_{0},w(y))=d(a,x_{0})+d(x_{0},w_{1}(y))=d(a,w_{1}(y)).$
Intuitively, $g_{1}$ sends $w(y)$ to a copy which is smaller than $b_{n}$ and fixes $A$ pointwise. Consequently, we have $w_{1}(D)<b_{n}$ and, for any $y\in D$, $w_{1}(y)<x_{0}$ iff $w(y)<x_{0}$.
Let $w_{2}(D)$ be a fresh isomorphic copy of $w(D)$ such that $w_{2}(D)$ and $w(D)$ has exactly one common point $x_{0}$. Let $D_{2}^{\prime}$ be the free amalgam of $D_{1}^{\prime}$ and $w_{2}(D)$ over $\{x_{0}\}$. Extend the ordering on $D_{1}^{\prime}$ to $D_{2}^{\prime}$ so that the following hold:
  1. (d)
    for any $y_{1},y_{2}\in D$, $w_{2}(y_{1})<w_{2}(y_{2})$ iff $w_{1}(y_{1})<w_{1}(y_{2})$;
  2. (e)
    for any $u\in A\cup B$ and $y\in D$, if $w(y)<a_{n}$, then $u<w_{2}(y)$ iff $u<w_{1}(y)$;
  3. (f)
    for any $y\in D$, if $w(y)>a_{n}$, then $b_{n-1}<w_{2}(y)<a_{n}$.
Define an order-preserving partial isometry $g_{2}\colon B\cup w_{1}(D)\to B\cup w_{2}(D)$ by letting $g_{2}(b)=b$ for all $b\in B$ and $g_{2}(w_{1}(y))=w_{2}(y)$ for all $y\in D$. Moreover, $w_{2}(D)<a_{n}$ and, for any $y\in D$, $w_{2}(y)<x_{0}$ iff $w_{1}(y)<x_{0}$.
Repeating this construction, we obtain, for $1\leq i\leq 2n$, $D_{i}^{\prime}=D^{\prime}\cup w_{1}(D)\cup\dots\cup w_{i}(D)$, which can be viewed as the free amalgam of $w_{i}(D)$ and the remaining part over $\{x_{0}\}$, an ordering on $D_{2n}^{\prime}$ extending that on $D^{\prime}$, and order-preserving partial isometries $g_{1},g_{2},g_{3},\dots,g_{2n}$ such that
  1. (i)
    for any $1\leq i\leq n$, $g_{2i-1}(a)=a$ for all $a\in A$, and $g_{2i}(b)=b$ for all $b\in B$;
  2. (ii)
    for any $1\leq i\leq n$ and $y\in D$, $g_{2i}(w_{2i-1}(y))=w_{2i}(y)$, and $g_{2i-1}(w_{2i-2}(y))=w_{2i-1}(y)$ if $i>1$.
Moreover, we have $w_{2n}(D)<a_{1}$ and, for any $y\in D$, $w_{2n}(y)<x_{0}$ iff $w(y)<x_{0}$.
Define a map
$r\colon A\cup w_{2n}(C)\to A\cup w_{2n}(p(C))$
by letting $r(a)=a$ for all $a\in A$ and $r(w_{2n}(c))=w_{2n}(p(c))$ for all $c\in C$. Then for $a\in A$ and $c\in C$, we have
$d(a,w_{2n}(c))$ $=d(a,x_{0})+d(x_{0},w_{2n}(c))=d(a,x_{0})+d(x_{0},w(c))$
$=d(a,x_{0})+d(x_{0},c)=d(a,x_{0})+d(x_{0},p(c))$
$=d(a,x_{0})+d(x_{0},w(p(c)))$
$=d(a,x_{0})+d(x_{0},w_{2n}(p(c)))=d(a,w_{2n}(p(c)))$
$=d(r(a),r(w_{2n}(c))).$
Thus $r$ is a partial isometry. Note that $w$ and $p$ are order-preserving and each fixes $x_{0}$. Thus for any $c\in C$, we have
$w_{2n}(c)<x_{0}$ $\iff w(c)<x_{0}\iff c<x_{0}$
$\iff p(c)<x_{0}\iff w(p(c))<x_{0}$
$\iff w_{2n}(p(c))<x_{0}.$
Since for all $y\in D$, $w_{2n}(y)<a_{1}$, we conclude that $r$ is an order-preserving partial isometry.
By Lemma 5.6 and our construction, the distances in $D_{2n}^{\prime}$ are all taken from the additive semigroup generated by the distances in $D$. Now redefine the metric on $D_{2n}^{\prime}$ by taking the minimum of the original distance value and $\sup\Delta$. Call the resulting metric space $D^{\prime\prime}$. Since $D$ is a $\Delta$-metric space, it follows that $D^{\prime\prime}$ is a $\Delta$-metric space, and, as an ordered metric space, $D^{\prime\prime}$ is an extension of $D$. Moreover, each of the order-preserving partial isometries $q_{1},q_{2},g_{1},g_{2},\dots,g_{2n},r$ continues to be order-preserving partial isometries of $D^{\prime\prime}$.
Now by the weak homogeneity of $\mathbb{U}_{\Delta}^{<}$, we can extend $D$ to an isomorphic copy of $D^{\prime\prime}$ as a substructure of $\mathbb{U}_{\Delta}^{<}$. By the ultrahomogeneity of $\mathbb{U}_{\Delta}^{<}$, we can extend all the above order-preserving partial isometries to full automorphisms of $\mathbb{U}_{\Delta}^{<}$. Denote the resulting automorphisms by the same notation. Then for $1\leq i\leq 2n$, we have $r,q_{1},g_{i}\in{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})$ if $i$ is odd, and $q_{2},g_{i}\in{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})$ if $i$ is even. Finally, for any $c\in C$,
$p(c)=w(q_{1},q_{2})^{-1}g_{1}^{-1}\dots g_{2n}^{-1}rg_{2n}\dots g_{1}w(q_{1},q_{2})(c).$
Thus if we let
$q=w(q_{1},q_{2})^{-1}g_{1}^{-1}\dots g_{2n}^{-1}rg_{2n}\dots g_{1}w(q_{1},q_{2}),$
then $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ and $p\upharpoonright C=q\upharpoonright C$ as required. ∎
Lemma 5.8.
Let $\Delta$ be a countable distance value set, and let $A,B\subseteq\mathbb{U}_{\Delta}^{<}$ be finite subsets. Suppose $A\cap B=\varnothing$. Then
$\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle\ is\ dense\ in\ {\rm Aut}(\mathbb{U}_{\Delta}^{<}).$
Proof.
Let $p\in{\rm Aut}(\mathbb{U}_{\Delta}^{<})$ and let $C\subseteq\mathbb{U}_{\Delta}^{<}$ be a finite subset. We show that there is a $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ such that $p\upharpoonright C=q\upharpoonright C$. Again we may assume $A\cup B\subseteq C$. Consider the ordered $\Delta$-metric space $C\cup p(C)$. Extend it to an ordered $\Delta$-metric space $C\cup p(C)\cup\{x_{0}\}$, where the metric is defined by
$d(x_{0},y)=\mbox{diam}(C\cup p(C))=\max\{d(y_{1},y_{2})\colon y_{1},y_{2}\in C\cup p(C)\}$
for any $y\in C\cup p(C)$, and the ordering is defined by $x_{0}<y$ for all $y\in C\cup p(C)$. Such an extension can be found in $\mathbb{U}_{\Delta}^{<}$ by the weak homogeneity of $\mathbb{U}^{<}_{\Delta}$, and so we may assume $x_{0}\in\mathbb{U}_{\Delta}^{<}$. Now consider $A^{\prime}=A\cup\{x_{0}\}$, $B^{\prime}=B\cup\{x_{0}\}$ and let $p^{\prime}\in{\rm Aut}_{x_{0}}(\mathbb{U}_{\Delta}^{<})$ be an extention of $p\upharpoonright C\cup\{(x_{0},x_{0})\}$. Applying Lemma 5.7 to $A^{\prime},B^{\prime}$ and $p^{\prime}$, we get a $q\in\langle{\rm Aut}_{A^{\prime}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B^{\prime}}(\mathbb{U}_{\Delta}^{<})\rangle$ such that $q\upharpoonright C\cup\{x_{0}\}=p^{\prime}\upharpoonright C\cup\{x_{0}\}$. Then in particular $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ and $q\upharpoonright C=p\upharpoonright C$ as required. ∎
We also need the following two lemmas about proper open subgroups of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$.
Lemma 5.9.
Let $V$ be a proper open subgroup of ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. Then there exists an element $x\in\mathbb{U}_{\Delta}^{<}$ such that $V\cdot x$ is finite.
Proof.
Toward a contradiction, assume $V\cdot x$ is infinite for all $x\in X$. We prove that $V$ is dense. Since $V$ is also closed, it follows that $V={\rm Aut}(\mathbb{U}_{\Delta}^{<})$, contradicting the properness of $V$.
Note that a basic open neighborhood of the identity element of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ has the form ${\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})$, where $A$ is a finite subset of $\mathbb{U}^{<}_{\Delta}$. Since $V$ is an open subgroup, there is a finite $A\subseteq\mathbb{U}^{<}_{\Delta}$ such that ${\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})\subseteq V$. By Neumann’s lemma (see Corollary 4.2.2 of Hodges [13]), there exists $g\in V$ such that $g(A)\cap A=\varnothing$.
By Lemma 5.8, $\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{g(A)}(\mathbb{U}_{\Delta}^{<})\rangle$ is dense in ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. This group is contained in $V$, and thus $V$ is dense. ∎
Lemma 5.10.
Let $V$ be a proper open subgroup of ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. The following are equivalent:
  1. (1)
    There exists a unique $x\in\mathbb{U}_{\Delta}^{<}$ such that $V={\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})$.
  2. (2)
    $V$ is maximal among proper closed subgroups of ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$.
Proof.
First suppose $V={\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})$. Assume $H$ is a closed subgroup of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ such that $V\subsetneq H\subsetneq\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$. Let $g\in H\setminus V$. Then $g(x)\neq x$ and ${\rm Aut}_{\{g(x)\}}(\mathbb{U}_{\Delta}^{<})=g{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})g^{-1}=gVg^{-1}\leq H$. By Lemma 5.8, $\langle{\rm Aut}_{\{g(x)\}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})\rangle$ is dense in ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. Since it is a subgroup of $H$ and $H$ is closed, we have $H={\rm Aut}(\mathbb{U}_{\Delta}^{<})$, contradicting the properness of $H$.
For the converse, assume (2) holds. By Lemma 5.9, there exists an element $x\in\mathbb{U}_{\Delta}^{<}$ such that $V\cdot x$ is finite. We show that $V\subseteq\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})$. Otherwise, let $g\in V$ be such that $g(x)\neq x$. Then either $g(x)>x$ or $g(x)<x$. Since $g$ is order-preserving, we get that either $g^{n+1}(x)>g^{n}(x)$ for all natural numbers $n$, or $g^{n+1}(x)<g^{n}(x)$ for all natural numbers $n$. In either case, $V\cdot x$ is infinite.
It is clear that $\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})\neq\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$. Thus by the the maximality of $V$, we have $V=\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})$.
For the uniqueness, assume ${\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})={\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<})$ but $x\neq y$. Then by Lemma 5.8, $\langle{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<})\rangle$ is dense. But
$\langle{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<})\rangle={\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})={\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<}),$
and thus they are both the whole group, which is impossible. ∎
We are now ready to finish the proof of Theorem 5.4. Recall that the remaining direction is (i) implying (ii).
Let $f\colon\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})\to\mbox{\rm Aut}(\mathbb{U}^{<}_{\Lambda})$ be a topological group isomorphism. By Lemma 5.10, for each $x\in\mathbb{U}_{\Delta}^{<}$, there is a unique $y\in\mathbb{U}_{\Lambda}^{<}$ such that $f({\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<}))={\rm Aut}_{\{y\}}(\mathbb{U}_{\Lambda}^{<})$. We write $y=g(x)$. The map $g$ is well-defined and bijective.
Note that for any $x,y,x^{\prime},y^{\prime}\in\mathbb{U}_{\Delta}^{<}$, ${\rm Aut}_{\{x,y\}}(\mathbb{U}_{\Delta}^{<})$ and ${\rm Aut}_{\{x^{\prime},y^{\prime}\}}(\mathbb{U}_{\Delta}^{<})$ are conjugate in $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ iff $d(x,y)=d(x^{\prime},y^{\prime})$. Also note that
${\rm Aut}_{\{x,y\}}(\mathbb{U}_{\Delta}^{<})=\mbox{\rm Aut}_{\{x\}}(\mathbb{U}^{<}_{\Delta})\cap\mbox{\rm Aut}_{\{y\}}(\mathbb{U}^{<}_{\Delta})$
and similarly for $\mbox{\rm Aut}_{\{x^{\prime},y^{\prime}\}}(\mathbb{U}^{<}_{\Delta})$. Thus by the definition of $g$, we have that for any $x,y,x^{\prime},y^{\prime}\in\mathbb{U}_{\Delta}^{<}$,
$d(x,y)=d(x^{\prime},y^{\prime})\iff d(g(x),g(y))=d(g(x^{\prime}),g(y^{\prime})).$
Now for any $d\in\Delta$, we define $\theta(d)\in\Lambda$ by finding $x,y\in\mathbb{U}_{\Delta}^{<}$ such that $d(x,y)=d$, and then setting $\theta(d)=d(g(x),g(y))$. By the above observation, $\theta$ is well-defined and bijective. We show that $\theta$ witnesses the equivalence of $\Delta$ and $\Lambda$. Suppose $(d_{1},d_{2},d_{3})$ is a $\Delta$-triangle. There exist $x,y,z\in\mathbb{U}_{\Delta}^{<}$ such that the distances among them are exactly $(d_{1},d_{2},d_{3})$. Then the distaces among $g(x),g(y),g(z)$ are $(\theta(d_{1}),\theta(d_{2}),\theta(d_{3}))$. Thus $(\theta(d_{1}),\theta(d_{2}),\theta(d_{3}))$ is a $\Lambda$-triangle. The converse also holds.
The proof of Theorem 5.4 is complete.
6. Some Descriptive Set Theoretic Results
Since $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ are Polish groups that are (isomorphic to) closed subgroups of $S_{\infty}$, our results in this paper fit very well in the following research program promoted by Kechris–Nies–Tent [14].
Program.
  1. (1)
    For natural classes of closed subgroups of $S_{\infty}$, determine if they are Borel as subsets of $\mathcal{F}(S_{\infty})$.
  2. (2)
    If a class is Borel, study the relative complexity of the topological isomorphism relation, using Borel reducibility $\leq_{B}$.
Here $\mathcal{F}(S_{\infty})$ is the standard Borel space consisting of all closed subsets of $S_{\infty}$. The $\sigma$-algebra of Borel sets on $\mathcal{F}(S_{\infty})$ is generated (via iterations of countable union, countable intersection, and complementation) by sets of the form
$\{F\in\mathcal{F}(S_{\infty})\colon F\cap U\neq\emptyset\},$
where $U\subseteq S_{\infty}$ is a basic open set, i.e., of the form
$N_{\bar{a},\bar{b}}=\{g\in S_{\infty}\colon g(\bar{a})=\bar{b}\}$
for some tuples $\bar{a},\bar{b}$ of natural numbers of the same length. For other undefined terms and concepts, see [7].
We now focus on the class of all closed subgroups of $S_{\infty}$ that are extremely amenable. We denote this class by $\mathsf{EA}$. The following theorem shows that $\mathsf{EA}$ is a Borel subset of $\mathcal{F}(S_{\infty})$.
Theorem 6.1.
The class $\mathsf{EA}$ of extremely amenable closed subgroups of $S_{\infty}$ is Borel.
Proof.
Given any closed subgroup $G$ of $S_{\infty}$, by a well-known construction (see Theorem 4.1.4 of [13] and Section 2 of [15]), we define a canonical structure $\mathcal{A}_{G}$ so that $\mbox{\rm Aut}(\mathcal{A}_{G})=G$. To do this, consider the diagonal action of $G$ on $\mathbb{N}^{n}$ for all postive integers $n\geq 1$, which is defined by
$g\cdot(a_{1},\dots,a_{n})=(g(a_{1}),\dots,g(a_{n})).$
For each $n\geq 1$, let $\mathcal{O}_{1}^{n},\mathcal{O}_{2}^{n},\dots$ enumerate the set
$\left\{G\cdot(a_{1},\dots,a_{n})\colon(a_{1},\dots,a_{n})\in\mathbb{N}^{n}\right\},$
which is the set of all orbits of $G$ for the diagonal action on $\mathbb{N}^{n}$. Consider the language $\mathcal{L}$ consisting of relation symbols $R_{n,i}$ for $n,i\geq 1$. Each $R_{n,i}$ is an $n$-ary relation symbol. The structure $\mathcal{A}_{G}$ is an $\mathcal{L}$-structure whose universe is $\mathbb{N}$, with $R_{n,i}^{\mathcal{A}_{G}}=\mathcal{O}_{i}^{n}$. Then it is easy to check that $\mbox{\rm Aut}(\mathcal{A}_{G})=G$. It is also straightforward to see that $\mathcal{A}_{G}$ is a Fraïssé structure, i.e., it is locally finite, countably infinite and ultrahomogeneous.
We note that the collection of all $\mathcal{L}$-structures with the universe $\mathbb{N}$ forms a Polish space $X_{\mathcal{L}}=\prod_{n,i\geq 1}X_{n,i}$, where each $X_{n,i}$ is a copy of the Cantor space $2^{\mathbb{N}^{n}}$. The map $G\mapsto\mathcal{A}_{G}$ from $\mathcal{F}(S_{\infty})$ to $X_{\mathcal{L}}$ is Borel. In fact, a subbasic open set in $X_{\mathcal{L}}$ is determined by a tuple $\bar{a}=(a_{1},\dots,a_{n})$ being assigned value 00 or value $1$. If $U$ is a basic open set in $X_{\mathcal{L}}$, then whether $\mathcal{A}_{G}\in U$ is determined by finitely many requirements of form $\exists g\in G\ g(\bar{a})=\bar{b}$ and its negation, where $\bar{a},\bar{b}$ are tuples of the same length. Thus $\{G\in\mathcal{F}(S_{\infty})\colon\mathcal{A}_{G}\in U\}$ is clearly a Borel set.
Let $\mathcal{K}$ be the age of $\mathcal{A}_{G}$. Then $\mathcal{A}_{G}=\mbox{Flim}\,\mathcal{K}$. By Theorem 2.2, $G=\mbox{\rm Aut}(\mathcal{A}_{G})=\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$ is extremely amenable iff $\mathcal{K}$ has the Ramsey property and every element of $\mathcal{K}$ is rigid. Thus to prove $\mathsf{EA}$ is Borel, it suffices to show both the following:
  1. (i)
    the collection of $\mathcal{L}$-structures $\mathcal{A}\in X_{\mathcal{L}}$ where the age of $\mathcal{A}$ satisfies the Ramsey property is a Borel subset of $X_{\mathcal{L}}$;
  2. (ii)
    the collection of $\mathcal{L}$-structures $\mathcal{A}\in X_{\mathcal{L}}$ where every finite substructure is rigid is a Borel subset of $X_{\mathcal{L}}$.
Both (i) and (ii) refer to finite subsets of $\mathcal{A}$ (whose universe is $\mathbb{N}$). Thus, despite the tediousness of the statements of the Ramsey property and the rigidity, they can be expressed as Borel statements in the topology of $X_{\mathcal{L}}$. This completes the proof. ∎
Thus we have accomplished the first goal of the above-mentioned research program for the class of extremely amenable closed subgroups of $S_{\infty}$. Next we address the second goal and state some consequences of our results from the previous sections. We let $\cong_{\mathsf{EA}}$ denote the topological group isomorphism relation among groups in $\mathsf{EA}$. As noted in [14], it follows from Theorem 6.1 that $\cong_{\mathsf{EA}}$ is an analytic equivalence relation. This gives an upper bound in the Borel reducibility hierarchy for $\cong_{\mathsf{EA}}$. In the rest of this section we focus on lowers bounds for $\cong_{\mathsf{EA}}$.
Definition 6.1.
Let $\mathsf{CDVS}\subseteq\mathbb{R}^{\mathbb{N}}$ consist of all sequences $(d_{i})_{i\in\mathbb{N}}$ of real numbers such that
  1. (a)
    $d_{i}\geq 0$ for all $i$; the set $\{i\in\mathbb{N}\colon d_{i}=0\}$ is infinite;
  2. (b)
    if $d_{i},d_{j}>0$, then $d_{i}=d_{j}$ iff $i=j$;
  3. (c)
    if $\sup_{k}d_{k}<\infty$, then $\sup_{k}d_{k}=d_{i}$ for some $i$;
  4. (d)
    for any $i,j\geq 0$, if $d_{i}+d_{j}<\sup_{k}d_{k}$ then there is $\ell$ such that $d_{\ell}=d_{i}+d_{j}$.
Let $\mathsf{CDVS}_{0}$ consist of all sequences $(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}$ such that
$\inf\{d_{i}>0\colon i\in\mathbb{N}\}=0.$
Intuitively, $\mathsf{CDVS}$ consists of codes of all countable distance value sets, which are enumerated nonrepeatedly as sequences of real numbers. We include zero entries in the codes to accommodate finite distance value sets. Similarly, $\mathsf{CDVS}_{0}$ consists of codes of all countable distance value sets $\Delta$ with $\inf\Delta=0$. Both $\mathsf{CDVS}$ and $\mathsf{CDVS}_{0}$ are Borel subsets of $\mathbb{R}^{\mathbb{N}}$, and hence they are both standard Borel spaces. Recall the following definition of Borel reducibility among equivalence relations on standard Borel spaces.
Definition 6.2.
Let $E,F$ be equivalence relations on standard Borel spaces $X,Y$, respectively. We say that $E$ is Borel reducible to $F$, and denote it as $E\leq_{B}F$, if there is a Borel function $f\colon X\to Y$ such that for all $x,y\in X$,
$xEy\iff f(x)Ff(y).$
We say that $E$ and $F$ are Borel bireducible if both $E\leq_{B}F$ and $F\leq_{B}E$. We say that $E$ is strictly Borel reducible to $F$, and write $E<_{B}F$, if $E\leq_{B}F$ and $F\not\leq_{B}E$.
Next we define the equivalence relations of interest which are lower bounds for $\cong_{\mathsf{EA}}$ in the sense of $\leq_{B}$.
Definition 6.3.
Let $\sim$ be the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined by
$(c_{i})_{i}\sim(d_{i})_{i}\iff\exists g\in S_{\infty}\ \exists r\in(0,+\infty)\ \forall i\ d_{g(i)}=r\cdot c_{i}.$
Let $\approx$ be the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined by
$\begin{array}[]{rcl}(c_{i})_{i}\approx(d_{i})_{i}&\iff&\exists g\in S_{\infty}\ [\forall i\ (c_{i}=0\longleftrightarrow d_{g(i)}=0)\mbox{ and }\\ &&\forall i,j,k\ (|c_{j}-c_{k}|\leq c_{i}\leq c_{j}+c_{k}\longleftrightarrow\\ &&|d_{g(j)}-d_{g(k)}|\leq d_{g(i)}\leq d_{g(j)}+d_{g(k)})].\end{array}$
For any $A\subseteq\mathbb{R}^{\mathbb{N}}$, let $\sim_{A}$ denote $\sim\upharpoonright(A\times A)$, and similarly, let $\approx_{A}$ denote $\approx\upharpoonright(A\times A)$.
Thus $\approx_{\mathsf{CDVS}}$ is the equivalence between countable distance value sets and $\approx_{\mathsf{CDVS}_{0}}$ is the equivalence between countable distance values sets without a positive lower bound. Theorem 4.5 can be restated as $\sim_{\mathsf{CDVS}_{0}}\,=\,\approx_{\mathsf{CDVS}_{0}}$.
The following theorem is a summary of our results about these equivalence relations.
Theorem 6.2.
$E_{GL_{2}(\mathbb{Q})}^{IR}\,\leq_{B}\ \sim_{\mathsf{CDVS}_{0}}\ \leq_{B}\ \approx_{\mathsf{CDVS}}\ \leq_{B}\ \cong_{\mathsf{EA}}$.
Proof.
$E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}\,\sim_{\mathsf{CDVS}_{0}}$ is witnessed by the map $\alpha\mapsto\Delta_{\alpha}$ and guaranteed by Theorem 4.6. We only need to see that there is a canonical Borel function that gives a coding sequence for $\Delta_{\alpha}=\{p\alpha+q\colon p,q\in\mathbb{Q}\}\cap(0,+\infty)$. One way to do this is as follows. Set the odd part of the sequence to zero. For the even part of the sequence, go through a canonical enumeration of all pairs $(p,q)$ where $p,q\in\mathbb{Q}$, and include $p\alpha+q$ only when it is a positive number.
$\sim_{\mathsf{CDVS}_{0}}\,\leq_{B}\,\approx_{\mathsf{CDVS}}$ is clear in view of Theorem 4.5, which states $\sim_{\mathsf{CDVS}_{0}}\,=\,\approx_{\mathsf{CDVS}_{0}}$, and the obvious $\approx_{\mathsf{CDVS}_{0}}\,\leq_{B}\,\approx_{\mathsf{CDVS}}$.
The last reduction $\approx_{\mathsf{CDVS}}\,\leq_{B}\,\cong_{\mathsf{EA}}$ is witnessed by the map $\Delta\mapsto\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ and guaranteed by Theorem 5.4. Again we only need to show that the map from a distance value set $\Delta$ to ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ is Borel. This is done in two steps. In the first step, we describe a Borel way to enumerate the class $\mathcal{K}^{<}_{\Delta}$, the class of all finite ordered $\Delta$-metric spaces, from a given sequence $(d_{i})_{i}$ coding $\Delta$. For this, we fix a countable language $\mathcal{L}$ consisting of countably many binary relation symbols, i.e., =$\mathcal{L}=\{R_{i}\}_{i\in\mathbb{N}}\cup\{<\}$. Enumerate those finite $\mathcal{L}$-structures $\mathcal{A}$ such that
  1. (i)
    for any distinct $x,y\in|\mathcal{A}|$, there is exactly one $i$ such that $d_{i}>0$ and $R^{\mathcal{A}}_{i}(x,y)$ holds;
  2. (ii)
    for any $x,y,z\in|\mathcal{A}|$ and $i,j,k\in\mathbb{N}$, if $R^{\mathcal{A}}_{i}(x,y)$, $R^{\mathcal{A}}_{j}(y,z)$ and $R^{\mathcal{A}}_{k}(x,z)$, then $|d_{j}-d_{k}|\leq d_{i}\leq d_{j}+d_{k}$;
  3. (iii)
    $<^{\mathcal{A}}$ is a linear order.
It is clear that such structures correspond to finite ordered $\Delta$-metric spaces.
Now, given an enumeration of $\mathcal{K}^{<}_{\Delta}$, one can construct the Fraïssé limit in a Borel way. This follows from the existence proof of Fraïssé’s theorem (see, e.g. Lemma 7.1.6 of Hodges [13]). Since the Fraïssé structure is ultrahomogeneous, its automorphism group intersects a basic open set $N_{\bar{a},\bar{b}}$ iff the substructures $\bar{a}$ and $\bar{b}$ are isomorphic. Thus the map from the Fraïssé structures to their automorphism groups is Borel. ∎
Finally, we compare the above equivalence relations with some benchmark equivalence relations in the theory of Borel reducibility for equivalence relations.
We list all relevant benchmark equivalence relations in the following. Discussions about the ones without references can be found in [7].
  1. (1)
    $=$ is the equality relation on $2^{\mathbb{N}}$.
  2. (2)
    $E_{0}$ is the eventual agreement relation on $2^{\mathbb{N}}$ defined as
    $(x_{i})_{i}E_{0}(y_{i})_{i}\iff\exists i\ \forall j\geq i\ x_{j}=y_{j}.$
  3. (3)
    $E_{\infty}$ is the universal countable Borel equivalence relation. It can be defined as the following equivalence relation on $2^{\mathbb{F}_{2}}$, where $\mathbb{F}_{2}$ is the free group with two generators:
    $(x_{g})_{g}E_{\infty}(y_{g})_{g}\iff\exists h\in\mathbb{F}_{2}\ \forall g\in\mathbb{F}_{2}\ x_{g}=y_{gh}.$
  4. (4)
    $=^{+}$ is the equivalence relation on $(2^{\mathbb{N}})^{\mathbb{N}}$ defined as
    $(x_{i})_{i}\ =^{+}(y_{i})_{i}\iff\{x_{i}\colon i\in\mathbb{N}\}=\{y_{i}\colon i\in\mathbb{N}\}.$
  5. (5)
    $=^{++}$ is the equivalence relation on $((2^{\mathbb{N}})^{\mathbb{N}})^{\mathbb{N}}$ defined as
    $(x_{i})_{i}\ =^{++}(y_{i})_{i}\iff\forall i\ \exists j\ x_{i}=^{+}y_{j}\mbox{ and }\forall j\ \exists i\ x_{i}=^{+}y_{j}.$
  6. (6)
    Let $\mathcal{P}$ be the powerset operation on sets. Let TC denote the transitive closure operation on sets. For $k=0,1$, let $\mathcal{Q}^{*}_{k}(\mathbb{N})$ be the collection of pairs $(A,R)$ such that:
    1. a)
      $A$ is a hereditarily countable set in $\mathcal{P}^{3}(\mathbb{N})$;
    2. b)
      $R$ is a ternary relation on $A\times A\times(\mathcal{P}^{k}(\mathbb{N})\cap\mbox{\rm TC}(A))$ such that
      • for any $a,b\in A$ there is some $r$ such that $R(a,b,r)$ holds;
      • given any $a\in A$, for any $b,b^{\prime}\in A$ and any $r$, if $R(a,b,r)$ and $R(a,b^{\prime},r)$ both hold then $b=b^{\prime}$.
    The equivalence relation $\cong^{*}_{3,k}$ is defined as the isomorphism relation of countable structures coding pairs $(A,R)$ in $Q^{*}_{k}(\mathbb{N})$. These equivalence relations were defined by Hjorth–Kechris–Louveau [12]. Our presentation follows [2].
The following Borel reducibility results are well-known (see [7]):
$=\ <_{B}E_{0}\,<_{B}E_{\infty}<_{B}\ =^{+}\,<_{B}\ =^{++}.$
It was also shown in Hjorth–Kechris–Louveau [12] that
$=^{+}\,<_{B}\ \cong^{*}_{3,0}\ <_{B}\ \cong^{*}_{3,1}\ <_{B}\ =^{++}.$
Corollary 5.5 can be restated as the equivalence relation $=$ being a lower bound for $\cong_{\mathsf{EA}}$ in the Borel reducibility hierarchy.
We note the following fact.
Lemma 6.3.
$E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$.
Proof.
$E_{GL_{2}(\mathbb{Q})}^{IR}$ is induced by a continuous action of a countable group on a standard Borel space. Thus $E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$. Note that every orbit is dense for this action. It follows from a theorem of Becker–Kechris (see Theorem 6.2.1 of [7]) that $E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}$. ∎
This gives $E_{0}$ as a lower bound for $\cong_{\mathsf{EA}}$, which is an improvement of Corollary 5.5.
Next we turn to equivalence relations $\sim_{\mathsf{CDVS}_{0}}$. Here we consider another equivalence relation defined and studied by Calderoni–Marker–Motto Ros–Shani [2]. Let $\mathsf{ArGp}$ be the space of all countable Archimedean ordered groups, and let $\cong_{\mathsf{ArGp}}$ denote the isomorphism relation on $\mathsf{ArGp}$. It was proved in [2] that
$=^{+}\ <_{B}\ \cong_{\mathsf{ArGp}}\ \leq_{B}\ \cong^{*}_{3,1}.$
Here we prove the following facts.
Lemma 6.4.
$\cong_{\mathsf{ArGp}}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$.
Proof.
By a theorem of Hölder (see Theorem 2.1 of [2]), every element of $\mathsf{ArGp}$ is isomorphic to a subgroup of $\mathbb{R}$ with the natural order of real numbers. Thus we may regard $\mathsf{ArGp}$ as the space of all countable subgroups of $\mathbb{R}$ with the natural order. Note that if $G\in\mathsf{ArGp}$ and $\inf[G\cap(0,+\infty)]\neq 0$, then $G$ is isomorphic to $\mathbb{Z}$. Let $\mathsf{ArGp}_{0}$ be the space of all elements of $\mathsf{ArGp}$ which are not isomorphic to $\mathbb{Z}$. Then $\mathsf{ArGp}\setminus\mathsf{ArGp}_{0}$ consists of exactly one isomorphism class of $\mathsf{ArGp}$, and it follows easily that $\cong_{\mathsf{ArGp}}\ \leq_{B}\ \cong_{\mathsf{ArGp}_{0}}$.
For each $G\in\mathsf{ArGp}_{0}$, let $\Delta(G)=G\cap(0,+\infty)$. Then $\Delta(G)$ is a countable distance value set and $\inf\Delta(G)=0$. It is easy to see that $\Delta\colon\mathsf{ArGp}_{0}\to\mathsf{CDVS}_{0}$ is a Borel map.
To see that $\Delta$ is a Borel reduction witnessing $\cong_{\mathsf{ArGp}}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$, we use a result of Hion (see Lemma 2.2 of [2]) which states that for any $G,H\in\mathsf{ArGp}$, $G\cong H$ iff there is a real number $\alpha$ such that $\alpha\cdot G=H$. It is then clear that $G\cong H$ iff there is a positive real number such that $\alpha\cdot G=H$ iff there is a positive real number such that $\alpha\cdot\Delta(G)=\Delta(H)$ iff $\Delta(G)\sim\Delta(H)$. ∎
Combining with the results of [2], we obtain $=^{+}$ as a non-optimal lower bound for $\sim_{\mathsf{CDVS}_{0}}$, as well as for $\cong_{\mathsf{EA}}$, in the Borel reducibility hierarchy. In symbols, we have $=^{+}\ <_{B}\ \cong_{\mathsf{EA}}$. This is a further improvement of Corollary 5.5 and Lemma 6.3. In particular, we have shown Theorem 1.7.
The next result will give an upper bound for $\sim_{\mathsf{CDVS}_{0}}$, and in particular implies
$=^{+}\ <_{B}\ \sim_{\mathsf{CDVS}_{0}}\ <_{B}\ =^{++}.$
Proposition 6.5.
$\sim_{\mathsf{CDVS}_{0}}\ \leq_{B}\ \cong^{*}_{3,1}$.
Proof.
Similar to Proposition 3.12 of [2], using Corollary 6.4 of [12], this can be proved by showing that $\sim_{\mathsf{CDVS}_{0}}$ is Borel bireducible with an isomorphism relation and it is itself ${\bf\Sigma}^{0}_{4}$.
We first show that $\sim$ is ${\bf\Sigma}^{0}_{4}$. For this, only note that for any $(c_{i})_{i},(d_{i})_{i}\in\mathbb{R}^{\mathbb{N}}$,
$\begin{array}[]{rcl}(c_{i})_{i}\sim(d_{i})_{i}&\iff&\exists i,j\ \left(c_{i}/d_{j}>0\mbox{ and }\right.\\ &&\forall n\ \exists m\ \left(c_{n}=d_{m}c_{i}/d_{j}\right)\mbox{ and }\\ &&\left.\forall m\ \exists n\ \left(c_{n}=d_{m}c_{i}/d_{j}\right)\right).\end{array}$
Next we show that $\sim_{\mathsf{CDVS}_{0}}$ is Borel bireducible with an isomorphism relation. For this we define a language $\mathcal{L}$ and an $\mathcal{L}_{\omega_{1}\omega}$-theory $T$ so that the isomorphism relation on all countable models of $T$ is Borel bireducible with $\sim_{\mathsf{CDVS}_{0}}$.
Let
=$\mathcal{L}=\{0,c,\leq,+\}\cup\{R_{q}\colon q\in\mathbb{Q}^{+}\}$
where
  • $0,c$ are constant symbols;
  • $\leq$ is a binary relation symbol;
  • $+$ is a binary function symbol;
  • for each positive rational number $q\in\mathbb{Q}^{+}$, $R_{q}$ is a binary relation symbol.
In a countable distance value set $\Delta$ with $\inf\Delta=0$, we interpret $c$ to be 00 if $\sup\Delta=+\infty$ and to be $\sup\Delta$ otherwise. For each $q\in\mathbb{Q}^{+}$, $R_{q}$ is interpreted as $R_{q}(x,y)$ iff $q<x/y$. Under these interpretations, $\Delta\cup\{0\}$ becomes a countable $\mathcal{L}$-structure.
Let $T$ be an $\mathcal{L}_{\omega_{1}\omega}$-theory describing the following properties of a countable $\mathcal{L}$-structure $\mathcal{M}$ (for notational simplicity, we omit the superscripts $\mathcal{M}$ in the interpretations of the symbols of $\mathcal{L}$ in $\mathcal{M}$):
  1. (1)
    For all $q\in\mathbb{Q}^{+}$ and $x\in|\mathcal{M}|$, $\lnot R_{q}(x,0)$ and $\lnot R_{q}(0,x)$.
  2. (2)
    For all $x,y\in|\mathcal{M}|\setminus\{0\}$, the set of $q\in\mathbb{Q}^{+}$ with $R_{q}(x,y)$ is nonempty and forms a Dedekind cut of positive rational numbers, i.e., it is not the whole positive rational numbers, is closed downward and has no maximum. Moreover, the real number defined by the Dedekind cut $\{q\in\mathbb{Q}^{+}\colon R_{q}(x,x)\}$ is $1$.
  3. (3)
    For all $x,x^{\prime},y\in|\mathcal{M}|\setminus\{0\}$, if $x\neq x^{\prime}$, then the Dedekind cuts defined by $x,y$ and $x^{\prime},y$ are different.
  4. (4)
    For all $x,y,z\in|\mathcal{M}|\setminus\{0\}$ and $p,q\in\mathbb{Q}^{+}$, if $R_{p}(x,y)$ and $R_{q}(y,z)$, then $R_{pq}(x,z)$, and if $\lnot R_{p}(x,y)$ and $\lnot R_{q}(y,z)$, then $\lnot R_{pq}(x,z)$.
  5. (5)
    $\leq$ is a linear ordering of $|\mathcal{M}|$, 00 is the $\leq$-least element, and for all $x,x^{\prime}\in|\mathcal{M}|\setminus\{0\}$, $x\leq x^{\prime}$ iff $x=x^{\prime}$ or $R_{1}(x^{\prime},x)$.
  6. (6)
    For all $x,x^{\prime},y\in|\mathcal{M}|\setminus\{0\}$ and $q\in\mathbb{Q}^{+}$, $R_{q}(x+x^{\prime},y)$ iff there are $q_{1},q_{2}\in\mathbb{Q}^{+}$ with $q_{1}+q_{2}=q$ such that $R_{q_{1}}(x,y)$ and $R_{q_{2}}(x^{\prime},y)$.
  7. (7)
    For all $x\in|\mathcal{M}|\setminus\{0\}$ and $q\in\mathbb{Q}^{+}$, there is $x^{\prime}\in|\mathcal{M}|\setminus\{0\}$ such that $\lnot R_{q}(x^{\prime},x)$.
Let $\mbox{Mod}(T)$ denote the Borel class of all countable models of $T$. Let $\cong_{T}$ denote the isomorphism relation on $\mbox{Mod}(T)$. It remains to verify that both $\sim_{\mathsf{CDVS}_{0}}\ \leq_{B}\ \cong_{T}$ and $\cong_{T}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$ hold.
Let $\Delta$ be a countable distance value set with $\inf\Delta=0$. We define a countable $\mathcal{L}$-structure $\mathcal{M}_{\Delta}$. Let $\mathbb{R}^{+}$ denote the semigroup of positive real numbers. Let $\bar{\Delta}\subseteq\mathbb{R}^{+}$ be the sub-semigroup of $\mathbb{R}^{+}$ generated by $\Delta$. Then either $\Delta=\bar{\Delta}$ or $\Delta=\bar{\Delta}\cap(0,{\rm sup}\Delta]$. Let $|\mathcal{M}_{\Delta}|=\{0\}\cup\bar{\Delta}$. Define $0,\leq,+$ in the obvious way. For $q\in\mathbb{Q}^{+}$ and $x,y\in|\mathcal{M}_{\Delta}|$, if either $x=0$ or $y=0$, then $\lnot R_{q}(x,y)$; otherwise, define $R_{q}(x,y)$ iff $q<x/y$. Finally, if $\sup\Delta=+\infty$, set $c=0$; otherwise, set $c={\rm sup}\Delta$. Now it is easy to verify that $\mathcal{M}_{\Delta}$ is a model of $T$. Note that clause (7) holds since $\Delta$ is dense in $(0,\sup\Delta)$. It is also clear that the map $\Delta\mapsto\mathcal{M}_{\Delta}$ gives a Borel reduction from $\sim_{\mathsf{CDVS}_{0}}$ to $\cong_{T}$.
Now given a countable model $\mathcal{M}$ of $T$, we define a map $f:\mathcal{M}\to[0,+\infty)$ as follows. We first choose a non-zero element $y_{0}\in\mathcal{M}$ canonically. Then define $f(0)=0$, and for $x\neq 0$, let $f(x)$ be the positive real number defined by the Dedekind cut $\{q\in\mathbb{Q}^{+}\colon R_{q}(x,y_{0})\}$. By clauses (2) and (3), $f$ is an injection. Let $\Delta_{\mathcal{M}}$ be $f(\mathcal{M})\setminus\{0\}$ if $c=0$ and $f(\mathcal{M})\cap(0,f(c)]$ otherwise.
By clause (7), we have $\inf\Delta_{\mathcal{M}}=0$. Thus to verify that $\Delta_{\mathcal{M}}$ is a countable distance value set, it suffices to check that $f(\mathcal{M})$ is closed under summation. For this just note that if $x,y\in\mathcal{M}$ are nonzero, then
$q<f(x+y)$ $\iff R_{q}(x+y,x^{\prime})$
$\iff\exists q_{1},q_{2}\ q_{1}+q_{2}=q\land R_{q_{1}}(x,x^{\prime})\land R_{q_{2}}(y,x^{\prime})$
$\iff\exists q_{1},q_{2}\ q_{1}+q_{2}=q\land q_{1}<f(x)\land q_{2}<f(y)$
$\iff q<f(x)+f(y).$
Thus we have $f(x+y)=f(x)+f(y)$, and so $f(\mathcal{M})$ is closed under summation.
Finally, we verify that the map $\mathcal{M}\mapsto\Delta_{\mathcal{M}}$ is a Borel reduction from $\cong_{T}$ to $\sim_{\mathsf{CDVS}_{0}}$. It is clear that the map is Borel. For the reduction, we first claim that for nonzero $x,y\in\mathcal{M}$, $R_{q}(x,y)$ iff $q<f(x)/f(y)$. Suppose $R_{q}(x,y)$. There is $q^{\prime}>q$ with $R_{q^{\prime}}(x,y)$. Let $p<q^{\prime}f(y)$. Then $p/q^{\prime}<f(y)$ and thus $R_{p/q^{\prime}}(y,y_{0})$. Since $R_{q^{\prime}}(x,y)$, by clause (4), we have $R_{p}(x,y_{0})$ and thus $p<f(x)$. This shows $q^{\prime}\leq f(x)/f(y)$ and thus $q<f(x)/f(y)$. Similarly we can prove that if $\lnot R_{q}(x,y)$, then $q\geq f(x)/f(y)$. This finishes the proof of the claim.
Now, given countable models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ of $T$, we denote the functions in the definitions of $\Delta_{\mathcal{M}_{1}}$ and $\Delta_{\mathcal{M}_{2}}$ by $f_{1}$ and $f_{2}$, respectively. Suppose $g$ is an isomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$. Let $\tilde{g}=f_{2}gf_{1}^{-1}$. Then for any $x,y\in f_{1}(\mathcal{M}_{1})$ and any $q\in\mathbb{Q}^{+}$,
$\begin{array}[]{rcl}q<x/y&\iff&R_{q}(f_{1}^{-1}(x),f_{1}^{-1}(y))\\ &\iff&R_{q}(gf_{1}^{-1}(x),gf_{1}^{-1}(y))\iff q<\tilde{g}(x)/\tilde{g}(y).\end{array}$
This means $\tilde{g}(x)/\tilde{g}(y)=x/y$ and thus there is a positive real number $\alpha$ such that $\tilde{g}(x)=\alpha x$ for all $x\in f_{1}(\mathcal{M}_{1})$. Note that $f_{2}(c^{\mathcal{M}_{2}})=\tilde{g}(f_{1}(c^{\mathcal{M}_{1}}))=\alpha f_{1}(c^{\mathcal{M}_{1}})$. Thus either
$\Delta_{\mathcal{M}_{2}}=f_{2}(\mathcal{M}_{2})=\alpha\cdot f_{1}(\mathcal{M}_{1})=\alpha\cdot\Delta_{\mathcal{M}_{1}}$
or
$\begin{array}[]{rcl}\Delta_{\mathcal{M}_{2}}&=&f_{2}(\mathcal{M}_{2})\cap(0,f_{2}(c^{\mathcal{M}_{2}})]\\ &=&\alpha\cdot\left\{f_{1}(\mathcal{M}_{1})\cap(0,f_{1}(c^{\mathcal{M}_{1}})]\right\}=\alpha\cdot\Delta_{\mathcal{M}_{1}}.\end{array}$
In either case, $\Delta_{\mathcal{M}_{1}}\sim\Delta_{\mathcal{M}_{2}}$. Conversely, suppose there is a positive real number $\alpha$ such that $\Delta_{\mathcal{M}_{2}}=\alpha\cdot\Delta_{\mathcal{M}_{1}}$. Then $f_{2}(\mathcal{M}_{2})=\alpha\cdot f_{1}(\mathcal{M}_{1})$ and $f_{2}(c^{\mathcal{M}_{2}})=\alpha f_{1}(c^{\mathcal{M}_{1}})$. Let $g=f_{2}^{-1}\alpha f_{1}\cup\{(0,0)\}$. Then a similar argument shows that $g$ is an isomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$. ∎
7. Concluding Remarks and Further Questions
We do not know the exact complexity of any of the equivalence relations considered in Theorem 6.2 in the Borel reducibility hierarchy. About $\approx_{\mathsf{CDVS}}$ we have the following observation.
Lemma 7.1.
$\approx_{\mathsf{CDVS}}$ is Borel reducible to an isomorphism relation.
Proof.
To each $\Delta\in\mathsf{CDVS}$ we associate a countable structure $\mathcal{A}_{\Delta}=(\Delta,R^{\mathcal{A}})$, where $R^{\mathcal{A}}$ is a ternary relation defined as
[…]
It is clear that for $\Delta,\Lambda\in\mathsf{CDVS}$, we have $\Delta\approx\Lambda$ iff $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$. This gives a Borel reduction from $\approx_{\mathsf{CDVS}}$ to an isomorphism relation. ∎
It follows that $\approx_{\mathsf{CDVS}}$ is Borel reducible to an equivalence relation induced by a Borel action of $S_{\infty}$. However, we do not know if $\approx_{\mathsf{CDVS}}$ or $\approx$ is Borel.
The main question left open is, of course, about the exact complexity of $\cong_{\mathsf{EA}}$ in the Borel reducibility hierarchy. Here we mention another class of automorphism groups known to be extremely amenable. Recall that an ultrametric is a metric satisfying $d(x,z)\leq\max\{d(x,y),d(y,z)\}$ for any $x,y,z$. An ordering $<$ on an ultrametric space $X$ is convex if for any $x\in X$ and $s>0$, the ball $B(x,s)=\{y\in X:d(x,y)\leq s\}$ is a convex set in the ordering $<$, i.e. if $y_{1}<z<y_{2}$ and $y_{1},y_{2}\in B(x,s)$, then $z\in B(x,s)$. An ultrametric space with a convex ordering is called a convexly-ordered ultrametric space.
Theorem 7.2 (Nguyen Van Thé [21]).
Let $S$ be a countable set of positive real numbers. Then the class of all finite convexly-ordered ultrametric spaces with nonzero distances all in $S$ is a Fraïssé class with the Ramsey property.
Denoting the Fraïssé limit of this class by $B_{S}^{c<}$, by the KPT correspondence, ${\rm Aut}(B_{S}^{c<})$ is extremely amenable. We note that if $S$ and $T$ are countable sets of positive real numbers that are order-isomorphic, then ${\rm Aut}(B_{S}^{c<})$ and ${\rm Aut}(B_{T}^{c<})$ are isomorphic as topological groups. If the converse is true, then we would obtain a new lower bound for the complexity of $\cong_{\mathsf{EA}}$, namely, the most complex one isomorphism relation for any class of countable structures. In particular, $\cong_{\mathsf{EA}}$ as a set would be complete analytic.
Finally, we note that there is a noticeable difference between Theorems 1.2 and 1.5, namely, in Theorem 1.2 we have the equivalence between the isomorphism of the topological groups and the isomorphism of them as abstract groups, while in Theorem 1.5 we do not have this equivalence. The reason for this equivalence to hold in Theorem 1.2 is that $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ has ample generics (see Corollary 4.1 of Solecki [30]), i.e., $G$ has a comeager orbit in the conjugacy action of $G$ on $G^{n}$, for each $n\geq 1$. By a result of Kechris–Rosendal [16], if a Polish group has ample generics, then every abstract homomorphism into a separable group is automatically continuous and thus it has a unique Polish topology. However, for the ordered case, Slutsky [29] showed that ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ does not have ample generics. Hence the question remains whether $\Delta$ and $\Lambda$ are equivalent if $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Lambda})$ are isomorphic as abstract groups.