### v2

# Extremely Amenable Automorphism Groups of Countable Structures

Abstract.
In this paper we address the question: How many pairwise non-isomorphic extremely amenable groups are there which are separable metrizable or even Polish? We show that there are continuum many such groups. In fact we construct continuum many pairwise non-isomorphic extremely amenable groups as automorphism groups of countable structures. We also consider this classification problem from the point of view of descriptive set theory by showing that the class of all extremely amenable closed subgroups of $S_{\infty}$ is Borel and their isomorphism relation is more complex than any isomorphism relation of countable structures in the Borel reducibility hierarchy.
2020 Mathematics Subject Classification:
22F05, 54H05
S. Gao and the F. Li acknowledge the partial support of their research by the National Natural Science Foundation of China (NSFC) grants 12271263 and 12250710128. R. Li acknowledges the partial support of his research by the NSFC grant 124B2001. S. Gao, F. Li and R. Li attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this research was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center.
1. Introduction
A topological group $G$ is extremely amenable if for every continuous action of $G$ on a compact Hausdorff space $X$, there is a fixed point $x\in X$, i.e., for all $g\in G$, $g\cdot x=x$. This is a remarkable property for a topological group to have. In fact, Veech [35] proved that no locally compact topological groups can be extremely amenable. This explained why extremely amenable groups were not present in the early development of topological group theory, for instance in the study of Hilbert’s fifth problem.
Ever since Mitchell [20] asked in 1970 about the existence of non-trivial extremely amenable topological groups, many examples of such groups have been discovered. At first, these groups were thought to be pathological since they had to be obtained via intricate constructions. However, as time goes by, more familiar groups have been shown to be extremely amenable. The extremely amenable groups we now know of can be loosely sorted into three classes as follows.
The first examples of extremely amenable groups were given by Herer–Christensen [13]. They constructed such groups which are abelian Polish groups of the form $L_{0}(\mu,\mathbb{R})$, which is the additive group of all $\mathbb{R}$-valued $\mu$-measurable functions, where $\mu$ is a pathological submeasure. Later, other examples of the more general form $L_{0}(\mu,G)$, where $G$ is a topological group, were shown to be extremely amenable. Among them, Furstenberg–Weiss (in unpublished work) and independently Glasner [10] showed that the group of measurable maps from the unit interval $\mathbb{I}$ with the Lebesgue measure $\lambda$ to the unit circle $\mathbb{T}$ is extremely amenable. Farah–Solecki [6] generalized it by showing that $L_{0}(\mu,G)$ is extremely amenable when $G$ is a compact solvable topological group and $\mu$ is a so-called diffused submeasure. Sabok [32] generalized further to arbitrary solvable topological groups $G$ and diffused submeaures $\mu$.
In a different thread of research, many operator groups were also shown to be extremely amenable. Famously, Gromov–Milman [12] showed that the unitary group $U(H)$ of the infinite dimensional separable Hilbert space $H$ is extremely amenable. Pestov proved the extreme amenability for the group $\mbox{Homeo}_{+}(\mathbb{I})$ of all order-preserving homeomorphisms of the unit interval ([28]) and for the group $\mbox{\rm Iso}(\mathbb{U})$ of all isometries of the universal Urysohn metric space $\mathbb{U}$ ([29]). Giordano–Pestov [9] showed the extreme amenability for $\mbox{\rm Aut}(\mathbb{I},\lambda)$, the group of all measure-preserving automorphisms of the Lebesgue space.
A third class of topological groups known to be extremely amenable in the literature comes from automorphism groups of countable structures. Pestov [28] gave the first example of such a group, which is ${\rm Aut}(\mathbb{Q},<)$, the automorphism group of a countable dense linear order without endpoints. Motivated by the idea in [28], Kechris–Pestov–Todorcevic [17] gave a complete characterization of when the automorphism group of a countable structure is extemely amenable, which is now called the KPT correspondence. The KPT correspondence connects the extreme amenability to the so-called structural Ramsey property of Fraïssé classes and gives a number of examples of extremely amenable groups. Through the work of Nešetřil–Rödl [22] [23] and Nešetřil [21] on Ramsey classes, we now have extremely amenable groups of the form $\mbox{\rm Aut}(M)$, where $M$ is the Fraïssé limit of a Fraïssé class of countable structures. Such structures include ordered graphs, ordered $K_{n}$-free graphs, ordered $\mathbb{Q}$-metric spaces, and ordered $\mathbb{N}$-metric spaces, etc. The KPT correspondence has also been generalized to the context of continuous logic by Melleray–Tsankov [19].
Despite the diversity of these examples, not much was known about the number of distinct isomorphism types of extremely amenable groups. Among the above examples, only finitely many pairwise non-isomorphic extremely amenable groups can be identified. This motivates us to study the following question: How many pairwise non-isomorphic extremely amenable groups are there?
The only result that explicitly addressed this question in the literature is a theorem of Pestov in [28], which states that for each infinite cardinal $\kappa$, there is an extremely amenable topological group of weight $\kappa$. Since the weight is an isomorphism invariant, we obtain many pairwise non-isomorphic extremely amenable groups. However, we would like to study this question in the context of more familiar groups, in particular separable metrizable groups and Polish groups. The theorem of Pestov gives only one such example.
Our main results of the paper confirm that there are continuum many non-isomorphic extremely amenable groups which are either Polish or separable metrizable. Our study focuses on the third class of examples in the above review.
The first main result of this paper confirms that there are continuum many pairwise non-isomorphic extremely amenable groups that are separable and metrizable. For each countable distance value set $\Delta$ (to be defined in Section 2), we consider the $\Delta$-metric Urysohn space $\mathbb{U}_{\Delta}$. The group $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ of all isometries of $\mathbb{U}_{\Delta}$, equipped with the topology of pointwise convergence in metric, is a separable metrizable group. We show the following theorems.
Theorem 1.1.
Let $\Delta$ be a countable distance value set. Then the following hold:
  1. (1)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable if and only if $\inf\Delta=0$.
  2. (2)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is Polish if and only if $\inf\Delta>0$.
  3. (3)
    If $\inf\Delta=0$, then $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of either $\mbox{\rm Iso}(\mathbb{U})$ or $\mbox{\rm Iso}(\mathbb{U}_{1})$, where $\mathbb{U}_{1}$ is the Urysohn sphere, i.e., the unit sphere of $\mathbb{U}$.
Theorem 1.2.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as topological groups.
  2. (ii)
    $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as abstract groups.
  3. (iii)
    $\Delta$ and $\Lambda$ are equivalent.
Theorem 1.3.
There are continnum many pairwise inequivalent countable distance value sets $\Delta$ that are dense in $(0,+\infty)$. Consequently, there are continuum many pairwise non-isomorphic extremely amenable groups $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ that are separable metrizable.
In fact, all groups in this class are dense in $\mbox{\rm Iso}(\mathbb{U})$; in particular, none of them are Polish.
[…]
Theorem 1.4.
[…]
We show the following theorems.
Theorem 1.5.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
[…]
  2. (ii)
    $\Delta$ and $\Lambda$ are equivalent.
Corollary 1.6.
[…]
The proofs of Theorems 1.2 and 1.5 follow the outline of the proof of a similar theorem in [5].
Our study of the isomorphism types of these extremely amenable groups goes much beyond counting the number of them. Another major effort we put in this paper is to study the isomorphism relation among these extremely amenable groups from the point of view of descriptive set theory. (For an overview of the descriptive set theory of equivalence relations, see Section 2). In view of Theorems 1.2 and 1.5, this is done by studying the complexity of the equivalence relation between countable distance value sets in the Borel reducibility hierarchy. We obtain the following penetrating results.
Theorem 1.7.
  1. (1)
    The equivalence between countable distance value sets $\Delta$ with $\inf\Delta>0$ and $\sup\Delta<\infty$ is Borel bireducible with the universal $S_{\infty}$-orbit equivalence relation.
  2. (2)
    The equivalence between countable distance value sets $\Delta$ with $\inf\Delta>0$ and $\sup\Delta=\infty$ is Borel bireducible with the equivalence relation $^{+}$.
  3. (3)
    The equivalence between countable distance value sets $\Delta$ with $\inf\Delta=0$ is strictly in between $^{+}$ and $^{++}$ in the Borel reducibility hierarchy.
These results significantly strengthen Theorem 1.3 and Corollary 1.6. Theorem 1.7 (3) is motivated by and related to the recent work of Calderoni–Marker–Motto Ros–Shani [2].
Finally, inspired by a research program initiated by Kechris–Nies–Tent [16], we show that the class of extremely amenable closed subgroups of $S_{\infty}$ is a Borel class (this was also independently proved by Iwanow–Majcher-Iwanow (unpublished); see [24, Section 3]), which implies that the isomorphism relation among these groups is an analytic equivalence relation. Regarding its complexity in the Borel reducibility hierarchy, our results yield the following.
Theorem 1.8.
The isomorphism relation among extremely amenable Polish groups that are isomorphic to closed subgroups of $S_{\infty}$ is more complex than the universal $S_{\infty}$-orbit equivalence relation. In particular, it is analytic and non-Borel.
The rest of the paper is organized as follows. In Section 2 we define the basic concepts and notions and give a review of the results we need to cite in this paper. In particular, we review the descriptive set theory of equivalence relations. In Section 3 we prove Theorems 1.1 and 1.4. In Section 4 we prove Theorems 1.2 and 1.5. In Section 5 we prove Theorem 1.7. Finally, in Section 6 we prove that the class of extremely amenable closed subgroups of $S_{\infty}$ is Borel, derive Theorem 1.8, and state some questions left open by this paper.
Acknowledgment. We thank Filippo Calderoni, Wei Dai, Aleksander Iwanow, Andre Nies, Marcin Sabok, Assaf Shani, Simon Thomas, and Víctor Hugo Yañez for their inspiring comments on an earlier version of the paper.
2. Preliminaries
In this section we review some of the basic concepts and notation we will use in the rest of the paper.
2.1. Topological dynamics
Our standard reference for topological dynamics is [1].
Our topological spaces and topological groups are all Hausdorff. Let $G$ be a topological group and let $X$ be a compact Hausdorff space. Let $a\colon G\times X\to X$ be a continuous action of $G$ on $X$. For $g\in G$ and $x\in X$, we also write $g\cdot x$ for $a(g,x)$. We say that $X$ is a $G$-flow. For $x\in X$, the orbit of $x$, denoted $x]_{G}$, is the set $\{g\cdot x\colon g\in G\}$. We say that the $G$-flow $X$ is minimal if for every $x\in X$, $x]_{G}$ is dense in $X$.
Suppose $G$ is a topological group and $X$, $Y$ are two $G$-flows. A continuous map $\varphi\colon X\to Y$ is equivariant or a homomorphism if for all $g\in G$ and $x\in X$, $\varphi(g\cdot x)=g\cdot\varphi(x)$. An isomorphism or a topological conjugacy from $X$ to $Y$ is an equivariant homeomorphism from $X$ onto $Y$. As usual, $X$ and $Y$ are isomorphic or topologically conjugate if there is an isomorphism from $X$ to $Y$.
A general result of topological dynamics states that for every topological group $G$, there is a universal minimal $G$-flow, i.e., a minimal $G$-flow which admits continuous equivariant maps onto all other minimal $G$-flows. Moreover, the universal minimal $G$-flow for a topological group $G$ is uniquely determined up to topological conjugacy.
A topological group $G$ is extremely amenable if every $G$-flow $X$ has a fixed point, i.e., there is $x\in X$ such that $g\cdot x=x$ for all $g\in G$. Equivalently, a topological group $G$ is extremely amenable if and only if the universal minimal $G$-flow is a singleton. By a theorem of Veech [35], no non-trivial locally compact topological group can be extremely amenable. Thus extreme amenability is a largeness property for topological groups.
2.2. Fraïssé Theory
Our standard reference for model theory and Fraïssé theory is [15].
A language $\mathcal{L}$ is a set consisting of the following:
  1. (a)
    a collection $(R_{i})_{i\in I}$ of relation symbols; for each $i\in I$, there is an associated arity $n(i)$, which is a positive integer, such that $R_{i}$ is an $n(i)$-ary relation symbol;
  2. (b)
    a collection $(F_{j})_{j\in J}$ of function symbols; for each $j\in J$, there is an associated arity $m(j)$, which is a natural number, such that $F_{j}$ is an $m(j)$-ary function symbol.
The 0-ary function symbols, if any, are also called constant symbols. We say $\mathcal{L}$ is countable if $I$ and $J$ are countable.
If $\mathcal{L}=((R_{i})_{i\in I},(F_{j})_{j\in J})$ is a language, an $\mathcal{L}$-structure is a tuple
$\mathcal{A}=(A;(R_{i}^{\mathcal{A}})_{i\in I},(F_{j}^{\mathcal{A}})_{j\in J})$
where
  1. (i)
    AA is a set known as the universe of $\mathcal{A}$ and denoted as $|\mathcal{A}|$;
  2. (ii)
    for each $i\in I$, $R_{i}^{\mathcal{A}}$ is the interpretation of $R_{i}$ in $\mathcal{A}$, which is an $n(i)$-ary relation on $A$, i.e., $R_{i}^{\mathcal{A}}\subseteq A^{n(i)}$;
  3. (iii)
    for each $j\in J$, $F_{j}^{\mathcal{A}}$ is the interpretation of $F_{j}$ in $\mathcal{A}$, which is an $m(j)$-ary function from $A^{m(j)}$ to $A$, i.e., $F_{j}^{\mathcal{A}}\colon A^{m(j)}\to A$.
If $m(j)=0$, $F_{j}^{\mathcal{A}}\in A$ is a constant. $\mathcal{A}$ is finite or countable if $|\mathcal{A}|$ is finite or countable, respectively.
Let $\mathcal{L}$ be a language and let $\mathcal{A}$ and $\mathcal{B}$ be $\mathcal{L}$-structures. By an embedding from $\mathcal{A}$ to $\mathcal{B}$, we mean an injection $f\colon|\mathcal{A}|\rightarrow|\mathcal{B}|$ such that for all $i\in I$ and $a_{1},\dots,a_{n(i)}\in A$,
$R_{i}^{\mathcal{A}}(a_{1},\dots,a_{n(i)})\iff R_{i}^{\mathcal{B}}(f(a_{1}),\dots,f(a_{n(i)})),$
and for all $j\in J$ and $a_{1},\dots,a_{m(j)}\in A$,
$f(F_{j}^{\mathcal{A}}(a_{1},\dots,a_{m(j)}))=F_{j}^{\mathcal{B}}(f(a_{1}),\dots,f(a_{m(j)})).$
If $f$ is the inclusion map, then we say $\mathcal{A}$ is a substructure of $\mathcal{B}$, denoted by $\mathcal{A}\subseteq\mathcal{B}$. If $S\subseteq|\mathcal{B}|$, then the substructure of $\mathcal{B}$ generated by $S$ is the smallest substructure of $\mathcal{B}$ whose universe contains $S$ as a subset. In this case $S$ is called a generating set for the substructure generated by $S$. A substructure is finitely generated if it has a finite generating set.
If there is an embedding from $\mathcal{A}$ to $\mathcal{B}$, we say $\mathcal{A}$ is embeddable to $\mathcal{B}$, denoted by $\mathcal{A}\leq\mathcal{B}$. Finally, if $f$ is bijective, we say it is an isomorphism; if there is an isomorphism between $\mathcal{A}$ and $\mathcal{B}$, we say they are isomorphic, and denote $\mathcal{A}\cong\mathcal{B}$.
For the rest of the discussion, fix a countable language $\mathcal{L}$. A class $\mathcal{K}$ of finite $\mathcal{L}$-structures is said to be a Fraïssé class if it satisfies the following conditions:
  1. (1)
    Up to isomorphism, there are only countably many different structures in $\mathcal{K}$.
  2. (2)
    There are structures of arbitarily large finite cardinality in $\mathcal{K}$.
  3. (3)
    Hereditary property (HP): If $\mathcal{A}\in\mathcal{K}$, then every finite substructure of $\mathcal{A}$ is in $\mathcal{K}$.
  4. (4)
    Joint embedding property (JEP): If $\mathcal{A},\mathcal{B}\in\mathcal{K}$, then there is a $\mathcal{C}\in\mathcal{K}$ such that $\mathcal{A}\leq\mathcal{C}$ and $\mathcal{B}\leq\mathcal{C}$.
  5. (5)
    Amalgamation property (AP): For any embeddings $f:\mathcal{A}\rightarrow\mathcal{B}$ and $g:\mathcal{A}\rightarrow\mathcal{C}$, there is a $\mathcal{D}\in\mathcal{K}$ and embeddings $r:\mathcal{B}\rightarrow\mathcal{D}$ and $s:\mathcal{C}\rightarrow\mathcal{D}$ such that $r\circ f=s\circ g$.
A countable $\mathcal{L}$-structure $\mathcal{A}$ is locally finite if every finitely generated substructure of $\mathcal{A}$ is finite. The age of $\mathcal{A}$, denoted by ${\rm Age}(\mathcal{A})$, is the class of finite substructures of $\mathcal{A}$. We say $\mathcal{A}$ is ultrahomogeneous if every isomorphism between two finite substructures can be extended to an automorphism of $\mathcal{A}$. We call $\mathcal{A}$ a Fraïssé structure if it is locally finite, countably infinite and ultrahomogeneous.
Fraïssé [7] (c.f. [15]) gave a correspondence between Fraïssé classes and Fraïssé structures, as follows. If $\mathcal{K}$ is a Fraïssé class, then there is a unique Fraïssé structure $\mathcal{A}$, up to isomorphism, such that ${\rm Age}(\mathcal{A})=\mathcal{K}$. This unique structure is called the Fraïssé limit of $\mathcal{K}$, denoted by $\mbox{Flim}\,{\mathcal{K}}$. Conversely, if $\mathcal{A}$ is a Fraïssé structure, then ${\rm Age}(\mathcal{A})$ is a Fraïssé class.
We remark here that the original setting of Fraïssé theory allows a more general definition of Fraïssé classes than we presented above. For example, in the above definition of Fraïssé classes, every mention of finite substructures can be replaced by finitely generated substructures and condition (2) can be omitted. Correspondingly, in the definition of Fraïssé structures, the structure does not need to be locally finite or infinite. We define these notions as above just because the KPT correspondence requires them.
The countable language $\mathcal{L}$ we will consider in this paper are all relational, i.e., they consists only of relation symbols and have no function symbols. In this case, every structure is locally finite, and the age of a structure consists of exactly all finitely generated substructures.
We note the following equivalent condition for a countable structure to be ultrahomogeneous.
Definition 2.1 ([15]).
Let $\mathcal{L}$ be a countable language. A countable $\mathcal{L}$-structure $\mathcal{A}$ is weakly homogeneous if for all $\mathcal{X}$ and $\mathcal{Y}$ in ${\rm Age}(\mathcal{A})$ with $\mathcal{X}\subseteq\mathcal{Y}$ and embedding $f$ from $\mathcal{X}$ to $\mathcal{A}$, $f$ can be extended to an embedding $g$ from $\mathcal{Y}$ to $\mathcal{A}$.
It turns out that for countable structures, ultrahomogeneity and weak homogeneity are equivalent. This equivalent condition will be used later.
2.3. The KPT correspondence
What has become known as the KPT correspondence was developed in Kechris–Pestov–Todorcevic [17]. We review the basic definitions and results here.
A Polish space is a separable, completely metrizable topological space. A Polish group is a topological group whose topology is Polish. An example of a Polish group is the infinite permutation group $_{\infty}. SS_{\infty}$ consists of all permutations of $\mathbb{N}$, where the group operation is composition. The topology on $S_{\infty}$ is the pointwise convergence topology, where the topology on $\mathbb{N}$ is discrete. Equivalently, $S_{\infty}$ has the subspace topology inherited from $\mathbb{N}^{\mathbb{N}}$, where $\mathbb{N}^{\mathbb{N}}$ has the product topology. Both the topologies on $\mathbb{N}^{\mathbb{N}}$ and on $S_{\infty}$ are Polish.
Throughout the rest of this subsection we fix a countable language $\mathcal{L}$. For a countable $\mathcal{L}$-structure $\mathcal{A}$, we use ${\rm Aut}(\mathcal{A})$ to denote the group of all automorphisms of $\mathcal{A}$, where the group operation is composition. Considering the discrete topology on $|\mathcal{A}|$, and equipping $\mbox{\rm Aut}(\mathcal{A})$ with the topology of pointwise convergence, $\mbox{\rm Aut}(\mathcal{A})$ becomes a Polish group. In fact, it can be regarded as a closed subgroup of the permutation group $S_{|\mathcal{A}|}$. When $|\mathcal{A}|$ is infinite, $S_{|\mathcal{A}|}$ is isomorphic to $S_{\infty}$. Conversely, if $G$ is a closed subgroup of $S_{\infty}$, then there is a countable structure $\mathcal{A}$ such that $G$ and $\mbox{\rm Aut}(\mathcal{A})$ are isomorphic as topological groups.
A Fraïssé class $\mathcal{K}$ is said to be ordered if there is a binary relation symbol << in $\mathcal{L}$ and, in every structure in $\mathcal{K}$, the interpretaton of << is a linear order. The KPT correspondence unravels a deep connection between the property of a Fraïssé ordered class $\mathcal{K}$ and the extreme amenability of $\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$. This property is called the (structural) Ramsey property, which we define below.
Let $\mathcal{K}$ be a class of finite structures and let $\mathcal{A},\mathcal{B}\in\mathcal{K}$. We use the following notation to denote the set of all substructures of $\mathcal{B}$ which are isomorphic to $\mathcal{A}$:
$\begin{pmatrix}\mathcal{B}\\ \mathcal{A}\end{pmatrix}=\{\mathcal{A^{\prime}}\colon\mathcal{A^{\prime}}\subseteq\mathcal{B},\mathcal{A^{\prime}}\cong\mathcal{A}\}.$
Then for $\mathcal{A}\leq\mathcal{B}\leq\mathcal{C}$ and $k\geq 2$, we use the Erdős-Rado arrow,
$\mathcal{C}\rightarrow(\mathcal{B})^{\mathcal{A}}_{k}$
to mean that for any function $c:\begin{pmatrix}\mathcal{C}\\ \mathcal{A}\end{pmatrix}\rightarrow k$, there is a $\mathcal{B^{\prime}}\in\begin{pmatrix}\mathcal{C}\\ \mathcal{B}\end{pmatrix}$ such that $c$ is constant on $\begin{pmatrix}\mathcal{B^{\prime}}\\ \mathcal{A}\end{pmatrix}$. Finally, we say $\mathcal{K}$ has the Ramsey property if for any $\mathcal{A}\leq\mathcal{B}\in\mathcal{K}$ and $k\geq 2$, there is a $\mathcal{C}\in\mathcal{K}$ such that $\mathcal{C}\rightarrow(\mathcal{B})^{\mathcal{A}}_{k}$.
Theorem 2.1 (Kechris-Pestov-Todorcevic [17]).
Let $\mathcal{K}$ be a Fraïssé ordered class and let $\mathcal{A}$ be the Fraïssé limit of $\mathcal{K}$. Then the following are equivalent:
  1. (1)
    ${\rm Aut}(\mathcal{A})$ is extremely amenable.
  2. (2)
    $\mathcal{K}$ has the Ramsey property.
The simplest example of a Fraïssé class with the Ramsey property is the class of all finite sets. In fact, the Ramsey property for this class is equivalent to the classical Ramsey theorem. Also, the class of all finite linear orders has the Ramsey property, which is again equivalent to the classical Ramsey theorem; since the Fraïssé limit of this class is the countable dense linear order without endpoints, it follows from the above theorem that ${\rm Aut}(\mathbb{Q},<)$ is extremely amenable. (Historically, Pestov [28] proved the extreme amenability of $\mbox{\rm Aut}(\mathbb{Q},<)$ first, which motivated the above theorem.) Many more examples of Fraïssé ordered classes with the Ramsey property are now known and the automorphism groups of their Fraïssé limits are all extremely amenable by the KPT correspondence. For an excellent survey of these results, see Nguyen Van Thé [27].
We will also use the following strengthening of Theorem 2.1, which is Theorem 3 of Nguyen Van Thé [27].
Theorem 2.2 ([27]).
Let $\mathcal{L}$ be a countable language, let $\mathcal{K}$ be a Fraïssé class of finite $\mathcal{L}$-structures, and let $\mathcal{A}$ be the Fraïssé limit of $\mathcal{K}$. Then the following are equivalent:
  1. (1)
    $\mbox{\rm Aut}(\mathcal{A})$ is extremely amenable.
  2. (2)
    $\mathcal{K}$ has the Ramsey property and every element of $\mathcal{K}$ is rigid, i.e., has trivial automorphism group.
2.4. Descriptive set theory of equivalence relations
In this paper we will be discussing the complexity of equivalence relations in the framework of descriptive set theory. We review the basic concepts and notions in this subsection. For other undefined terms and concepts, see [8].
Let $X$ be a set and let $\mathcal{B}$ be a $\sigma$-algebra of subsets of $X$. $(X,\mathcal{B})$ is a standard Borel space if there is a Polish topology $\tau$ on $X$ such that $\mathcal{B}$ is the $\sigma$-algebra of Borel sets generated (via iterations of countable union, countable intersection, and complementation) by $\tau$. Any Polish space is a standard Borel space. If $X$ is a standard Borel space and $Y$ is a Borel subset of $X$, then $Y$ is a standard Borel space. A subset $A$ of a standard Borel space is analytic if there is a standard Borel space $Y$ and there is a Borel function $f\colon Y\to X$ with $A=f(Y)$. An equivalence relation $E$ on a standard Borel space $X$ is Borel (or analytic) if $E$ is a Borel (analytic) subset of $X^{2}$.
For equivalence relations $E,F$ on standard Borel spaces $X,Y$, respectively, we say that $E$ is Borel reducible to $F$, denoted $E\leq_{B}F$, if there is a Borel function $f\colon X\to Y$ such that for all $x,x^{\prime}\in X$, we have
$xEx^{\prime}\iff f(x)Ff(x^{\prime}).$
We say that $E$ is Borel bireducible with $F$, denoted $E\sim_{B}F$, if both $E\leq_{B}F$ and $F\leq_{B}E$ hold. We say that $E$ is strictly Borel reducible to $F$, denoted $E<_{B}F$, if $E\leq_{B}F$ but $F\not\leq_{B}E$. The Borel reducibility preorder $\leq_{B}$ gives rise to a relative complexity hierarchy for equivalence relations on standard Borel spaces. Specifically, given an equivalence relation $E$, we attempt to determine its exact complexity by comparing it, in terms of Borel reducibility, with some benchmark equivalence relations previously studied and understood in the theory. If there is a benchmark equivalence relation $F$ so that we can show $E\sim_{B}F$, then we consider the exact complexity of $E$ to be completely determined. Short of doing this, we can establish upper bounds $E\leq_{B}F$ or $E<_{B}F$ for $E$, or lower bounds $F\leq_{B}E$ or $F<_{B}E$ for $E$, as much as we can, which provide partial information on the complexity of $E$.
We will use the following benchmark equivalence relations in this paper. Discussions about the ones without references can be found in [8].
  1. (1)
    $_{\mathbb{N}}$ is the equality relation on $\mathbb{N}$, which is a Polish space with the discrete topology. If $_{\mathbb{N}}\ \leq_{B}E$, then $E$ contains infinitely many equivalence classes.
  2. (2)
    $_{\mathbb{R}}$ is the equality relation on $2^{\mathbb{N}}$. If $_{\mathbb{R}}\ \leq_{B}E$, then $E$ contains continuum many equivalence classes.
  3. (3)
    $E_{0}$ is the eventual agreement relation on $2^{\mathbb{N}}$ defined as (x_i)_iE_0(y_i)_i⇔∃i ∀j≥i x_j=y_j.
  4. (4)
    $E_{\infty}$ is the universal countable Borel equivalence relation. It can be defined as the following equivalence relation on $2^{\mathbb{F}_{2}}$, where $\mathbb{F}_{2}$ is the free group with two generators: (x_g)_gE_∞(y_g)_g ⇔∃h∈F_2 ∀g∈F_2 x_g=y_gh.
  5. (5)
    $^{+}$ is the equivalence relation on $(2^{\mathbb{N}})^{\mathbb{N}}$ defined as (x_i)_i =^+ (y_i)_i⇔{x_i:i∈N}={y_i:i∈N}.
  6. (6)
    $^{++}$ is the equivalence relation on $((2^{\mathbb{N}})^{\mathbb{N}})^{\mathbb{N}}$ defined as (x_i)_i =^++ (y_i)_i⇔∀i ∃j x_i=^+ y_j and ∀j ∃i x_i=^+y_j.
  7. (7)
    Let $\mathcal{P}$ be the powerset operation on sets. Let TC denote the transitive closure operation on sets. For $k=0,1$, let $\mathcal{Q}^{*}_{k}(\mathbb{N})$ be the collection of pairs $(A,R)$ such that:
    1. a)
      AA is a hereditarily countable set in $\mathcal{P}^{3}(\mathbb{N})$;
    2. b)
      $R$ is a ternary relation on $A\times A\times(\mathcal{P}^{k}(\mathbb{N})\cap\mbox{\rm TC}(A))$ such that
      • for any $a,b\in A$ there is some $r$ such that $R(a,b,r)$ holds;
      • given any $a\in A$, for any $b,b^{\prime}\in A$ and any $r$, if $R(a,b,r)$ and $R(a,b^{\prime},r)$ both hold then $b=b^{\prime}$.
    The equivalence relation $\cong^{*}_{3,k}$ is defined as the isomorphism relation of countable structures coding pairs $(A,R)$ in $\mathcal{Q}^{*}_{k}(\mathbb{N})$. These equivalence relations were defined by Hjorth–Kechris–Louveau [14]. Our presentation above follows [2].
  8. (8)
    $E^{\infty}_{S_{\infty}}$ is the universal equivalence relation among all $S_{\infty}$-orbit equivalence relations, i.e., for any Borel action of $S_{\infty}$ on a standard Borel space $X$, the orbit equivalence relation $E^{X}_{S_{\infty}}=\{(x,y)\in X\colon\exists g\in S_{\infty}\;g\cdot x=y\}$ is Borel reducible to $E^{\infty}_{S_{\infty}}$. This equivalence relation is known to be universal among all isomorphism relations of countable structures. That is, given any Borel class of countable structures, the isomorphism relation is Borel reducible to $E^{\infty}_{S_{\infty}}$. Moreover, $E^{\infty}_{S_{\infty}}$ is also known to be Borel bireducible with many isomorphism relations of countable structures. For example, $E^{\infty}_{S_{\infty}}$ is Borel bireducible with the isomorphism relation for all countable linear orders.
The following Borel reducibility results are well known (see [8]):
$_{\mathbb{N}}\ <_{B}\ =_{\mathbb{R}}\ <_{B}E_{0}\,<_{B}E_{\infty}<_{B}\ =^{+}\,<_{B}\ =^{++}\ <_{B}E^{\infty}_{S_{\infty}}.$
It was also shown in Hjorth–Kechris–Louveau [14] that
$^{+}\,<_{B}\ \cong^{*}_{3,0}\ <_{B}\ \cong^{*}_{3,1}\ <_{B}\ =^{++}.$
The equivalence relations in (1) through (7) above are Borel equivalence relations, and $E^{\infty}_{S_{\infty}}$ is analytic and non-Borel.
Kechris–Nies–Tent [16] initiated the following research program.
Program.
  1. (1)
    For natural classes of closed subgroups of $S_{\infty}$, determine if they are Borel as subsets of $\mathcal{F}(S_{\infty})$.
  2. (2)
    If a class is Borel, study the relative complexity of the topological isomorphism relation, using Borel reducibility $\leq_{B}$.
Here $\mathcal{F}(S_{\infty})$ is the standard Borel space consisting of all closed subsets of $S_{\infty}$. The $\sigma$-algebra of Borel sets on $\mathcal{F}(S_{\infty})$ is generated by sets of the form
$\{F\in\mathcal{F}(S_{\infty})\colon F\cap U\neq\emptyset\},$
where $U\subseteq S_{\infty}$ is a basic open set, i.e., of the form
$N_{\bar{a},\bar{b}}=\{g\in S_{\infty}\colon g(\bar{a})=\bar{b}\}$
for some tuples $\bar{a},\bar{b}$ of natural numbers of the same length.
In this paper we will consider the class of closed subgroups of $S_{\infty}$ which are extremely amenable. We will show that it is a Borel subset of $\mathcal{F}(S_{\infty})$ (Theorem 6.1), thus this class fits well in the KNT research program.
3. The Extreme Amenability of Automorphism Groups
3.1. $\Delta$-metric spaces
We work with the following notion.
Definition 3.1 ([5]).
Let $\Delta$ be a nonempty set of positive real numbers. We call $\Delta$ a distance value set if for all $x,y\in\Delta$,
$\min\{x+y,\sup\Delta\}\in\Delta.$
Briefly speaking, a distance value set is exactly an initial segment of a subsemigroup of positive reals. Given a distance value set $\Delta$, a metric space $(X,d)$ is a $\Delta$-metric space if $d(x,y)\in\Delta$ for all $x\neq y\in X$.
Fix a countable distance value set $\Delta$. We can view each countable $\Delta$-metric space as a countable structure, as follows. Let $\mathcal{L}_{\Delta}=\{R_{s}\colon s\in\Delta\}$, where each $R_{s}$ is a binary relation symbol. Given any countable $\Delta$-metric space $(X,d)$, set $R^{X}_{s}(x,y)$ if and only if $d(x,y)=s$. Thus we turn $(X,d)$ into a countable $\mathcal{L}_{\Delta}$-structure $(X;(R^{X}_{s})_{s\in\Delta})$. Let $\mbox{\rm Aut}(X,d)$, or simply $\mbox{\rm Aut}(X)$ if $d$ is understood, denote the automorphism group $\mbox{\rm Aut}(X;(R^{X}_{s})_{s\in\Delta})$. Then $\mbox{\rm Aut}(X)$ is a Polish group.
However, in this section we will consider another topology, and thus we will use a different notation. Let $\mbox{\rm Iso}(X,d)$ or simply $\mbox{\rm Iso}(X)$ if $d$ is understood, denote the group of all autoisometries of $X$, i.e., $f\in\mbox{\rm Iso}(X)$ if $f\colon X\to X$ satisfies $d(f(x),f(y))=d(x,y)$ for all $x,y\in X$. Equip $\mbox{\rm Iso}(X)$ with the topology of pointwise convergence in metric, i.e., the topology given by basic open sets of the form
$\{f\in\mbox{\rm Iso}(X)\colon d(f(x_{1}),y_{1})<\epsilon_{1},\dots,d(f(x_{n}),y_{n})<\epsilon_{n}\}$
where $x_{1},\dots,x_{n},y_{1},\dots,y_{n}\in X$ and $\epsilon_{1},\dots,\epsilon_{n}$ are positive real numbers. Then $\mbox{\rm Iso}(X)$ is a separable metrizable group, and it is Polish when $d$ is a complete metric on $X$. Let $(\bar{X},\bar{d})$ be the completion of $(X,d)$. Then $\mbox{\rm Iso}(X)$ is isomorphic to a subgroup of $\mbox{\rm Iso}(\bar{X})$, and $\mbox{\rm Iso}(\bar{X})$ is Polish.
Importantly, note that $\mbox{\rm Aut}(X)$ and $\mbox{\rm Iso}(X)$ are identical as sets, and the identity map is a group isomorphism.
Fix a countable distance value set $\Delta$. Let $\mathcal{K}_{\Delta}$ be the class of all finite $\Delta$-metric spaces. Then $\mathcal{K}_{\Delta}$ is a Fraïssé class. We denote the Fraïssé limit of $\mathcal{K}_{\Delta}$ by $\mathbb{U}_{\Delta}$ and call it the $\Delta$-Urysohn space. In the case that $\Delta=\mathbb{Q}$, the resulting Fraïssé limit $\mathbb{U}_{\mathbb{Q}}$ is the well-known rational Urysohn space and has been denoted by $\mathbb{QU}$ in the literature. The completion of $\mathbb{QU}$ is the universal Urysohn space and has been denoted by $\mathbb{U}$ in the literature. $\mathbb{U}$ is universal for all separable complete metric spaces, i.e., for any separable complete metric space $(X,d)$, there is an isometric embedding from $X$ into $\mathbb{U}$.
3.2. The extreme amenability of $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$
To apply the KPT correspondence, we need to consider a Fraïssé ordered class. Let $\Delta$ be a countable distance value set. Let $\mathcal{K}_{\Delta}^{<}$ be the class of all finite ordered $\Delta$-metric spaces, i.e., each element of $\mathcal{K}_{\Delta}^{<}$ is a structure $(X;(R_{s}^{X})_{s\in\Delta},<^{X})$, where $^{X}$ is a linear order on $X$. It is easy to see that $\mathcal{K}_{\Delta}^{<}$ is a Fraïssé (ordered) class. We denote its Fraïssé limit by $\mathbb{U}_{\Delta}^{<}$. It can be shown that $\mathbb{U}_{\Delta}^{<}\cong(\mathbb{U}_{\Delta},<^{\mathbb{U}_{\Delta}})$ where $^{\mathbb{U}_{\Delta}}$ is a linear order isomorphic to $(\mathbb{Q},<)$ (see, e.g., Proposition 5.2 of [17]).
The following is a theorem of Nešetřil [21]; Ngyen Van Thé [26] stated the result more explicitly (as Theorem 13 of [26]) and gave a full proof.
Theorem 3.1 (Nešetřil).
Let $\Delta$ be a countable distance value set. Then $\mathcal{K}_{\Delta}^{<}$ has the Ramsey property.
By the KPT correspondence, we get that $\mbox{\rm Aut}(\mathbb{U}_{\Delta}^{<})$ is extremely amenable, which is Theorem 1.4.
3.3. The extreme amenability of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
In this subsection we prove Theorem 1.1. We first consider a countable distance value set $\Delta$ with $\inf\Delta=0$. Note that this condition is equivalent to that $\Delta$ is dense in $(0,\sup\Delta)$. In this case we show that $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable. Our proof follows the outline of the proof of Theorem 6.17 of [17], which argues the extreme amenability of ${\rm Iso}(\mathbb{U})$ from the extreme amenability of ${\rm Aut}(\mathbb{U}_{\mathbb{Q}}^{<})$. We use the following basic fact (Lemma 6.18 of [17]).
Lemma 3.2 ([17]).
Let $G,H$ be topological groups and let $f$ be a continuous group homomorphism from $G$ into $H$ with dense image. If $G$ is extremely amenable, so is $H$.
[…]
Proposition 3.3.
Suppose $\Delta$ is a countable distance value set with $\inf\Delta=0$. Given $x_{1},\dots,x_{n},y_{1},\dots,y_{n}\in\mathbb{U}_{\Delta}$ such that $x_{i}\mapsto y_{i}$, $1\leq i\leq n$, is a partial isometry, and given $\epsilon>0$, there exist $y_{1}^{\prime},\dots,y_{n}^{\prime}\in\mathbb{U}_{\Delta}$ so that $x_{i}\mapsto y_{i}^{\prime}$, $1\leq i\leq n$, is an order-preserving (with respect to $^{\mathbb{U}_{\Delta}}$) partial isometry, and for all $1\leq i\leq n$, $d(y_{i},y_{i}^{\prime})<\epsilon$.
Proof.
Let $Z=\{y_{1},\dots,y_{n},z_{1},\dots,z_{n}\}$ be a finite ordered $\Delta$-metric space with the following properties:
  1. (i)
    $d_{Z}(y_{i},y_{j})=d_{Z}(z_{i},z_{j})=d_{\mathbb{U}_{\Delta}}(y_{i},y_{j})$ for $1\leq i<j\leq n$,
  2. (ii)
    $d_{Z}(y_{i},z_{j})=\delta+d_{\mathbb{U}_{\Delta}}(y_{i},y_{j})$ for $1\leq i\leq j\leq n$,
  3. (iii)
    $y_{i}<^{Z}y_{j}$ if and only if $y_{i}<y_{j}$ for $1\leq i<j\leq n$,
  4. (iv)
    $y_{i}<^{Z}z_{j}$ for $1\leq i,j\leq n$,
  5. (v)
    $z_{i}<^{Z}z_{j}$ if and only if $x_{i}<x_{j}$ for $1\leq i<j\leq n$,
where $\delta\in\Delta$ and $\delta<\epsilon$. Note that $(Z,d_{Z})$ is a metric space. Thus, by the universality and the ultrahomogeneity of $\mathbb{U}_{\Delta}^{<}\cong(\mathbb{U}_{\Delta},<^{\mathbb{U}_{\Delta}})$, there are $\{y_{i}^{\prime}\}_{i=1}^{n}$ in $\mathbb{U}_{\Delta}$ such that $y_{i}\mapsto y_{i}$ and $z_{i}\mapsto y_{i}^{\prime}$ is an order-preserving isometry from $Z$ into $\mathbb{U}_{\Delta}^{<}$. These $y_{i}^{\prime}$ for $1\leq i\leq n$ are as desired. ∎
We are now ready to prove Theorem 1.1, which we restate below.
Theorem 3.4.
Let $\Delta$ be a countable distance value set. Then:
  1. (1)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable if and only if $\inf\Delta=0$.
  2. (2)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ is Polish if and only if $\inf\Delta>0$.
  3. (3)
    If $\inf\Delta=0$, then ${\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of either ${\rm Iso}(\mathbb{U})$ or $\mbox{\rm Iso}(\mathbb{U}_{1})$.
Proof.
[…]
On the other hand, assume $\inf\Delta>0$. Then $\mathbb{U}_{\Delta}$ has discrete topology, and it follows that $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ are isomorphic as topological groups. It is clear that not every finite $\Delta$-metric space is rigid. By Theorem 2.2, $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ is not extremely amenable. Hence $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is not extremely amenable.
Next we prove (3). Recall that Vershik [36] (Theorem 3) has shown that if $\Delta$ is dense in $(0,+\infty)$, then $\mathbb{U}_{\Delta}$ is dense in $\mathbb{U}$, hence the metric completion of $\mathbb{U}_{\Delta}$ is $\mathbb{U}$. Similarly, if $\Delta$ is bounded and dense in $(0,\sup\Delta)$, then the metric completion of $\mathbb{U}_{\Delta}$ is just $\mathbb{U}_{a}$ where $a=\sup\Delta$ and $\mathbb{U}_{a}$ is a sphere of $\mathbb{U}$ with diameter $a$. Thus, given any isometry $f\in\mbox{\rm Iso}(\mathbb{U}_{\Delta})$, there is a unique isometry $\hat{f}\in\mbox{\rm Iso}(\mathbb{U})$ (or $\mbox{\rm Iso}(\mathbb{U}_{a})$ if $a=\sup\Delta<+\infty$) extending $f$. Therefore we can view $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ as the subgroup of $\mbox{\rm Iso}(\mathbb{U})$ (or of $\mbox{\rm Iso}(\mathbb{U}_{a})$) consisting of exactly those isometries fixing $\mathbb{U}_{\Delta}$ setwise. It was proved in [3] that if $\Delta=\mathbb{Q}$, then $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U})$. The argument in [3] can be generalized to prove that if $\Delta$ is dense in $(0,+\infty)$ then ${\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of ${\rm Iso}(\mathbb{U})$. The same argument also shows that when $\inf\Delta=0$ and $\Delta$ is bounded, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{a})$, where $a=\sup\Delta$. However, for any $0<a<\infty$, $\mbox{\rm Iso}(\mathbb{U}_{a})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$ are isomorphic. Hence in this case $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is isomorphic to a dense subgroup of $\mbox{\rm Iso}(\mathbb{U}_{1})$.
(2) If $\inf\Delta>0$, then ${\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ are isomorphic as topological groups, while the latter is Polish. Hence $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is Polish. On the other hand, suppose $\inf\Delta=0$. First suppose $\Delta$ is unbounded, i.e., $\Delta$ is dense in $(0,+\infty)$. Then by the above argument ${\rm Iso}(\mathbb{U}_{\Delta})$ is dense in ${\rm Iso}(\mathbb{U})$. Note that $\mbox{\rm Iso}(\mathbb{U})$ is Polish. So if ${\rm Iso}(\mathbb{U}_{\Delta})$ was Polish, it would be a closed subgroup of $\mbox{\rm Iso}(\mathbb{U})$ and thus $\mbox{\rm Iso}(\mathbb{U}_{\Delta})=\mbox{\rm Iso}(\mathbb{U})$. This is impossible since we can easily find an isomertry of $\mathbb{U}$ not setwise fixing $\mathbb{U}_{\Delta}$. Next suppose $\Delta$ is bounded and let $a=\sup\Delta$. Then by the above argument $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is dense in $\mbox{\rm Iso}(\mathbb{U}_{a})$. If $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ was Polish, we again have $\mbox{\rm Iso}(\mathbb{U}_{\Delta})=\mbox{\rm Iso}(\mathbb{U}_{a})$, which yields a similar contradiction. ∎
By Theorem 3.4 (1) and (2), for any countable distance value set $\Delta$, the separable metrizable group ${\rm Iso}(\mathbb{U}_{\Delta})$ can never be both Polish and extremely amenable. When $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is not Polish, one typically takes the metric completion of $\mathbb{U}_{\Delta}$ and considers its isometry group, which is Polish. Clause (3) of Theorem 3.4 says that this obtains essentially only two Polish groups, $\mbox{\rm Iso}(\mathbb{U})$ and $\mbox{\rm Iso}(\mathbb{U}_{1})$, both known to be extremely amenable by Pestov [29].
We remark here that Theorem 3.4 (1) follows from a less well-known result of Gould. In [11, Theorem A.2.2 (4)] he showed that a topological group is extremely amenable if and only if all of its dense subgroups are extremely amenable. By this result, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ is extremely amenable when $\inf\Delta=0$ following the density observations and results of Pestov mentioned in the preceding paragraph. The authors thank Yañez for pointing out the reference [11] to us.
In addition, Dai–Gao–Yañez [4] recently proved that for countable distance value sets $\Delta$ with $\inf\Delta=0$, $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ has the so-called strong Lévy property. This also implies that they are extremely amenable by a result of Gromov–Milman [12].
4. Isomorphism Types of Extremely Amenable Groups
[…]
Definition 4.1 ([5]).
Let $\Delta$ and $\Lambda$ be distance value sets.
  1. (i)
    A triple $(x,y,z)$ of positive real numbers is a $\Delta$-triangle if $x,y,z\in\Delta$ and $|x-y|\leq z\leq x+y$.
  2. (ii)
    We say that $\Delta$ and $\Lambda$ are equivalent if there is a bijection $f\colon\Delta\to\Lambda$ such that for any positive real numbers $x,y,z$, $(x,y,z)$ is a $\Delta$-triangle if and only if $(f(x),f(y),f(z))$ is a $\Lambda$-triangle.
4.1. Isomorphism types of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
In this subsection we characterize the topological isomorphism relation of groups of the form $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ for countable distance value sets $\Delta$ with $\inf\Delta=0$. We use the following theorem from [5].
Theorem 4.1 ([5]).
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    ${\rm Aut}(\mathbb{U}_{\Delta})$ and ${\rm Aut}(\mathbb{U}_{\Lambda})$ are isomorphic as abstract groups.
  2. (ii)
    ${\rm Aut}(\mathbb{U}_{\Delta})$ and ${\rm Aut}(\mathbb{U}_{\Lambda})$ are isomorphic as topological groups.
  3. (iii)
    $\Delta$ and $\Lambda$ are equivalent.
Our result is similar. The following is a restatement of Theorem 1.2.
Theorem 4.2.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ and ${\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as abstract groups.
  2. (ii)
    ${\rm Iso}(\mathbb{U}_{\Delta})$ and ${\rm Iso}(\mathbb{U}_{\Lambda})$ are isomorphic as topological groups.
  3. (iii)
    $\Delta$ and $\Lambda$ are equivalent.
Proof.
To see that (i) implies (iii), the key oberservation is that $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ are identical abstract groups for any $\Delta$. Thus this implication follows from the implication of (iii) from (i) in Theorem 4.1.
The implication of (i) from (ii) is obvious.
To see that (iii) implies (ii), fix a bijection $f$ from $\Delta$ to $\Lambda$ which witnesses the equivalence of the two distance value sets. Let $d$ be the metric on $\mathbb{U}_{\Delta}$. Define a new metric $d_{f}$ on $\mathbb{U}_{\Delta}$ by setting $d_{f}(x,y)=f(d(x,y))$. Then $(\mathbb{U}_{\Delta},d_{f})$ is isometric to $\mathbb{U}_{\Lambda}$, since $(\mathbb{U}_{\Delta},d_{f})$ is a Fraïssé structure and its age is exactly the class of all finite $\Lambda$-metric spaces. Now the identity map from $(\mathbb{U}_{\Delta},d)$ to $(\mathbb{U}_{\Delta},d_{f})$ induces a topological isomorphism from $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$ to $\mbox{\rm Iso}(\mathbb{U}_{\Lambda})$. ∎
[…]
Theorem 4.3.
Let $\Delta$ and $\Lambda$ be countable distance value sets. Then the following are equivalent:
  1. (i)
    ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ and ${\rm Aut}(\mathbb{U}_{\Lambda}^{<})$ are isomorphic as topological groups.
  2. (ii)
    $\Delta$ and $\Lambda$ are equivalent.
The rest of this subsection is devoted to a proof of Theorem 4.3. The outline of the proof follows the proof of Theorem 4.1 in [5].
The argument for the implication of (i) from (ii) is similar to the proof of Theorem 4.2.
To prove the other direction, we need some results due to Slutsky [33]. We work with the following concept from [33] in the Fraïssé theory of finite metric spaces.
Definition 4.2.
Let $\mathcal{A},\mathcal{B},\mathcal{C}$ be finite metric spaces. For simplicity assume $\mathcal{A}$ is a substructure of both $\mathcal{B}$ and $\mathcal{C}$. Also assume $|\mathcal{B}|\cap|\mathcal{C}|=|\mathcal{A}|$. Define a finite metric space $\mathcal{D}$ as follows. Let $|\mathcal{D}|$ = $|\mathcal{B}|\cup|\mathcal{C}|$. For $x,y\in|\mathcal{D}|$, define $d_{\mathcal{D}}(x,y)$ as follows. If $x,y\in|\mathcal{B}|$, then let $d_{\mathcal{D}}(x,y)=d_{\mathcal{B}}(x,y)$. Similarly, if $x,y\in\mathcal{C}$, then let $d_{\mathcal{D}}(x,y)=d_{\mathcal{C}}(x,y)$. If $x\in|\mathcal{B}|$ and $y\in|\mathcal{C}|$, then let
$d_{\mathcal{D}}(x,y)=\left\{\begin{array}[]{ll}\min\{d_{\mathcal{B}}(x,z)+d_{\mathcal{C}}(z,y)\colon z\in|\mathcal{A}|\},&\mbox{ if $|\mathcal{A}|\neq\varnothing$,}\\ \mbox{diam}(\mathcal{B})+\mbox{diam}(\mathcal{C}),&\mbox{ if $|\mathcal{A}|=\varnothing$.}\end{array}\right$.
Then $\mathcal{D}$ is called the free amalgam of $\mathcal{B}$ and $\mathcal{C}$ over $\mathcal{A}$.
Since (ordered) metric spaces can be regarded as relational structures, for an (ordered) metric space $\mathcal{B}$, a substructure $\mathcal{A}$ of $\mathcal{B}$ generated by a set $A\subseteq|\mathcal{B}|$ satisfies $|\mathcal{A}|=A$. Thus we will identify $\mathcal{A}$ with $A$, and speak of $A$ as a substructure. This does not cause confusion, and will simplify our notation.
Definition 4.3.
Let $A$ be a metric space and $p$ a partial isometry on $A$. A point $x\in{\rm dom}(p)$ is called periodic if there is a natural number $n>0$ such that
$x,p(x),\dots,p^{n}(x)\in{\rm dom}(p)\ {\rm and}\ p^{n}(x)=x.$
The set of all periodic points of $p$ is denoted by $Z(p)$. A point $x\in{\rm dom}(p)$ is called a fixed point if $p(x)=x$. The set of all fixed points of $p$ is denoted by $F(p)$.
Lemma 4.4 ([33]).
Let $D$ be a finite ordered metric space. Let $p_{1}$ and $p_{2}$ be order-preserving partial isometries of $D$ such that $Z(p_{1})=F(p_{1})$ and $Z(p_{2})=F(p_{2})$. Suppose $F(p_{1})\cap F(p_{2})\neq\emptyset$. Then there exist a finite ordered metric space $D^{\prime}$ extending $D$, order-preserving partial isometries $q_{1}$ and $q_{2}$ of $D^{\prime}$ extending $p_{1},p_{2}$ respectively, and an element $w\in\mathsf{F}(s,t)$, the free group generated by s,t{s,t}, such that
  1. (1)
    $Z(q_{1})=Z(p_{1})$ and $Z(q_{2})=Z(p_{2})$;
  2. (2)
    $D\subseteq{\rm dom}(q_{1})$;
  3. (3)
    ${\rm dom}(q_{1})\cup w(q_{1},q_{2})(D)$ is the free amalgam of ${\rm dom}(q_{1})$ and $w(q_{1},q_{2})(D)$ over $F(p_{1})\cap F(p_{2})$.
Moreover, the distances in $D^{\prime}$ are taken from the additive semigroup generated by the distances in $D$.
This result was originally proved for metric spaces as Theorem 4.12 of [33]. Nevertheless, as Slutsky remarked (Remark 4.13 of [33]), the above ordered version also holds.
We use the following notation in our further discussions. For a countable structure $\mathcal{A}$ and subset $B\subseteq|\mathcal{A}|$, we use ${\rm Aut}_{B}(\mathcal{A})$ to denote the subgroup of $\mbox{\rm Aut}(\mathcal{A})$ consisting of the automorphisms of $\mathcal{A}$ which fix each element of $B$. Note that if $B$ is finite, then ${\rm Aut}_{B}(\mathcal{A})$ is an open subgroup of $\mbox{\rm Aut}(\mathcal{A})$.
The following two lemmas are key technical results in our proof. The combination of their unordered versions was a theorem of Slutsky [33].
Lemma 4.5.
Let $\Delta$ be a countable distance value set, and let $A,B\subseteq\mathbb{U}_{\Delta}^{<}$ be finite subsets. Suppose $A\cap B\neq\varnothing$. Then
$\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle\ is\ dense\ in\ {\rm Aut}_{A\cap B}(\mathbb{U}_{\Delta}^{<}).$
Proof.
For notational simplicity, we assume $A\cap B=\{x_{0}\}$. Suppose $A=\{a_{0},\dots,a_{n}\}$ and $B=\{b_{0},\dots,b_{m}\}$. Again, for definiteness and notational simplicity, assume $m=n$ and $a_{0}=b_{0}=x_{0}<a_{1}<b_{1}<\dots<a_{n}<b_{n}$. Note that in general, $A\cap B=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{0}<x_{1}<\cdots<x_{k}$ and the rest of the points in $A,B$ can be in the gaps between $x_{i}$s including below $x_{0}$ and above $x_{k}$. For the general case, the argument is similar.
Let $C$ be a finite subset of $\mathbb{U}_{\Delta}^{<}$ and let $p\in{\rm Aut}_{A\cap B}(\mathbb{U}_{\Delta}^{<})$. It suffices to show that there is a $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ such that $p\upharpoonright C=q\upharpoonright C$. Without loss of generality, we may assume $A\subseteq C$ and $B\subseteq C$. Let $D=C\cup p(C)$. We find (to be explicitly defined later)
$q=w(q_{1},q_{2})^{-1}g_{1}^{-1}\dots g_{2n}^{-1}rg_{2n}\dots g_{1}w(q_{1},q_{2}),$
where $g_{2i-1}\in{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})$ and $g_{2i}\in{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})$. Roughly speaking, we start from an isometric copy of $D$ in $\mathbb{U}_{\Delta}^{<}$ where all the new points (other than $A\cap B$) are appropriately distributed between $x_{i}$s. Using the combination of $g_{i}$s, we propagate the points in the initial copy of $D$ to create the correct ordering of points in the final copy of $D$.
As a substructure of $\mathbb{U}^{<}_{\Delta}$, $D$ is a finite ordered metric space. Define two order-preserving partial isometries $p_{1},p_{2}$ of $D$ by letting $p_{1}(x)=x$ for all $x\in A$ and $p_{2}(x)=x$ for all $x\in B$. Now applying Lemma 4.4 to $D$ and $p_{1},p_{2}$, we get a finite ordered metric space $D^{\prime}$ extending $D$, order-preserving partial isometries $q_{1},q_{2}$ of $D^{\prime}$ which are extensions of $p_{1},p_{2}$, respectively, and a word $w\in\mathsf{F}(s,t)$ such that
  1. (1)
    $Z(q_{1})=Z(p_{1})=F(p_{1})=A$ and $Z(q_{2})=Z(p_{2})=F(p_{2})=B$;
  2. (2)
    $D\subseteq{\rm dom}(q_{1})$;
  3. (3)
    ${\rm dom}(q_{1})\cup w(q_{1},q_{2})(D)$ is the free amalgam of $\mbox{dom}(q_{1})$ and $w(q_{1},q_{2})(D)$ over $\{x_{0}\}$.
For notational simplicity we write $w(q_{1},q_{2})$ as $w$.
We note that $(A\cup B)\cap w(D)=\{x_{0}\}$. This is because $x_{0}\in A\cup B\subseteq C\subseteq D\subseteq\mbox{dom}(q_{1})$ and by (3), $\mbox{dom}(q_{1})\cup w(D)$ is the free amalgam of $\mbox{dom}(q_{1})$ and $w(D)$ over $\{x_{0}\}$.
Let $w_{1}(D)$ be an isomorphic copy of $w(D)$ so that $w_{1}(D)$ and $w(D)$ have exactly one common point $x_{0}$. For each $y\in D$, the point in $w_{1}(D)$ corresponding to $w(y)\in w(D)$ will be denoted by $w_{1}(y)$. Let $D_{1}^{\prime}$ be the free amalgam of $D^{\prime}$ and $w_{1}(D)$ over $\{x_{0}\}$. Figure 1 illustrates the containment relationship among the sets considered so far.
[…]
Figure 1. The composition of $D_{1}^{\prime}$.
Extend the ordering on $D^{\prime}$ to $D_{1}^{\prime}$ so that the following hold:
  1. (a)
    for any $y_{1},y_{2}\in D$, $w_{1}(y_{1})<w_{1}(y_{2})$ if and only if $w(y_{1})<w(y_{2})$;
  2. (b)
    for any $u\in A\cup B$ and $y\in D$, if $w(y)<b_{n}=\max(A\cup B)$, then $u<w_{1}(y)$ if and only if $u<w(y)$;
  3. (c)
    for any $y\in D$, if $w(y)>b_{n}$, then $a_{n}<w_{1}(y)<b_{n}$.
Such an extension clearly exists.
Now define $g_{1}\colon A\cup w(D)\to A\cup w_{1}(D)$ by letting $g_{1}(a)=a$ for all $a\in A$ and $g_{1}(w(y))=w_{1}(y)$ for all $y\in D$. Then $g_{1}$ is order-preserving by the above requirements (a) through (c). To see that $g_{1}$ is a partial isometry, let $a\in A$ and $y\in D$. Since $\mbox{dom}(q_{1})\cup w(D)$ is the free amalgam of $\mbox{dom}(q_{1})$ and $w(D)$ over $\{x_{0}\}$ and $D_{1}^{\prime}$ is the free amalgam of $D^{\prime}$ and $w_{1}(D)$ over $\{x_{0}\}$, we have
$d(a,w(y))=d(a,x_{0})+d(x_{0},w(y))=d(a,x_{0})+d(x_{0},w_{1}(y))=d(a,w_{1}(y)).$
Intuitively, $g_{1}$ sends $w(y)$ to a copy which is smaller than $b_{n}$ and fixes $A$ pointwise. Consequently, we have $w_{1}(D)<b_{n}$ and, for any $y\in D$, $w_{1}(y)<x_{0}$ if and only if $w(y)<x_{0}$.
Let $w_{2}(D)$ be a fresh isomorphic copy of $w(D)$ such that $w_{2}(D)$ and $w(D)$ has exactly one common point $x_{0}$. Let $D_{2}^{\prime}$ be the free amalgam of $D_{1}^{\prime}$ and $w_{2}(D)$ over $\{x_{0}\}$. Extend the ordering on $D_{1}^{\prime}$ to $D_{2}^{\prime}$ so that the following hold:
  1. (d)
    for any $y_{1},y_{2}\in D$, $w_{2}(y_{1})<w_{2}(y_{2})$ if and only if $w_{1}(y_{1})<w_{1}(y_{2})$;
  2. (e)
    for any $u\in A\cup B$ and $y\in D$, if $w(y)<a_{n}$, then $u<w_{2}(y)$ if and only if $u<w_{1}(y)$;
  3. (f)
    for any $y\in D$, if $w(y)>a_{n}$, then $b_{n-1}<w_{2}(y)<a_{n}$.
Define an order-preserving partial isometry $g_{2}\colon B\cup w_{1}(D)\to B\cup w_{2}(D)$ by letting $g_{2}(b)=b$ for all $b\in B$ and $g_{2}(w_{1}(y))=w_{2}(y)$ for all $y\in D$. Moreover, $w_{2}(D)<a_{n}$ and, for any $y\in D$, $w_{2}(y)<x_{0}$ if and only if $w_{1}(y)<x_{0}$.
Repeating this construction, we obtain, for $1\leq i\leq 2n$, $D_{i}^{\prime}=D^{\prime}\cup w_{1}(D)\cup\dots\cup w_{i}(D)$, which can be viewed as the free amalgam of $w_{i}(D)$ and the remaining part over $\{x_{0}\}$, an ordering on $D_{2n}^{\prime}$ extending that on $D^{\prime}$, and order-preserving partial isometries $g_{1},g_{2},g_{3},\dots,g_{2n}$ such that
  1. (i)
    for any $1\leq i\leq n$, $g_{2i-1}(a)=a$ for all $a\in A$, and $g_{2i}(b)=b$ for all $b\in B$;
  2. (ii)
    for any $1\leq i\leq n$ and $y\in D$, $g_{2i}(w_{2i-1}(y))=w_{2i}(y)$, and $g_{2i-1}(w_{2i-2}(y))=w_{2i-1}(y)$ if $i>1$.
Moreover, we have $w_{2n}(D)<a_{1}$ and, for any $y\in D$, $w_{2n}(y)<x_{0}$ if and only if $w(y)<x_{0}$.
Define a map
$r\colon A\cup w_{2n}(C)\to A\cup w_{2n}(p(C))$
by letting $r(a)=a$ for all $a\in A$ and $r(w_{2n}(c))=w_{2n}(p(c))$ for all $c\in C$. Then for $a\in A$ and $c\in C$, we have
$d(a,w_{2n}(c))$ $=d(a,x_{0})+d(x_{0},w_{2n}(c))=d(a,x_{0})+d(x_{0},w(c))$
$=d(a,x_{0})+d(x_{0},c)=d(a,x_{0})+d(x_{0},p(c))$
$=d(a,x_{0})+d(x_{0},w(p(c)))$
$=d(a,x_{0})+d(x_{0},w_{2n}(p(c)))=d(a,w_{2n}(p(c)))$
$=d(r(a),r(w_{2n}(c))).$
Thus $r$ is a partial isometry. Note that $w$ and $p$ are order-preserving and each fixes $x_{0}$. Thus for any $c\in C$, we have
$w_{2n}(c)<x_{0}$ $\iff w(c)<x_{0}\iff c<x_{0}$
$\iff p(c)<x_{0}\iff w(p(c))<x_{0}$
$\iff w_{2n}(p(c))<x_{0}.$
Since for all $y\in D$, $w_{2n}(y)<a_{1}$, we conclude that $r$ is an order-preserving partial isometry.
By Lemma 4.4 and our construction, the distances in $D_{2n}^{\prime}$ are all taken from the additive semigroup generated by the distances in $D$. Now redefine the metric on $D_{2n}^{\prime}$ by taking the minimum of the original distance value and $\sup\Delta$. Call the resulting metric space $D^{\prime\prime}$. Since $D$ is a $\Delta$-metric space, it follows that $D^{\prime\prime}$ is a $\Delta$-metric space, and, as an ordered metric space, $D^{\prime\prime}$ is an extension of $D$. Moreover, each of the order-preserving partial isometries $q_{1},q_{2},g_{1},g_{2},\dots,g_{2n},r$ continues to be order-preserving partial isometries of $D^{\prime\prime}$.
Now by the weak homogeneity of $\mathbb{U}_{\Delta}^{<}$, we can extend $D$ to an isomorphic copy of $D^{\prime\prime}$ as a substructure of $\mathbb{U}_{\Delta}^{<}$. By the ultrahomogeneity of $\mathbb{U}_{\Delta}^{<}$, we can extend all the above order-preserving partial isometries to full automorphisms of $\mathbb{U}_{\Delta}^{<}$. Denote the resulting automorphisms by the same notation. Then for $1\leq i\leq 2n$, we have $r,q_{1},g_{i}\in{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<})$ if $i$ is odd, and $q_{2},g_{i}\in{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})$ if $i$ is even. Finally, for any $c\in C$,
$p(c)=w(q_{1},q_{2})^{-1}g_{1}^{-1}\dots g_{2n}^{-1}rg_{2n}\dots g_{1}w(q_{1},q_{2})(c).$
Thus if we let
$q=w(q_{1},q_{2})^{-1}g_{1}^{-1}\dots g_{2n}^{-1}rg_{2n}\dots g_{1}w(q_{1},q_{2}),$
then $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ and $p\upharpoonright C=q\upharpoonright C$ as required. ∎
Lemma 4.6.
Let $\Delta$ be a countable distance value set, and let $A,B\subseteq\mathbb{U}_{\Delta}^{<}$ be finite subsets. Suppose $A\cap B=\varnothing$. Then
$\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle\ is\ dense\ in\ {\rm Aut}(\mathbb{U}_{\Delta}^{<}).$
Proof.
Let $p\in{\rm Aut}(\mathbb{U}_{\Delta}^{<})$ and let $C\subseteq\mathbb{U}_{\Delta}^{<}$ be a finite subset. We show that there is a $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ such that $p\upharpoonright C=q\upharpoonright C$. Again we may assume $A\cup B\subseteq C$. Consider the ordered $\Delta$-metric space $C\cup p(C)$. Extend it to an ordered $\Delta$-metric space $C\cup p(C)\cup\{x_{0}\}$, where the metric is defined by
$d(x_{0},y)=\mbox{diam}(C\cup p(C))=\max\{d(y_{1},y_{2})\colon y_{1},y_{2}\in C\cup p(C)\}$
for any $y\in C\cup p(C)$, and the ordering is defined by $x_{0}<y$ for all $y\in C\cup p(C)$. Such an extension can be found in $\mathbb{U}_{\Delta}^{<}$ by the weak homogeneity of $\mathbb{U}^{<}_{\Delta}$, and so we may assume $x_{0}\in\mathbb{U}_{\Delta}^{<}$. Now consider $A^{\prime}=A\cup\{x_{0}\}$, $B^{\prime}=B\cup\{x_{0}\}$ and let $p^{\prime}\in{\rm Aut}_{x_{0}}(\mathbb{U}_{\Delta}^{<})$ be an extention of $p\upharpoonright C\cup\{(x_{0},x_{0})\}$. Applying Lemma 4.5 to $A^{\prime},B^{\prime}$ and $p^{\prime}$, we get a $q\in\langle{\rm Aut}_{A^{\prime}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B^{\prime}}(\mathbb{U}_{\Delta}^{<})\rangle$ such that $q\upharpoonright C\cup\{x_{0}\}=p^{\prime}\upharpoonright C\cup\{x_{0}\}$. Then in particular $q\in\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{B}(\mathbb{U}_{\Delta}^{<})\rangle$ and $q\upharpoonright C=p\upharpoonright C$ as required. ∎
[…]
Lemma 4.7.
Let $V$ be a proper open subgroup of ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. Then there exists an element $x\in\mathbb{U}_{\Delta}^{<}$ such that $V\cdot x$ is finite.
Proof.
Toward a contradiction, assume $V\cdot x$ is infinite for all $x\in\mathbb{U}_{\Delta}^{<}$. We prove that $V$ is dense. Since $V$ is also closed, it follows that $V={\rm Aut}(\mathbb{U}_{\Delta}^{<})$, contradicting the properness of $V$.
[…]
By Lemma 4.6, $\langle{\rm Aut}_{A}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{g(A)}(\mathbb{U}_{\Delta}^{<})\rangle$ is dense in ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. This group is contained in $V$, and thus $V$ is dense. ∎
Lemma 4.8.
Let $V$ be a proper open subgroup of ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$. The following are equivalent:
  1. (1)
    There exists a unique $x\in\mathbb{U}_{\Delta}^{<}$ such that $V={\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})$.
  2. (2)
    $V$ is maximal among proper closed subgroups of ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$.
Proof.
[…]
For the uniqueness, assume ${\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})={\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<})$ but $x\neq y$. Then by Lemma 4.6, $\langle{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<})\rangle$ is dense. But
$\langle{\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<}),{\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<})\rangle={\rm Aut}_{\{x\}}(\mathbb{U}_{\Delta}^{<})={\rm Aut}_{\{y\}}(\mathbb{U}_{\Delta}^{<}),$
and thus they are both the whole group, which is impossible. ∎
We are now ready to finish the proof of Theorem 4.3. Recall that the remaining direction is (i) implying (ii).
[…]
$d(x,y)=d(x^{\prime},y^{\prime})\iff d(g(x),g(y))=d(g(x^{\prime}),g(y^{\prime})).$
Now for any $d\in\Delta$, we define $\theta(d)\in\Lambda$ by finding $x,y\in\mathbb{U}_{\Delta}^{<}$ such that $d(x,y)=d$, and then setting $\theta(d)=d(g(x),g(y))$. By the above observation, $\theta$ is well-defined and bijective. We show that $\theta$ witnesses the equivalence of $\Delta$ and $\Lambda$. Suppose $(d_{1},d_{2},d_{3})$ is a $\Delta$-triangle. There exist $x,y,z\in\mathbb{U}_{\Delta}^{<}$ such that the distances among them are exactly $(d_{1},d_{2},d_{3})$. Then the distaces among $g(x),g(y),g(z)$ are $(\theta(d_{1}),\theta(d_{2}),\theta(d_{3}))$. Thus $(\theta(d_{1}),\theta(d_{2}),\theta(d_{3}))$ is a $\Lambda$-triangle. The converse also holds.
The proof of Theorem 4.3 is complete.
5. The Complexity of the Equivalence between Distance Value Sets
In this section we study the complexity of the equivalence (see Definition 4.1) between countable distance value sets from the point of view of descriptive set theory. We first establish some notation.
Definition 5.1.
Let $\mathsf{CDVS}\subseteq\mathbb{R}^{\mathbb{N}}$ consist of all sequences $(d_{i})_{i\in\mathbb{N}}$ of real numbers such that
  1. (a)
    $d_{i}\geq 0$ for all $i$; the set $\{i\in\mathbb{N}\colon d_{i}=0\}$ is infinite; $d_{i}>0$ for some $i$;
  2. (b)
    if $d_{i},d_{j}>0$, then $d_{i}=d_{j}$ if and only if $i=j$;
  3. (c)
    if $\sup_{k}d_{k}<\infty$, then $\sup_{k}d_{k}=d_{i}$ for some $i$;
  4. (d)
    for any $i,j\geq 0$, if $d_{i}+d_{j}<\sup_{k}d_{k}$ then there is $\ell$ such that $d_{\ell}=d_{i}+d_{j}$.
Intuitively, $\mathsf{CDVS}$ consists of codes of all countable distance value sets, which are enumerated as sequences of real numbers without repetitions. We include zero entries in the codes to accommodate finite distance value sets. $\mathsf{CDVS}$ is a Borel subset of $\mathbb{R}^{\mathbb{N}}$, and hence it is a standard Borel space. We also use the following notation:
$\begin{array}[]{lcl}\mathsf{CDVS}_{0}&=&\left\{(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}\colon\inf\{d_{i}>0\colon i\in\mathbb{N}\}=0\right\},\\ \mathsf{CDVS}_{+}&=&\left\{(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}\setminus\mathsf{CDVS}_{0}\colon\sup_{i}d_{i}<\infty\right\},\\ \mathsf{CDVS}_{\infty}&=&\left\{(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}\setminus\mathsf{CDVS}_{0}\colon\sup_{i}d_{i}=\infty\right\}.\end{array}$
These sets form a partition of $\mathsf{CDVS}$ into Borel subsets, and in particular they are all standard Borel spaces. $\mathsf{CDVS}_{0}$ consists of codes of all countable distance value sets $\Delta$ with $\inf\Delta=0$. Similarly, $\mathsf{CDVS}_{+}$ codes countable distance value sets $\Delta$ with $\inf\Delta>0$ and $\sup\Delta<\infty$, and $\mathsf{CDVS}_{\infty}$ codes those $\Delta$ with $\inf\Delta>0$ and $\sup\Delta=\infty$. In our discussions below, we will not distinguish a countable distance value set $\Delta$ from a code of $\Delta$ as an element of $\mathsf{CDVS}$.
Definition 5.2.
Let $\approx$ be the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined by
$\begin{array}[]{rcl}(c_{i})_{i}\approx(d_{i})_{i}&\iff&\exists g\in S_{\infty}\ [\forall i\ (c_{i}=0\longleftrightarrow d_{g(i)}=0)\mbox{ and }\\ &&\forall i,j,k\ (|c_{j}-c_{k}|\leq c_{i}\leq c_{j}+c_{k}\longleftrightarrow\\ &&|d_{g(j)}-d_{g(k)}|\leq d_{g(i)}\leq d_{g(j)}+d_{g(k)})].\end{array}$
For any $A\subseteq\mathbb{R}^{\mathbb{N}}$, let $\approx_{A}$ denote $\approx\upharpoonright(A\times A)$.
Thus $\approx_{\mathsf{CDVS}}$ is the equivalence between countable distance value sets. Note that $\mathsf{CDVS}_{0}$, $\mathsf{CDVS}_{+}$ and $\mathsf{CDVS}_{\infty}$ are all invariant under the equivalence relation $\approx$.
In the rest of this section we will determine the exact complexity of $\approx_{\mathsf{CDVS}_{+}}$, $\approx_{\mathsf{CDVS}_{\infty}}$ and $\approx_{\mathsf{CDVS}}$ and give some upper and lower bounds for $\approx_{\mathsf{CDVS}_{0}}$ in the Borel reducibility hierarchy.
5.1. The exact complexity of $\approx_{\mathsf{CDVS}_{+}}$ and $\approx_{\mathsf{CDVS}}$
Lemma 5.1.
$\approx_{\mathsf{CDVS}}$ is Borel reducible to an isomorphism relation of countable structures. Consequently, $\approx_{\mathsf{CDVS}}$ is Borel reducible to $E^{\infty}_{S_{\infty}}$.
Proof.
To each $\Delta\in\mathsf{CDVS}$ we associate a countable structure $\mathcal{A}_{\Delta}=(\Delta,R^{\mathcal{A}})$, where $R^{\mathcal{A}}$ is a ternary relation defined as
$R^{\mathcal{A}}(d_{1},d_{2},d_{3})\iff\mbox{$(d_{1},d_{2},d_{3})$ is a $\Delta$-triangle.}$
It is clear that for $\Delta,\Lambda\in\mathsf{CDVS}$, we have $\Delta\approx\Lambda$ if and only if $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$. This gives a Borel reduction from $\approx_{\mathsf{CDVS}}$ to an isomorphism relation of countable structures. ∎
Theorem 5.2.
The isomorphism relation of all countable linear orders is Borel reducible to $\approx_{\mathsf{CDVS}_{+}}$. Consequently, $E^{\infty}_{S_{\infty}}$ is Borel reducible to $\approx_{\mathsf{CDVS}_{+}}$.
Proof.
For a countable linear order $\mathcal{O}$, let $\psi(\mathcal{O})\subseteq(3,4)$ be an isomorphic copy of $\mathcal{O}$, and let
$\varphi(\mathcal{O})=\{2,4,6\}\cup\psi(\mathcal{O})\cup(2+\psi(\mathcal{O})).$
Then $\varphi(\mathcal{O})\in\mathsf{CDVS}_{+}$. We verify that $\varphi$ is a Borel reduction from the isomorphism relation of all countable linear orders to $\approx_{\mathsf{CDVS}_{+}}$.
First suppose $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are isomorphic countable linear orders. Then there is an order-preserving bijection $f$ between $\psi(\mathcal{O}_{1})$ and $\psi(\mathcal{O}_{2})$. For $x\in\varphi(\mathcal{O}_{1})$, let
$F(x)=\left\{\begin{array}[]{ll}f(x),&\mbox{ if $x\in\psi(\mathcal{O}_{1})$,}\\ f(x-2)+2,&\mbox{ if $x\in 2+\psi(\mathcal{O}_{1})$,}\\ x,&\mbox{ otherwise.}\end{array}\right$.
It is straightforward to check that $F$ is a bijection between $\varphi(\mathcal{O}_{1})$ and $\varphi(\mathcal{O}_{2})$ witnessing their equivalence. Thus $\varphi(\mathcal{O}_{1})\approx\varphi(\mathcal{O}_{2})$.
Conversely, suppose $\varphi(\mathcal{O}_{1})\approx\varphi(\mathcal{O}_{2})$ and this equivalence is witnessed by a bijection $g$. Let $\Delta=\varphi(\mathcal{O}_{1})$ and $\Lambda=\varphi(\mathcal{O}_{2})$, and let $\mathcal{A}_{\Delta}=(\Delta,R^{\Delta})$ and $\mathcal{A}_{\Lambda}=(\Lambda,R^{\Lambda})$ be the countable structures defined in the proof of Lemma 5.1. Then $g$ is an isomorphism witnessing $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$.
Now note that $2$ is definable in both $\mathcal{A}_{\Delta}$ and $\mathcal{A}_{\Lambda}$ by the formula $\exists y\neg R(x,x,y)$. Thus $g(2)=2$. Moreover, $\psi(\mathcal{O}_{1})$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$x\neq 2\wedge R(2,2,x)\wedge\exists y\neg R(2,x,y).$
Similarly, $\psi(\mathcal{O}_{2})$ is definable in $\mathcal{A}_{\Lambda}$ by the same formula. Thus $g$ sends $\psi(\mathcal{O}_{1})$ to $\psi(\mathcal{O}_{2})$. Finally, on $\psi(\mathcal{O}_{1})$, the $x<y$ relation is definable by the formula
$\exists z\,[\neg R(2,x,z)\wedge R(2,y,z)].$
Similarly for $\psi(\mathcal{O}_{2})$. Hence $g$ preserves the order << as a bijection between $\psi(\mathcal{O}_{1})$ and $\psi(\mathcal{O}_{2})$. It follows that the linear orders $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are isomorphic.
It is routine to check that $\varphi$ is Borel. ∎
Corollary 5.3.
Both $\approx_{\mathsf{CDVS}_{+}}$ and $\approx_{\mathsf{CDVS}}$ are Borel bireducible with $E^{\infty}_{S_{\infty}}$. In particular, they are analytic and non-Borel.
Proof.
This follows immediately from Lemma 5.1 and Theorem 5.2, in view of the trivial fact that $\approx_{\mathsf{CDVS}_{+}}$ is Borel reducible to $\approx_{\mathsf{CDVS}}$ via the identity function. ∎
5.2. The exact complexity of $\approx_{\mathsf{CDVS}_{\infty}}$
Theorem 5.4.
Let $\Delta$ and $\Lambda$ be countable distance value sets with $\inf\Delta=\inf\Lambda=1$ and $\sup\Delta=\sup\Lambda=\infty$. Then $\Delta$ and $\Lambda$ are equivalent if and only if $\Delta=\Lambda$.
Proof.
The implication $\Leftarrow$ is obvious. We only show the $\Rightarrow$ direction. For this, let $f$ be a bijection between $\Delta$ and $\Lambda$ witnessing their equivalence.
For a subset $I\subseteq\mathbb{R}$, let $\Delta_{I}=\Delta\cap I$ and $\Lambda_{I}=\Lambda\cap I$. Let $\mathcal{A}_{\Delta}=(\Delta,R^{\Delta})$ and $\mathcal{A}_{\Lambda}=(\Lambda,R^{\Lambda})$ be the countable structures defined in the proof of Lemma 5.1. Then $f$ is an isomorphism witnessing $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$.
First note that $\inf\Delta=1\in\Delta$ if and only if $\mathcal{A}_{\Delta}$ satisfies the sentence
$\exists x\,\forall y\,[R(x,x,y)\rightarrow\forall z\,R(z,z,y)].$
Similarly for $\Lambda$. Thus $1\in\Delta$ if and only if $1\in\Lambda$. We consider two cases.
Case 1: $1\in\Delta\cap\Lambda$. Since $\Delta$ and $\Lambda$ are distance value sets with $\sup\Delta=\inf\Delta=\infty$, we have that $k\in\Delta\cap\Lambda$ for all positive integers $k$. Note that $1$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{1}(x)=\forall y\,[R(x,x,y)\rightarrow\forall z\,R(z,z,y)].$
Similarly for $\mathcal{A}_{\Lambda}$. Thus $f(1)=1$. Also, we have that $\Delta_{[1,2]}$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{[1,2]}(x)=\forall y\,R(y,y,x).$
Similarly for $\Lambda_{[1,2]}$. Hence $f$ sends $\Delta_{[1,2]}$ to $\Lambda_{[1,2]}$.
Now for $k\geq 2$, define inductively
$\chi_{[1,k+1]}(x)=\exists y\,[\chi_{[1,k]}(y)\wedge R(1,y,x)].$
Then $\Delta_{[1,k+1]}$ is definable in $\mathcal{A}_{\Delta}$ by $\chi_{[1,k+1]}$, and similarly for $\Lambda_{[1,k+1]}$. It follows that $f$ sends $\Delta_{[1,k+1]}$ to $\Lambda_{[1,k+1]}$ for all $k\geq 1$.
Next we define, by induction on $k\in\mathbb{N}^{+}$ and $m\in\mathbb{N}$, a formula $\chi_{[1,1+\frac{k}{2^{m}}]}$ as follows.
$\begin{array}[]{rcl}\chi_{[1,1+\frac{k}{2^{0}}]}(x)&=&\chi_{[1,k+1]}(x)\mbox{ for all $k\geq 1$,}\\ \chi_{[1,1+\frac{k}{2^{m+1}}]}(x)&=&\forall y\,[R(x,x,y)\rightarrow\chi_{[1,1+\frac{k+2^{m}}{2^{m}}]}(y)].\end{array}$
For any $k\in\mathbb{N}^{+}$ and $m\in\mathbb{N}$, the formula $\chi_{[1,1+\frac{k}{2^{m}}]}$ defines $\Delta_{[1,1+\frac{k}{2^{m}}]}$ in $\mathcal{A}_{\Delta}$ and $\Lambda_{[1,1+\frac{k}{2^{m}}]}$ in $\mathcal{A}_{\Lambda}$. Hence $f$ sends $\Delta_{[1,1+\frac{k}{2^{m}}]}$ to $\Lambda_{[1,1+\frac{k}{2^{m}}]}$. This implies that for any dyadic rationals $1<a<b$, $f$ sends $\Delta_{[a,b]}$ to $\Lambda_{[a,b]}$. Hence $f$ must be the identity function, and so $\Delta=\Lambda$.
Case 2: $1\not\in\Delta\cup\Lambda$. The formula $\chi_{[1,2]}$ now defines $\Delta_{(1,2]}$ in $\mathcal{A}_{\Delta}$ and $\Lambda_{(1,2]}$ in $\mathcal{A}_{\Lambda}$. Thus $f$ sends $\Delta_{(1,2]}$ to $\Lambda_{(1,2]}$.
Let
$\varphi_{2}(x)=\exists y\,[\chi_{[1,2]}(y)\wedge\forall z\,(\chi_{[1,2]}(z)\rightarrow R(x,y,z))]\wedge\neg\chi_{[1,2]}(x)$
and let $D_{2}$ be defined by $\varphi_{2}$ in $\mathcal{A}_{\Delta}$. Then $D_{2}\subseteq\Delta_{(2,3]}$. We claim that $D_{2}\neq\varnothing$. To see this, let $a,b\in\Delta_{(1,2]}$ be such that $a<b$ and 2(a1)<b12(a-1)<b-1. Such $a,b$ exist since $\inf\Delta=1\not\in\Delta$. Then $2a\in\Delta$ since $\Delta$ is a distance value set, and $\varphi_{2}[2a]$ holds in $\mathcal{A}_{\Delta}$ as witnessed by $b$. Thus $2a\in D_{2}$. A similar argument shows $\inf D_{2}=2$. It follows that $\Delta_{(2,3]}$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{(2,3]}(x)=\forall y\,\forall z\,[(\chi_{[1,2]}(y)\wedge\varphi_{2}(z))\rightarrow R(x,y,z)]\wedge\neg\chi_{[1,2]}(x)$
and $\Delta_{(1,3]}$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{(1,3]}(x)=\chi_{[1,2]}(x)\vee\chi_{(2,3]}(x).$
Similarly for $\mathcal{A}_{\Lambda}$. Therefore $f$ sends $\Delta_{(2,3]}$ to $\Lambda_{(2,3]}$ and sends $\Delta_{(1,3]}$ to $\Lambda_{(1,3]}$.
Now for $k\geq 2$, define inductively
$\begin{array}[]{rcl}\varphi_{k+1}(x)&=&\exists y\,[\varphi_{k}(y)\wedge\forall z\,(\chi_{[1,2]}(z)\rightarrow R(x,y,z))]\wedge\neg\chi_{(1,k+1]}(x),\\ \chi_{(k+1,k+2]}(x)&=&\forall y\,\forall z\,[(\chi_{[1,2]}(y)\wedge\varphi_{k+1}(z))\rightarrow R(x,y,z)]\wedge\neg\chi_{(1,k+1]}(x),\\ \chi_{(1,k+2]}(x)&=&\chi_{(1,k+1]}(x)\vee\chi_{(k+1,k+2]}(x).\end{array}$
Then for all $k\geq 1$, $\chi_{(1,k+2]}$ defines $\Delta_{(1,k+2]}$ in $\mathcal{A}_{\Delta}$ and $\Lambda_{(1,k+2]}$ in $\mathcal{A}_{\Lambda}$. Hence for all $k\geq 2$, $f$ sends $\Delta_{(1,k+1]}$ to $\Lambda_{(1,k+1]}$.
The rest of the proof is similar to Case 1. We conclude that $f$ must be the identity function, and so $\Delta=\Lambda$. ∎
Theorem 5.5.
$\approx_{\mathsf{CDVS}_{\infty}}$ is Borel bireducible with $^{+}$.
Proof.
To each $\Delta\in\mathsf{CDVS}_{\infty}$ we associate a code for
$\Delta^{*}=\displaystyle\frac{\Delta}{\inf\Delta}.$
Then $\Delta^{*}\in\mathsf{CDVS}_{\infty}$ and $\inf\Delta^{*}=1$. By Theorem 5.4, the map $\Delta\mapsto\Delta^{*}$ is a Borel function witnessing that $\approx_{\mathsf{CDVS}_{\infty}}\,\leq_{B}\ =^{+}$.
Conversely, given any nonempty countable subset $A$ of the interval $(0,1)$, let $\Delta_{A}$ be the subsemigroup of $(\mathbb{R}^{+},+)$ generated by $\{1\}\cup(1+A)$. Then $\Delta_{A}$ is a countable distance value set with $\inf\Delta_{A}=1$ and $\sup\Delta_{A}=\infty$. Thus $\Delta_{A}\in\mathsf{CDVS}_{\infty}$. Moreover, $\Delta_{A}\cap(1,2)=1+A$. By Theorem 5.4, if $A$ and $B$ are two nonempty countable subsets of $(0,1)$, then $\Delta_{A}$ and $\Delta_{B}$ are equivalent if and only if $\Delta_{A}=\Delta_{B}$, if and only if $1+A=1+B$, and if and only if $A=B$. Thus $A\mapsto\Delta_{A}$ is a Borel function witnessing that $^{+}\,\leq_{B}\ \approx_{\mathsf{CDVS}_{\infty}}$. ∎
5.3. The complexity of $\approx_{\mathsf{CDVS}_{0}}$
In this subsection we give some upper and lower bounds for the complexity of $\approx_{\mathsf{CDVS}_{0}}$. We also discuss some related equivalence relations.
Theorem 5.6.
Let $\Delta$ and $\Lambda$ be countable distance value sets with $\inf\Delta=\inf\Lambda=0$. Then the following are equivalent:
  1. (1)
    $\Delta$ and $\Lambda$ are equivalent.
  2. (2)
    There is a positive real number $r$ such that $\Lambda=r\cdot\Delta$.
Moreover, if $\Delta$ and $\Lambda$ are equivalent, then every witness of their equivalence has the form of multiplication by a positive real number.
Proof.
[…]
Fix a sequence $(\delta_{i})_{i\in\mathbb{N}}$ in $\Delta$ such that $\delta_{i+1}<{\delta_{i}}/{2}$ for all $i$. Such a sequence exists since $\mathcal{A}_{\Delta}$ satisfies the sentence $\forall x\,\exists y\,\neg R(y,y,x)$. Since $f$ is an isomorphism, we have $f(\delta_{i+1})<{f(\delta_{i})}/{2}$ for all $i$.
We claim that $f$ is continuous as a function from $\Delta\subseteq\mathbb{R}$ to $\Lambda\subseteq\mathbb{R}$. To see this, fix $x\in\Delta$. For any $\epsilon>0$, find an $i$ such that $\delta_{i}<x$ and $f(\delta_{i})<\epsilon$. For any $y\in\Delta$ with $|x-y|<\delta_{i}$, $(x,y,\delta_{i})$ is a $\Delta$-triangle. Thus $(f(x),f(y),f(\delta_{i}))$ is a $\Lambda$-triangle. In particular, $|f(x)-f(y)|<f(\delta_{i})<\epsilon$. The actually proves that for all positive $\delta$, $f$ is uniformly continuous on $\Delta_{>\delta}\triangleq\Delta\cap(\delta,+\infty)$.
Next we show that $f$ is in fact linear. In the proof we use notation such as $\Delta_{>\delta}$, $\Delta_{<\delta}$, and $\Delta_{[a,b]}$ with obvious meanings.
To show that $f$ is linear, assume $x,y,x+y\in\Delta$. We claim that
  1. (a)
    $f(\Delta_{[|x-y|,x+y]})\subseteq\Lambda_{[|f(x)-f(y)|,f(x)+f(y)]}$;
  2. (b)
[…]
  3. (c)
[…]
Let $z\in\Delta$. (a) is equivalent to saying that if $(x,y,z)$ is a $\Delta$-triangle, then $(f(x),f(y),f(z))$ is a $\Lambda$-triangle, which is immediate from our assumption. Also immediate is that if $z>x+y$, then $(x,y,z)$ is not a $\Delta$-triangle. Thus $(f(x),f(y),f(z))$ is not a $\Lambda$-triangle, and so either $f(z)>f(x)+f(y)$ or f(z)<|f(x)f(y)|f(z)<|f(x)-f(y)|. Similarly for z<|xy|z<|x-y|. Now, to prove (b), assume toward a contradiction that there are $z_{1},z_{2}\in\Delta$ with $z_{1},z_{2}>x+y$ such that $f(z_{1})<|f(x)-f(y)|$ and $f(z_{2})>f(x)+f(y)$. Without loss of generality assume $z_{1}<z_{2}$ (the argument for the case $z_{1}>z_{2}$ is similar). By the continuity of $f$ and the density of both $\Delta$ and $\Lambda$, we may find $z_{1}=a_{1},\dots,a_{n}=z_{2}$ in $\Delta$ such that $a_{k}>x+y$ and $|f(a_{k+1})-f(a_{k})|<f(x)+f(y)-|f(x)-f(y)|$ for all $1\leq k\leq n$. Then there must be an $a_{k}$ with $|f(x)-f(y)|\leq f(a_{k})\leq f(x)+f(y)$, contradicting $a_{k}>x+y$. This proves (b). A similar argument for $z_{1},z_{2}<|x-y|$ proves (c).
Since $f$ is a bijection from $\Delta$ to $\Lambda$, the claimed inclusions are actually equalities. By considering sufficiently small $\delta_{i}$, we see that $f(\Delta_{<|x-y|})=\Lambda_{<|f(x)-f(y)|}$, and thus $f(\Delta_{>x+y})=\Lambda_{>f(x)+f(y)}$. By the continuty of $f$ again, we have $f(x+y)=f(x)+f(y)$.
Finally, we argue that for any $x,t\in\Delta$, we have $f(x)/x=f(t)/t$. This proves (2).
For each $k>0$, find $0<\delta_{k}<1/2^{k}$ by the continuity of $f$ such that for all $z\in\Delta$, $|t-z|<\delta_{k}$ implies |f(t)f(z)|<1/k|f(t)-f(z)|<1/k. Now for each $k$, find positive rational number $n_{k}/m_{k}$ such that $0<t-(n_{k}/m_{k})x<\delta_{k}/2$. Note that by the density of $\Delta$ and bijectivity of $f$, $\delta_{k}\to 0$ as $k\to\infty$. Thus $n_{k}/m_{k}\to t/x$ as $k\to\infty$.
Fix $k$. By the density of $\Delta$, find an increasing sequence $(t_{j})_{j\in\mathbb{N}}$ with $t_{j}\in\Delta$ and $t_{j}\to x/m_{k}$ as $j\to\infty$. Then $m_{k}t_{j}\to x$ as $j\to\infty$, and $m_{k}t_{j}<x\in\Delta$. Thus $m_{k}t_{j}\in\Delta$ and $m_{k}f(t_{j})=f(m_{k}t_{j})\to f(x)$, which gives $f(t_{j})\to f(x)/m_{k}$. On the other hand, $n_{k}t_{j}<(n_{k}/m_{k})x<t\in\Delta$ and $n_{k}t_{j}\to(n_{k}/m_{k})x$ as $j\to\infty$. Hence $n_{k}t_{j}\in\Delta$ and $f(n_{k}t_{j})=n_{k}f(t_{j})\to(n_{k}/m_{k})f(x)$ as $j\to\infty$. Thus we can find a $j$ such that $|n_{k}t_{j}-(n_{k}/m_{k})x|<\delta_{k}/2$ and $|f(n_{k}t_{j})-(n_{k}/m_{k})f(x)|<1/k$. For this fixed $j$, we have $|t-n_{k}t_{j}|\leq|t-(n_{k}/m_{k})x|+|n_{k}t_{j}-(n_{k}/m_{k})x|<\delta_{k}$. Thus $|f(t)-f(n_{k}t_{j})|<1/k$, and consequently $|f(t)-(n_{k}/m_{k})f(x)|<2/k$.
Letting $k\to\infty$, we get $f(t)=(t/x)f(x)$ and thus $f(x)/x=f(t)/t$. ∎
In view of Theorem 5.6, we define the following notion.
Definition 5.3.
Let $\sim$ be the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined by
$(c_{i})_{i}\sim(d_{i})_{i}\iff\exists g\in S_{\infty}\ \exists r\in(0,+\infty)\ \forall i\ d_{g(i)}=r\cdot c_{i}.$
For any $A\subseteq\mathbb{R}^{\mathbb{N}}$, let $\sim_{A}$ denote $\sim\upharpoonright(A\times A)$.
Thus Theorem 5.6 can be restated as $\sim_{\mathsf{CDVS}_{0}}\,=\ \approx_{\mathsf{CDVS}_{0}}$.
To obtain some upper and lower bounds for $\sim_{\mathsf{CDVS}_{0}}$ in the Borel reducibility hierarchy, we consider another equivalence relation defined and studied by Calderoni–Marker–Motto Ros–Shani [2]. Let $\mathsf{ArGp}$ be the space of all countable Archimedean ordered groups, and let $\cong_{\mathsf{ArGp}}$ denote the isomorphism relation on $\mathsf{ArGp}$. It was proved in [2] that
$^{+}\ <_{B}\ \cong_{\mathsf{ArGp}}\ \leq_{B}\ \cong^{*}_{3,1}.$
Here we prove the following facts.
Lemma 5.7.
$\cong_{\mathsf{ArGp}}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$.
Proof.
By a theorem of Hölder (see Theorem 2.1 of [2]), every element of $\mathsf{ArGp}$ is isomorphic to a subgroup of $\mathbb{R}$ with the natural order of real numbers. Thus we may regard $\mathsf{ArGp}$ as the space of all countable subgroups of $\mathbb{R}$ with the natural order. Note that if $G\in\mathsf{ArGp}$ and $\inf[G\cap(0,+\infty)]\neq 0$, then $G$ is isomorphic to $\mathbb{Z}$. Let $\mathsf{ArGp}_{0}$ be the space of all elements of $\mathsf{ArGp}$ which are not isomorphic to $\mathbb{Z}$. Then $\mathsf{ArGp}\setminus\mathsf{ArGp}_{0}$ consists of exactly one isomorphism class of $\mathsf{ArGp}$, and it follows easily that $\cong_{\mathsf{ArGp}}\ \leq_{B}\ \cong_{\mathsf{ArGp}_{0}}$.
For each $G\in\mathsf{ArGp}_{0}$, let $\Delta(G)=G\cap(0,+\infty)$. Then $\Delta(G)$ is a countable distance value set and $\inf\Delta(G)=0$. It is easy to see that $\Delta\colon\mathsf{ArGp}_{0}\to\mathsf{CDVS}_{0}$ is a Borel map.
To see that $\Delta$ is a Borel reduction witnessing $\cong_{\mathsf{ArGp}}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$, we use a result of Hion (see Lemma 2.2 of [2]) which states that for any $\in\mathsf{ArGp}, GHG\cong H$ if and only if there is a real number $r$ such that $r\cdot G=H$. It is then clear that $G\cong H$ if and only if there is a positive real number $r$ such that $r\cdot G=H$, if and only if there is a positive real number $r$ such that $r\cdot\Delta(G)=\Delta(H)$, and if and only if $\Delta(G)\sim\Delta(H)$. ∎
Theorem 5.8.
$\sim_{\mathsf{CDVS}_{0}}\ \leq_{B}\ \cong^{*}_{3,1}$.
Proof.
Similar to Proposition 3.12 of [2], using Corollary 6.4 of [14], this can be proved by showing that $\sim_{\mathsf{CDVS}_{0}}$ is Borel bireducible with an isomorphism relation and it is itself ${\bf\Sigma}^{0}_{4}$.
We first show that $\sim$ is ${\bf\Sigma}^{0}_{4}$. For this, only note that for any $(c_{i})_{i},(d_{i})_{i}\in\mathbb{R}^{\mathbb{N}}$,
$\begin{array}[]{rcl}(c_{i})_{i}\sim(d_{i})_{i}&\iff&\exists i,j\ \left(c_{i}/d_{j}>0\mbox{ and }\right.\\ &&\forall n\ \exists m\ \left(c_{n}=d_{m}c_{i}/d_{j}\right)\mbox{ and }\\ &&\left.\forall m\ \exists n\ \left(c_{n}=d_{m}c_{i}/d_{j}\right)\right).\end{array}$
Next we show that $\sim_{\mathsf{CDVS}_{0}}$ is Borel bireducible with an isomorphism relation. For this we define a language $\mathcal{L}$ and an $\mathcal{L}_{\omega_{1}\omega}$-theory $T$ so that the isomorphism relation on all countable models of $T$ is Borel bireducible with $\sim_{\mathsf{CDVS}_{0}}$.
Let
$\mathcal{L}=\{0,c,\leq,+\}\cup\{R_{q}\colon q\in\mathbb{Q}^{+}\}$
where
  • $0,c$ are constant symbols;
  • $\leq$ is a binary relation symbol;
  • ++ is a binary function symbol;
  • for each positive rational number $q\in\mathbb{Q}^{+}$, $R_{q}$ is a binary relation symbol.
In a countable distance value set $\Delta$ with $\inf\Delta=0$, we interpret $c$ to be 0 if $\sup\Delta=+\infty$ and to be $\sup\Delta$ otherwise. For each $q\in\mathbb{Q}^{+}$, $R_{q}$ is interpreted as $R_{q}(x,y)$ if and only if $q<x/y$. Under these interpretations, $\Delta\cup\{0\}$ becomes a countable $\mathcal{L}$-structure.
Let $T$ be an $\mathcal{L}_{\omega_{1}\omega}$-theory describing the following properties of a countable $\mathcal{L}$-structure $\mathcal{M}$ (for notational simplicity, we omit the superscripts $\mathcal{M}$ in the interpretations of the symbols of $\mathcal{L}$ in $\mathcal{M}$):
  1. (1)
    For all $q\in\mathbb{Q}^{+}$ and $x\in|\mathcal{M}|$, $\lnot R_{q}(x,0)$ and $\lnot R_{q}(0,x)$.
  2. (2)
    For all $x,y\in|\mathcal{M}|\setminus\{0\}$, the set of $q\in\mathbb{Q}^{+}$ with $R_{q}(x,y)$ is nonempty and forms a Dedekind cut of positive rational numbers, i.e., it is not the whole positive rational numbers, is closed downward and has no maximum. Moreover, the real number defined by the Dedekind cut $\{q\in\mathbb{Q}^{+}\colon R_{q}(x,x)\}$ is $1$.
  3. (3)
    For all $x,x^{\prime},y\in|\mathcal{M}|\setminus\{0\}$, if $x\neq x^{\prime}$, then the Dedekind cuts defined by $x,y$ and $x^{\prime},y$ are different.
  4. (4)
    For all $x,y,z\in|\mathcal{M}|\setminus\{0\}$ and $p,q\in\mathbb{Q}^{+}$, if $R_{p}(x,y)$ and $R_{q}(y,z)$, then $R_{pq}(x,z)$, and if $\lnot R_{p}(x,y)$ and $\lnot R_{q}(y,z)$, then $\lnot R_{pq}(x,z)$.
  5. (5)
    $\leq$ is a linear ordering of $|\mathcal{M}|$, 0 is the $\leq$-least element, and for all $^{\prime}\in|\mathcal{M}|\setminus\{0\}, xxx\leq x^{\prime}$ if and only if $x=x^{\prime}$ or $R_{1}(x^{\prime},x)$.
  6. (6)
    For all $x,x^{\prime},y\in|\mathcal{M}|\setminus\{0\}$ and $q\in\mathbb{Q}^{+}$, $R_{q}(x+x^{\prime},y)$ if and only if there are $q_{1},q_{2}\in\mathbb{Q}^{+}$ with $q_{1}+q_{2}=q$ such that $R_{q_{1}}(x,y)$ and $R_{q_{2}}(x^{\prime},y)$.
  7. (7)
    For all $x\in|\mathcal{M}|\setminus\{0\}$ and $q\in\mathbb{Q}^{+}$, there is $x^{\prime}\in|\mathcal{M}|\setminus\{0\}$ such that $\lnot R_{q}(x^{\prime},x)$.
Let $\mbox{Mod}(T)$ denote the Borel class of all countable models of $T$. Let $\cong_{T}$ denote the isomorphism relation on $\mbox{Mod}(T)$. It remains to verify that both $\sim_{\mathsf{CDVS}_{0}}\ \leq_{B}\ \cong_{T}$ and $\cong_{T}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$ hold.
Let $\Delta$ be a countable distance value set with $\inf\Delta=0$. We define a countable $\mathcal{L}$-structure $\mathcal{M}_{\Delta}$. Let $\mathbb{R}^{+}$ denote the semigroup of positive real numbers. Let $\bar{\Delta}\subseteq\mathbb{R}^{+}$ be the sub-semigroup of $\mathbb{R}^{+}$ generated by $\Delta$. Then either $\Delta=\bar{\Delta}$ or $\Delta=\bar{\Delta}\cap(0,{\rm sup}\Delta]$. Let $|\mathcal{M}_{\Delta}|=\{0\}\cup\bar{\Delta}$. Define $0,\leq,+$ in the obvious way. For $q\in\mathbb{Q}^{+}$ and $x,y\in|\mathcal{M}_{\Delta}|$, if either $x=0$ or $y=0$, then $\lnot R_{q}(x,y)$; otherwise, define $R_{q}(x,y)$ if and only if $q<x/y$. Finally, if $\sup\Delta=+\infty$, set $c=0$; otherwise, set $c={\rm sup}\Delta$. Now it is easy to verify that $\mathcal{M}_{\Delta}$ is a model of $T$. Note that clause (7) holds since $\Delta$ is dense in $(0,\sup\Delta)$. It is also clear that the map $\Delta\mapsto\mathcal{M}_{\Delta}$ gives a Borel reduction from $\sim_{\mathsf{CDVS}_{0}}$ to $\cong_{T}$.
Now given a countable model $\mathcal{M}$ of $T$, we define a map $f:\mathcal{M}\to[0,+\infty)$ as follows. We first choose a non-zero element $y_{0}\in\mathcal{M}$ canonically. Then define $f(0)=0$, and for $x\neq 0$, let $f(x)$ be the positive real number defined by the Dedekind cut $\{q\in\mathbb{Q}^{+}\colon R_{q}(x,y_{0})\}$. By clauses (2) and (3), $f$ is an injection. Let $\Delta_{\mathcal{M}}$ be $f(\mathcal{M})\setminus\{0\}$ if $c=0$ and $f(\mathcal{M})\cap(0,f(c)]$ otherwise.
By clause (7), we have $\inf\Delta_{\mathcal{M}}=0$. Thus to verify that $\Delta_{\mathcal{M}}$ is a countable distance value set, it suffices to check that $f(\mathcal{M})$ is closed under summation. For this just note that if $x,y\in\mathcal{M}$ are nonzero, then
$q<f(x+y)$ $\iff R_{q}(x+y,y_{0})$
$\iff\exists q_{1},q_{2}\ q_{1}+q_{2}=q\land R_{q_{1}}(x,y_{0})\land R_{q_{2}}(y,y_{0})$
$\iff\exists q_{1},q_{2}\ q_{1}+q_{2}=q\land q_{1}<f(x)\land q_{2}<f(y)$
$\iff q<f(x)+f(y).$
Thus we have $f(x+y)=f(x)+f(y)$, and so $f(\mathcal{M})$ is closed under summation.
Finally, we verify that the map $\mathcal{M}\mapsto\Delta_{\mathcal{M}}$ is a Borel reduction from $\cong_{T}$ to $\sim_{\mathsf{CDVS}_{0}}$. It is clear that the map is Borel. For the reduction, we first claim that for nonzero $x,y\in\mathcal{M}$, $R_{q}(x,y)$ if and only if $q<f(x)/f(y)$. Suppose $R_{q}(x,y)$. There is $q^{\prime}>q$ with $R_{q^{\prime}}(x,y)$. Let $p<q^{\prime}f(y)$. Then $p/q^{\prime}<f(y)$ and thus $R_{p/q^{\prime}}(y,y_{0})$. Since $R_{q^{\prime}}(x,y)$, by clause (4), we have $R_{p}(x,y_{0})$ and thus $p<f(x)$. This shows $q^{\prime}\leq f(x)/f(y)$ and thus $q<f(x)/f(y)$. Similarly we can prove that if $\lnot R_{q}(x,y)$, then $q\geq f(x)/f(y)$. This finishes the proof of the claim.
Now, given countable models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ of $T$, we denote the functions in the definitions of $\Delta_{\mathcal{M}_{1}}$ and $\Delta_{\mathcal{M}_{2}}$ by $f_{1}$ and $f_{2}$, respectively. Suppose $g$ is an isomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$. Let $\tilde{g}=f_{2}gf_{1}^{-1}$. Then for any $x,y\in f_{1}(\mathcal{M}_{1})$ and any $q\in\mathbb{Q}^{+}$,
$\begin{array}[]{rcl}q<x/y&\iff&R_{q}(f_{1}^{-1}(x),f_{1}^{-1}(y))\\ &\iff&R_{q}(gf_{1}^{-1}(x),gf_{1}^{-1}(y))\iff q<\tilde{g}(x)/\tilde{g}(y).\end{array}$
This means $\tilde{g}(x)/\tilde{g}(y)=x/y$ and thus there is a positive real number $r$ such that $\tilde{g}(x)=rx$ for all $x\in f_{1}(\mathcal{M}_{1})$. Note that $f_{2}(c^{\mathcal{M}_{2}})=\tilde{g}(f_{1}(c^{\mathcal{M}_{1}}))=rf_{1}(c^{\mathcal{M}_{1}})$. Thus either
$\Delta_{\mathcal{M}_{2}}=f_{2}(\mathcal{M}_{2})=r\cdot f_{1}(\mathcal{M}_{1})=r\cdot\Delta_{\mathcal{M}_{1}}$
or
$\begin{array}[]{rcl}\Delta_{\mathcal{M}_{2}}&=&f_{2}(\mathcal{M}_{2})\cap(0,f_{2}(c^{\mathcal{M}_{2}})]\\ &=&r\cdot\left\{f_{1}(\mathcal{M}_{1})\cap(0,f_{1}(c^{\mathcal{M}_{1}})]\right\}=r\cdot\Delta_{\mathcal{M}_{1}}.\end{array}$
In either case, $\Delta_{\mathcal{M}_{1}}\sim\Delta_{\mathcal{M}_{2}}$. Conversely, suppose there is a positive real number $r$ such that $\Delta_{\mathcal{M}_{2}}=r\cdot\Delta_{\mathcal{M}_{1}}$. Then $f_{2}(\mathcal{M}_{2})=r\cdot f_{1}(\mathcal{M}_{1})$ and $f_{2}(c^{\mathcal{M}_{2}})=rf_{1}(c^{\mathcal{M}_{1}})$. Let $g=f_{2}^{-1}rf_{1}\cup\{(0,0)\}$. Then a similar argument shows that $g$ is an isomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$. ∎
Corollary 5.9.
$^{+}\ <_{B}\ \approx_{\mathsf{CDVS}_{0}}\ <_{B}\ =^{++}$.
Proof.
Combining with the results of [2] and Theorem 5.6, Lemma 5.7 implies $^{+}\ <_{B}\ \approx_{\mathsf{CDVS}_{0}}$ and Theorem 5.8 implies $\approx_{\mathsf{CDVS}_{0}}\ <_{B}\ =^{++}$. ∎
We also consider a special kind of countable distance value sets for which the equivalence is more explicit.
Definition 5.4.
Let $\alpha$ be a positve irrational number. Let
$\Delta_{\alpha}=\{p\alpha+q:p,q\in\mathbb{Q}\}\cap(0,+\infty).$
Define an action of $GL_{2}(\mathbb{Q})$ on the space of irrational numbers by the linear fractional transformation:
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot\alpha=\frac{a\alpha+b}{c\alpha+d}.$
We denote the equivalence relation induced by this action by $E_{GL_{2}(\mathbb{Q})}^{IR}$.
For any positive irrational number $\alpha$, $\Delta_{\alpha}$ is a countable distance value set with $\inf\Delta_{\alpha}=0$. The equivalence between distance value sets of the form $\Delta_{\alpha}$ can be characterized by $E_{GL_{2}(\mathbb{Q})}^{IR}$, as the following theorem shows.
Theorem 5.10.
Let $\alpha,\beta$ be positive irrational numbers. Then the following are equivalent:
  1. (1)
    $\Delta_{\alpha}$ and $\Delta_{\beta}$ are equivalent.
  2. (2)
    $\alpha$ and $\beta$ are $E_{GL_{2}(\mathbb{Q})}^{IR}$-equivalent.
Proof.
For the implication of (2) from (1), let $r\in\mathbb{R}^{+}$ be such that the function defined by $f(t)=r\cdot t$ witnesses the equivalence of $\Delta_{\alpha}$ and $\Delta_{\beta}$. Then
$\beta=\frac{\beta}{1}=\frac{f(\beta)}{f(1)}=\frac{a\alpha+b}{c\alpha+d},$
where $a,b,c,d\in\mathbb{Q}$. Note that the matrix formed by $a,b,c,d$ is invertible; otherwise there is an $s\in\mathbb{Q}$ such that $a=sc$ and $b=sd$, and then $\beta=s$, contradicting that $\beta$ is irrational.
For the other implication, suppose $\alpha=({a\beta+b})/({c\beta+d})$, where $a,b,c,d\in\mathbb{Q}$. We can assume $c\beta+d>0$. It is easy to verify that the function $f(t)=(c\beta+d)t$ is a triangle-preserving bijection from $\Delta_{\alpha}$ to $\Delta_{\beta}$. ∎
We note the following fact.
Lemma 5.11.
$E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$.
Proof.
$E_{GL_{2}(\mathbb{Q})}^{IR}$ is induced by a continuous action of a countable group on a Polish space. Thus $E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$. Note that every orbit is dense for this action. It follows from a theorem of Becker–Kechris (see Theorem 6.2.1 of [8]) that $E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}$. ∎
6. Isomorphism of Extremely Amenable Groups in the Framework of Descriptive Set Theory
[…]
Theorem 6.1.
The class $\mathsf{EA}$ of extremely amenable closed subgroups of $S_{\infty}$ is Borel. Consequently, $\cong_{\mathsf{EA}}$ is an analytic equivalence relation.
Proof.
Given any closed subgroup $G$ of $S_{\infty}$, by a well-known construction (see Theorem 4.1.4 of [15] and Section 2 of [17]), we define a canonical structure $\mathcal{A}_{G}$ so that $\mbox{\rm Aut}(\mathcal{A}_{G})=G$. To do this, consider the diagonal action of $G$ on $\mathbb{N}^{n}$ for all postive integers $n\geq 1$, which is defined by
$g\cdot(a_{1},\dots,a_{n})=(g(a_{1}),\dots,g(a_{n})).$
For each $n\geq 1$, let $\mathcal{O}_{1}^{n},\mathcal{O}_{2}^{n},\dots$ enumerate the set
$\left\{G\cdot(a_{1},\dots,a_{n})\colon(a_{1},\dots,a_{n})\in\mathbb{N}^{n}\right\},$
which is the set of all orbits of $G$ for the diagonal action on $\mathbb{N}^{n}$. Consider the language $\mathcal{L}$ consisting of relation symbols $R_{n,i}$ for $n,i\geq 1$. Each $R_{n,i}$ is an $n$-ary relation symbol. The structure $\mathcal{A}_{G}$ is an $\mathcal{L}$-structure whose universe is $\mathbb{N}$, with $R_{n,i}^{\mathcal{A}_{G}}=\mathcal{O}_{i}^{n}$. Then it is easy to check that $\mbox{\rm Aut}(\mathcal{A}_{G})=G$. It is also straightforward to see that $\mathcal{A}_{G}$ is a Fraïssé structure, i.e., it is locally finite, countably infinite and ultrahomogeneous.
We note that the collection of all $\mathcal{L}$-structures with the universe $\mathbb{N}$ forms a Polish space $X_{\mathcal{L}}=\prod_{n,i\geq 1}X_{n,i}$, where each $X_{n,i}$ is a copy of the Cantor space $2^{\mathbb{N}^{n}}$. The map $G\mapsto\mathcal{A}_{G}$ from $\mathcal{F}(S_{\infty})$ to $X_{\mathcal{L}}$ is Borel. In fact, a subbasic open set in $X_{\mathcal{L}}$ is determined by a tuple $\bar{a}=(a_{1},\dots,a_{n})$ being assigned value 0 or value $1$. If $U$ is a basic open set in $X_{\mathcal{L}}$, then whether $\mathcal{A}_{G}\in U$ is determined by finitely many requirements of form $\exists g\in G\ g(\bar{a})=\bar{b}$ and its negation, where $\bar{a},\bar{b}$ are tuples of the same length. Thus $\{G\in\mathcal{F}(S_{\infty})\colon\mathcal{A}_{G}\in U\}$ is clearly a Borel set.
Let $\mathcal{K}$ be the age of $\mathcal{A}_{G}$. Then $\mathcal{A}_{G}=\mbox{Flim}\,\mathcal{K}$. By Theorem 2.2, $G=\mbox{\rm Aut}(\mathcal{A}_{G})=\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$ is extremely amenable if and only if $\mathcal{K}$ has the Ramsey property and every element of $\mathcal{K}$ is rigid. Thus to prove $\mathsf{EA}$ is Borel, it suffices to show both the following:
  1. (i)
    the collection of $\mathcal{L}$-structures $\mathcal{A}\in X_{\mathcal{L}}$ where the age of $\mathcal{A}$ satisfies the Ramsey property is a Borel subset of $X_{\mathcal{L}}$;
  2. (ii)
    the collection of $\mathcal{L}$-structures $\mathcal{A}\in X_{\mathcal{L}}$ where every finite substructure is rigid is a Borel subset of $X_{\mathcal{L}}$.
Both (i) and (ii) refer to finite subsets of $\mathcal{A}$ (whose universe is $\mathbb{N}$). Thus, despite the tediousness of the statements of the Ramsey property and the rigidity, they can be expressed as Borel statements in the topology of $X_{\mathcal{L}}$. This completes the proof. ∎
After the completion of an earlier version of this paper, we learned from Aleksander Iwanow that Theorem 6.1 had been proved by Iwanow–Majcher-Iwanow in response to a question asked by Nies in the Logic Blog in 2020. The proof is included in the manuscript [24] and is different from the one above.
The following theorem is a corollary of our results about countable distance value sets in the preceding section (in particular, Corollary 5.3).
Theorem 6.2.
$E^{\infty}_{S_{\infty}}$ is Borel reducible to $\cong_{\mathsf{EA}}$. In particular, $\cong_{\mathsf{EA}}$ is not Borel.
Proof.
[…]
  1. (i)
    for any distinct $x,y\in|\mathcal{A}|$, there is exactly one $i$ such that $d_{i}>0$ and $R^{\mathcal{A}}_{i}(x,y)$ holds;
  2. (ii)
    for any $x,y,z\in|\mathcal{A}|$ and $i,j,k\in\mathbb{N}$, if $R^{\mathcal{A}}_{i}(x,y)$, $R^{\mathcal{A}}_{j}(y,z)$ and $R^{\mathcal{A}}_{k}(x,z)$, then $|d_{j}-d_{k}|\leq d_{i}\leq d_{j}+d_{k}$;
  3. (iii)
    $^{\mathcal{A}}$ is a linear order.
It is clear that such structures correspond to finite ordered $\Delta$-metric spaces.
Now, given an enumeration of $\mathcal{K}^{<}_{\Delta}$, one can construct the Fraïssé limit in a Borel way. This follows from the existence proof of Fraïssé’s theorem (see, e.g. Lemma 7.1.6 of Hodges [15]). Since the Fraïssé structure is ultrahomogeneous, its automorphism group intersects a basic open set $N_{\bar{a},\bar{b}}$ if and only if the substructures $\bar{a}$ and $\bar{b}$ are isomorphic. Thus the map from the Fraïssé structures to their automorphism groups is Borel. ∎
Corollary 6.3.
There are continuum many pairwise non-isomorphic Polish groups that are extremely amenable. There are continuum many pairwise non-isomorphism separable metrizable groups what are not Polish and are extremely amenable.
Proof.
For both results, note that
$_{\mathbb{R}}\ <_{B}\,E_{0}\ \leq_{B}E^{IR}_{GL_{2}(\mathbb{Q})}\ \leq_{B}\ \approx_{\mathsf{CDVS}_{0}}$
by Lemma 5.11 and Theorem 5.10. Thus there are continuum many pairwise $\approx_{\mathsf{CDVS}_{0}}$-inequivalent elements. Now the first conclusion follows from Theorem 4.3, and the second conclusion follows from Theorems 4.2 and 3.4 (1). ∎
We mention some open problems.
Let us first mention another class of automorphism groups known to be extremely amenable. Recall that an ultrametric is a metric satisfying $d(x,z)\leq\max\{d(x,y),d(y,z)\}$ for any $x,y,z$. An ordering << on an ultrametric space $X$ is convex if for any $x\in X$ and $s>0$, the ball $B(x,s)=\{y\in X:d(x,y)\leq s\}$ is a convex set in the ordering <<, i.e. if $y_{1}<z<y_{2}$ and $y_{1},y_{2}\in B(x,s)$, then $z\in B(x,s)$. An ultrametric space with a convex ordering is called a convexly-ordered ultrametric space.
Theorem 6.4 (Nguyen Van Thé [25]).
Let $S$ be a countable set of positive real numbers. Then the class of all finite convexly-ordered ultrametric spaces with nonzero distances all in $S$ is a Fraïssé class with the Ramsey property.
Denoting the Fraïssé limit of this class by $B_{S}^{c<}$, by the KPT correspondence, ${\rm Aut}(B_{S}^{c<})$ is extremely amenable. We note that if $S$ and $T$ are countable sets of positive real numbers that are order-isomorphic, then ${\rm Aut}(B_{S}^{c<})$ and ${\rm Aut}(B_{T}^{c<})$ are isomorphic as topological groups.
Question 6.5.
Suppose $S$ and $T$ are countable sets of positive real numbers such that ${\rm Aut}(B_{S}^{c<})$ and ${\rm Aut}(B_{T}^{c<})$ are isomorphic as topological groups. Are $S$ and $T$ order-isomorphic?
If the answer is positive, then we would obtain another proof of Theorem 6.2.
We note that there is a prominent difference between Theorems 1.2 and 1.5, namely, in Theorem 1.2 we have the equivalence between the isomorphism of the topological groups and the isomorphism of them as abstract groups, while in Theorem 1.5 we do not have this equivalence. The reason for this equivalence to hold in Theorem 1.2 is that $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ has ample generics (see Corollary 4.1 of Solecki [34]), i.e., $G$ has a comeager orbit in the conjugacy action of $G$ on $G^{n}$, for each $n\geq 1$. By a result of Kechris–Rosendal [18], if a Polish group has ample generics, then every abstract homomorphism into a separable group is automatically continuous and thus it has a unique Polish topology. However, for the ordered case, Slutsky [33] showed that ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ does not have ample generics. Hence the following question remains.
Question 6.6.
[…]
Of course, the main question left open by this paper is the following.
Question 6.7.
What is the exact complexity of $\cong_{\mathsf{EA}}$ in the Borel reducibility hierarchy?
Recall from the Introduction that there are many examples of extremely amenable Polish groups of the form $L_{0}(\mu,G)$, and in particular many of them are abelian. However, there is currently no known method to distinguish their isomorphism types. Thus we still have the following question.
Question 6.8.
Are threre continnum many pairwise non-isomorphic extremely amenable abelian Polish groups?